A Complete Bibliography of Publications in the Bulletin of Mathematical Biology: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

11 March 2024 Version 1.11

Title word cross-reference

+ [118]. 3 [38]. = [118]. _{2} [94, 64]. _{3} [5]. β [334]. f [169]. G [1, 81]. Gα_{i} [1, 81]. Gβγ [1, 81]. IC_{50} [70]. k [107]. κ [551]. N [144]. θ [171].

-Circularity [107]. -diffusion [171]. -Hydroxypropionic [38]. -Statistics [169]. -Strain [144].

/Discrete [411].

1 [194]. 19 [282, 351, 454, 156, 326, 370, 367, 323, 422, 519, 392, 234, 311, 120, 476, 116, 382, 398, 384, 375, 366, 448, 428, 435, 307, 309]. 1BB [551].

2 [331].

3D [327, 295].

5/6 [184].

6 [184].

Ablation [456]. Aboveground [109]. Abundance [452, 541]. Abuse [467, 236]. Accelerated [280]. Account [180]. Accounting [44]. Accumulation [136]. Achieve [337]. Acid [38]. Acidity [238]. Acinar [186]. Across [513, 490, 233]. Activated [185]. Activation [227, 193, 219, 34, 551]. Activator [416]. Active [146, 301]. Actomyosin [192]. Acute [577, 37]. Acyclic [305]. Adaptation [83, 214]. Adapted [323]. Adaptive [355, 531, 124, 52, 126]. Additional [24, 352]. adherence [487]. Adhesion [510]. Adhesion-independent [510]. Adjuvant [332]. Administration [465]. Adult [66]. Adults [263]. Advances [62]. Advancing [344]. Advective [264]. Affect [523, 16]. Affected [300]. African [311]. Against [302, 14]. Agar [483]. Age [464, 450, 245, 14, 364, 349, 516, 137, 384, 463, 263, 337]. Age-Dependent [14]. Age-Related [364, 137]. Age-Specific [384, 349]. Age-Structured [450, 245, 463, 337]. Agent [557, 103, 491, 258, 514, 348, 522]. Agent-Based [557, 103, 491, 258, 514, 348, 522]. Aggregate [484]. Aggregation [574, 389]. Agnostic [244]. Aided [347]. AIDS [226, 210, 463, 408]. Airborne [339]. al [54]. Alcohol [427]. Algae [53]. Algal [485]. Algebra [413]. Algebraic [117, 4]. Algebras [153]. Algorithms [443]. Allee [76, 336, 508, 543, 84]. Allopolyploidy [462]. Allow [224]. Along [157, 164]. Alopecia [36]. Alpha [577]. Altered [15]. Alternative [35, 297, 255, 244]. Amyloid [521]. Anaerobic [277]. Analysing [381]. Analysis [151, 2, 412, 176, 68, 232, 558, 38, 36, 97, 313, 293, 148, 328, 306, 370, 363, 49, 523, 422, 358, 405, 5, 327, 295, 479, 281, 437, 400, 32, 189, 343, 551, 48, 433, 487, 101, 110, 435, 239]. Analytical [547, 24, 393]. Ancestral [369, 516, 23]. Anchoring [525]. Andalusia [476]. Androgen [545]. Angiogenesis [449, 25, 466]. Anti [459, 417, 214]. Anti-inflammatory [459]. Anti-predation [214]. Anti-tumour [417]. Antibiotic [335, 8, 330]. Antibiotics [407]. Antibody [240, 161, 15, 278]. Antibody-Dependent [240, 15]. Antigen [195]. Antigen-Competitive [195]. Antimicrobial [547, 191]. Antiretrovial [527]. Antiviral [31]. Anything [89]. Apocalypse [177]. Apparent [206]. Application [489, 127, 14, 304, 427, 548]. Applications [228, 516, 202, 174]. Approach [282, 417, 557, 312, 538, 561, 11, 147, 77, 482, 208, 311, 124, 247, 150, 91, 384, 393, 380, 373]. Approach-Avoidance [538]. Approaches [111, 176]. Approximate [115, 517]. Approximating [283]. Approximation [346, 286]. Approximations [504, 21]. Arabidopsis [225, 172]. Arcs [462, 305, 352]. Areas [577]. Areata [36]. Arise [550]. Array [246]. Array-RQMC [246]. Arrest [356, 496]. Arteriole [472]. Asbestos [209]. Asia [368]. Aspects [550]. Assay [576]. Assess [160, 37]. Assessing [162, 359, 536, 384, 294]. Assessment [427, 478]. Assimilation [156]. Association [188]. Assortment [402]. Astrobiological [228].

Asymmetric [184]. Asymptomatic [44, 474, 317]. Asymptotes [5]. Asymptotic [38, 363, 523, 437, 574]. Asynchronous [546]. Atherosclerosis [251]. Atherosclerotic [459, 507]. ATM [193]. ATR [193, 258]. ATR-Inhibiting [258]. Atrial [456]. Attention [71]. Atto [259]. Atto-Foxes [259]. Attractor [250]. Attractors [84]. aureus [373]. Auto [48]. Auto-Regulated [48]. Automata [105]. Autophosphorylation [219]. Autopolyploidy [462]. Autoregulation [479]. Autotroph [284]. Auxin [19]. Averaging [489]. Avoidance [538]. Ayala [396].

B [82, 551]. Babesiosis [66]. Back [323]. Background [383]. Bacteria [165, 284, 407]. Bacterial [506, 203, 29, 457, 288, 485]. Bactericidal [285]. Bacterium [213]. Balance [313, 118]. Balancing [355]. Barrier [527]. Based [111, 98, 227, 557, 103, 490, 403, 450, 72, 533, 491, 258, 483, 213, 395, 440, 499, 320, 477, 134, 446, 514, 554, 283, 80, 493, 348, 522, 389, 396, 368, 326]. Basic [317, 244, 113, 372]. Basin [255]. Basis [304, 136]. Bat [562]. Bayesian [489, 178, 281, 248]. Bayesian-Weighted [248]. be [562, 497]. Bed [254]. Bee [222, 534]. Beetle [530, 220, 409]. Behavior [513, 420, 403, 304, 427, 301, 3]. Behaviour [459, 423, 283, 382, 275]. Behavioural [16]. Belowground [109]. Benefits [455, 69]. Berry [5]. Bespoke [196]. Best [302]. Bet [52]. Bet-Hedging [52]. Better [517]. Between [162, 359, 240, 20, 206, 193, 107, 315, 574, 350, 439, 48, 204, 448, 428]. Between-Host [428]. Bias [505]. Bifurcation [390, 97, 293, 416, 377]. Bilayer [559]. Binary [336]. Binding [432]. Biochemical [470]. Biofilm [203, 42, 215, 165, 29, 324]. Biofilms [260, 485]. Biogenesis [137]. Biogeochemistry [228]. Biological [512, 533, 345, 499, 121, 574, 246, 174, 235, 424]. Biology [62, 111, 103, 322, 61, 119, 122, 117, 292, 65, 129, 100]. Biomarker [223]. Biomolecular [155]. Biomolecules [110]. Biophysical [383, 319]. Bioreduction [260]. Biosorption [485]. Bipolarity [308]. Birth [516]. Births [571]. Bistable [300]. Biting [491, 14, 301]. Black [256]. Blockades [261]. Blood [527, 361, 205]. Bluetongue [491]. Bone [136, 434]. Boolean [276, 451, 511]. Borne [96, 452, 400, 301, 310]. Borrelia [221]. Boundaries [313, 533, 140]. Boundary [183, 79, 313, 363, 150, 377]. Bovine [276, 66]. BRAF [10]. Brain [527, 558, 194]. Branches [560]. Branching [564, 294]. Brazil [14]. Breaks [13]. Breast [566, 388, 429, 553, 434]. Bridging [333]. Brief [539]. Brugada [444]. Buffers [432]. Building [111, 262, 372]. Bulb [383]. Bulk [32].

C [5]. C. [319]. CA1 [531]. Caemmerer [5]. Calcium [432, 185, 186]. Calcium-Activated [185]. Calculate [11]. Calibrated [258]. cAMP [81, 1]. Campus [435]. Can [28, 406, 181, 325, 5, 423, 562]. Cancer [566, 388, 105, 429, 553, 406, 93, 238, 22, 258, 147, 494, 358, 430, 281, 545, 270, 480, 67, 415, 434, 17, 334, 564, 578, 469]. Capillaries [472]. Capillary [130]. Capsular [553]. Capturing [20]. CAR [566, 197, 334, 551]. CAR-4-1BB-Mediated [551]. CAR-T [334]. Carcinoma [223]. Cardiac [456]. Cardiomyocytes [243]. Care [8, 127]. Carnivores [526]. Carrier [278]. Carriers [44]. Cartilage [20]. Cascades [208, 267]. Case [102, 467, 486, 117, 74, 379, 442, 368]. Cases [75]. Cassava [242]. Catalyzed [286]. Catheter [456]. Cattle [66]. Cause [377]. Caused [370, 478]. CD200 [358]. CD200R [358]. Celebrating [292]. Cell [417, 447, 83, 520, 30, 390, 217, 501, 321, 42, 421, 510, 22, 492, 199, 523, 181, 482, 51, 515, 13, 37, 132, 430, 197, 379, 184, 67, 415, 192, 564, 223]. Cell-Driven [415]. Cell-Scale [67]. Cells [566, 390, 544, 383, 97, 362, 149, 1, 152, 230, 186, 360, 334, 86, 81]. Cellular [20, 105, 287, 258, 188, 528]. Center [149, 93]. Centre [134]. Centre-Based [134]. centrosomal [308]. Cephalosporin [373]. Cephalosporin-Resistant [373]. Chagas [43]. Chain [456]. Chains [118]. Challenges [62, 119, 212]. Change [63, 530, 427, 382]. Changed [89]. Changes [119, 364]. Changing [62, 129]. Channel [548]. Channels [185, 27]. Characterization [524]. Characterizing [378, 106]. Characters [12]. Checkpoint [261]. Chemical [461, 470, 498, 231, 246]. Chemostat [78, 159]. Chemostats [386]. Chemotactic [520, 419, 389]. Chemotaxis [321, 550, 347, 41, 288]. Chemotherapies [429]. Chemotherapy [106, 545]. Chemotherapy-Induced [106]. Child [556, 401]. China [82, 370, 75, 463, 337, 408]. Chinese [309]. Chinese-Spring-Festival [309]. Chlamydia [179]. Cholera [477, 74]. Cholesterol [137]. Chromosomes [385]. Chronic [139, 166, 330]. CI [371]. Circadian [167]. Circuit [189]. Circular [50, 446, 529]. Circularity [107]. Circulating [162, 359, 221]. Circulation [449, 338, 522]. City [522]. Class [113]. Clean [201]. Clearance [34]. Climate [530, 28, 280]. Climate-Induced [28]. Clinical [565]. Clonal [430]. Close [22]. Closed [143]. Closures [448, 428]. Clot [85]. Cluster [401, 319, 380]. Clustering [308, 379, 32]. Co [221, 144, 554]. Co-circulating [221]. Co-colonization [144]. Co-Culture [554]. CoA [38]. Coagulation [205]. Coastal [569]. Cocktails [77]. Cocoa [256]. Code [50, 446]. Codes [50, 107]. Coding [502]. Codon [446]. Coevolutionary [450]. Coexistence [455, 268, 144, 17]. Coexisting [155]. Coinfection [271, 15, 442]. coli [451, 483, 373]. Collection [292]. Collective [217, 127, 327, 295, 69, 92, 473]. College [63, 435, 89]. Colonies [222, 534]. colonization [144]. Colony [483]. Coloured [46]. Combination [31, 429]. Combinatorial [500]. Combinatorics [443]. Combined [160, 216]. Combining [356]. Commentary [54]. Communities [221, 119]. Community [111]. Community-Based [111]. Commuters [425]. Compact [232]. Comparing [431, 243, 8]. Comparison [70, 461, 233, 499, 55, 320, 391, 281, 427, 254]. Comparisons [540]. Compartment [554, 487]. Compartment-Based [554]. Compartmental [329, 234, 393]. Compartments [509, 523]. Compatible [12]. Compete [127]. Competition [206, 330, 560, 207, 177, 132, 484, 316]. Competitive [195, 168, 204, 159]. Complement [35, 297]. Complex [426, 95, 41, 118, 253, 247]. Complexity [391]. Composition [221, 77]. Computational [566, 1, 81, 69, 410, 399, 294, 473]. computationally [23]. Compute [24]. Computer [92]. Concentration [436]. Concentration-Dependent [436]. Condensates [155]. Condition [472]. Conditionally [27]. Conditioning [236]. Conditions [79, 470, 377]. conducting [149]. Confined [510]. Confinement [180]. Conflict [412, 538]. Congo [412]. Connecting [557, 100]. Connections [119]. Consider [301]. Considerations [52]. Consistent [125, 23]. Constant [280]. Constant-Speed [280]. Constitutive [235]. Constraints [57]. Constrictions [143]. Consumer [378, 90]. Contact [22, 272]. Contagion [445, 403]. Context [60]. Continuation [512, 8]. Continuous [411]. Continuous- [411]. Continuous-/Discrete-Time [411]. Continuum [417, 514]. Contraction [410, 399]. Contracture [553]. Contribution [266]. Contributions [292]. Control [112, 47, 7, 326, 201, 213, 453, 218, 272, 77, 289, 517, 37, 120, 270, 511, 541, 373, 525, 398, 4, 384, 521, 575, 163, 307]. Control-Theoretic [77]. Controlling [187, 382, 424]. Conversion [515]. Cooperation [402]. Coordinated [69]. Coordination [92]. Coral [325, 481]. Corals [69, 92]. Coronavirus [577]. Correction [359, 297, 239, 274, 225, 324, 166, 257, 94, 81, 226, 327, 439, 410, 448]. Corrections [183]. Cortical [525, 319]. Cottontail [206]. Could [497]. Countable [564]. Countable-Type [564]. Counter [332]. Counting [249]. Countries [453, 311]. Coupled [217, 271]. Coupled-Cell [217]. Coupling [530, 229, 550, 267, 545, 410, 399, 60, 428, 448]. Course [16]. CoV [331]. Covert [222]. COVID [54, 282, 454, 156, 326, 370, 367, 422, 519, 392, 234, 311, 120, 476, 116, 382, 398, 384, 375, 366, 448, 428, 307, 309, 351, 323, 435]. COVID-19 [54, 282, 454, 156, 326, 370, 367, 422, 519, 392, 234, 311, 120, 476, 116, 382, 398, 384, 375, 366, 448, 428, 307, 309, 351, 323, 435]. Crampin [314]. Crisis [467]. criterial [488]. Critical [448, 428]. Cross [221, 559, 546, 503, 115, 114]. Cross-Diffusion [559, 115, 114]. Cross-Immunity [546, 503]. Cross-Reaction [221]. Crosstalk [193]. Crowd [445]. CSF [79]. Cue [321]. Culture [554]. Cultures [520, 70]. Cumulants [21]. Curbing [307]. Current [25]. Cycle [276, 199, 13, 118]. Cycles [329, 498, 546]. Cyclic [341, 207]. Cytoplasmic [184]. Czech [351].

D [181, 292]. Dairy [191]. Damage [193, 258, 136]. Darn [517]. Data [557, 565, 490, 403, 176, 486, 156, 475, 345, 552, 121, 484, 252, 129, 427, 174, 278, 463]. Data-Driven [252]. Data-Validated [475]. Datasets [80]. DDE [371]. De-escalation [8]. Death [354, 516, 469]. Death-Rate [469]. Deaths [571]. Decision [225, 172]. Decisions [315]. Decompositions [231]. Decoys [43]. Deep [174]. Defense [378]. Deformations [51]. Degenerating [575]. Degeneration [137]. Degradation [576, 67]. Dehydration [483]. Delay [276, 544, 536, 493, 236]. Delayed [48, 261]. Delays [351, 31, 374, 216].

Delineating [251]. Delivery [278]. Delta [577]. Demand [454, 54]. Democratic [412]. Demographic [154]. Dendritic [185]. Dengue [26, 240, 546, 414, 14, 15, 163]. dense [23]. Density [526, 2, 239, 370, 263, 3]. Density-Dependent [2, 239, 263]. Dependent [526, 2, 239, 240, 508, 436, 14, 15, 91, 507, 263]. Depletion [37]. Derivation [131, 502, 343]. Derivatives [414]. Dermal [518, 139, 166]. Describe [482]. Described [562]. Description [418, 288]. Descriptions [247]. Design [379]. Designing [490, 299]. Designs [548]. Detailed [118]. Determination [91]. Determine [41, 69, 473]. Deterministic [320, 391]. Development [459, 449, 335, 251, 273, 528]. Developmental [216]. Diabetes [139, 166]. Diel [471]. Differences [290, 370]. Different [249, 344, 233, 377]. Differential [36, 535, 469]. Diffusion [46, 157, 128, 559, 7, 550, 483, 457, 115, 237, 436, 198, 543, 471, 484, 374, 53, 114, 204, 60, 200, 113, 396, 18, 171]. Diffusive [123, 192]. Dilemma [68]. Dilution [43]. Dimension [250]. Dimensional [557, 232, 215, 324, 260, 380]. Dimensionally [42]. Dimensions [449]. Dimorphism [563]. Direct [532]. Directionality [150]. Discrepancy [548]. Discrete [417, 378, 403, 323, 343]. Discrete-Time [411]. Discretization [568]. Disease [445, 242, 431, 222, 464, 240, 256, 201, 573, 571, 182, 453, 296, 272, 104, 320, 577, 349, 154, 244, 301, 88, 310, 382, 333, 60, 43]. Disease-Induced [296, 272]. Diseases [298, 96, 339, 400]. Disorder [427]. Dispersal [387, 175, 133, 264, 569, 348, 416]. Dispersing [28]. Dispersion [444]. Disruption [480]. Dissection [83]. Dissipation [27]. Dissipative [85]. Distal [145]. Distance [125, 99]. Distancing [355]. Distinct [401]. Distinct-Cluster [401]. Distinguishing [574]. Distributed [354, 374, 394]. Distribution [483, 152, 300, 516]. Distributions [516, 484]. Disturbance [268]. Disturbances [398]. Divergence [516]. Diverse [76]. Diversification [344, 24]. Diversity [175, 500]. Division [184]. DNA [23, 193, 258, 13, 86]. Does [249]. Domain [436]. Domains [138]. Dormancy [225, 172]. Dormancy-Germination [172]. Dose [37]. Double [13]. Double-Strand [13]. Drift [264]. Drive [430]. Driven [145, 73, 27, 143, 252, 415, 192, 216]. Driver [546, 129]. Drug [465, 70, 299, 10, 278, 487]. Drugs [258, 194, 236]. Drusen [137]. Dual [189, 478]. Duct [360]. Due [548, 45]. Dues [61]. During [563, 227, 420, 287, 141, 194]. Dynamic [160, 188, 71, 48]. Dynamical [290, 91]. Dynamics [489, 131, 445, 282, 96, 242, 151, 520, 555, 526, 532, 222, 530, 20, 412, 287, 68, 36, 79, 450, 486, 539, 179, 156, 76, 97, 362, 502, 229, 238, 421, 322, 491, 336, 345, 11, 572, 571, 367, 182, 115, 219, 456, 226, 210, 198, 560, 577, 9, 161, 514, 349, 124, 283, 143, 195, 168, 476, 80, 69, 400, 254, 316, 186, 88, 371, 310, 382, 114, 398, 236, 85, 333, 375, 366, 564, 279, 204, 284, 458, 101, 309]. Dysfunction [410, 399].

E. [483]. Early [404, 319]. Eastern [206]. Ebola [412, 104]. Eco [328].

Eco-evo [328]. Ecological [262, 427, 481]. Ecology [303, 144]. Economic [325]. Ecosystem [151, 140, 90, 522]. Ecosystems [262]. Edge [98]. Edge-Based [98]. Edmund [314]. Educate [122]. Education [62, 111, 103, 61, 119, 117, 129]. Effect [96, 404, 180, 420, 222, 206, 412, 501, 317, 518, 339, 95, 475, 508, 296, 432, 226, 210, 543, 104, 410, 399, 338, 84, 126, 263, 408, 389, 3, 381, 473, 368, 309, 275]. Effects [242, 472, 447, 455, 70, 183, 383, 76, 466, 336, 331, 75, 494, 325, 194, 116, 310, 578, 60, 200, 481]. Efficacy [342]. efficient [23]. Efflux [332]. Efflux-Mediated [332]. Effort [495]. EGFR [223]. EGFR-Mutated [223]. Egg [555]. Elastic [51]. Elasticity [447, 501]. Elderly [351]. Electrical [130]. Electro [149]. Electro-physiology [149]. Electrodiffusion [185]. elegans [319]. Elimination [393]. Elongation [158]. Embedding [493]. Embryo [181]. Embryogenesis [319]. Emergence [503, 154]. Emergent [506, 327, 295, 114]. Empirical [548]. Encompassing [508]. Encounters [22]. Endemic [220, 266]. Endosymbiotic [213]. Engineered [38, 363]. Engineering [447]. Engineers [151, 140]. England [467]. Enhanced [45]. Enhancement [240, 15]. Enough [567]. Ensemble [548]. Ensure [567]. Entrainment [167]. Environment [214, 348]. Environmental [392, 481]. Environments [510, 264, 27, 391, 171]. Enzyme [286]. Epicardial [444]. Epidemic [187, 404, 426, 180, 420, 474, 540, 232, 317, 339, 450, 156, 342, 354, 344, 355, 367, 218, 440, 552, 272, 422, 519, 289, 154, 124, 488, 39, 252, 476, 253, 382, 478, 398, 16, 433, 279, 408, 113, 458, 381]. Epidemics [98, 157, 346, 201, 493, 463]. Epidemiological [44, 229, 306, 173, 271, 80]. Epidemiology [513, 403, 144, 452]. Epigenetic [293]. Epilimnion [284]. Epistatic [350, 439]. Epithelial [132]. Equation [36, 543, 144, 208, 124, 389]. Equations [276, 9, 5, 121, 484, 235, 535, 469]. Equilibria [369]. Erdos [346]. ERK [413]. Erythrocytic [298]. escalation [8]. Escaping [503]. Escherichia [451, 373]. Estimate [250, 178]. Estimates [249, 116]. Estimation [532, 123, 125, 256, 73, 211, 477, 484, 74, 244, 407]. Estimators [512]. Estrous [276]. European [206]. Evacuation [445]. Evaluating [337]. Evaluation [441, 307]. Evaporation [73]. Evaporation-Driven [73]. Events [529]. evo [328]. Evolution [420, 464, 385, 50, 466, 61, 133, 264, 436, 214, 430, 316, 263]. Evolutionarily [302]. Evolutionary [238, 33]. Ex [537]. Exacerbate [364]. Exact [115]. Examine [60]. Example [203]. Excitability [243]. Excitation [410, 399]. Exclusion [159]. Exercise [563]. Exhaustion [528]. Exhibits [12]. Exosomal [434]. Expand [313, 122]. Expected [477]. Experience [93, 322]. Experimental [379, 548, 278]. Experiments [209, 199, 537, 554]. Explains [136]. Explore [417, 367]. Expose [122]. Exposed [407]. exposure [87]. Expression [374]. Extended [362, 421]. Extinction [249, 503, 300, 320, 477, 394]. Extracellular [51, 67].

Facilitate [434]. Facilitator [17]. Factors [75]. Family [76]. Far [41].

Far-Field [41]. Farm [191]. Farms [191]. Farquhar [5]. Fast [157, 274, 265, 340]. fastidiosa [489]. Fear [391, 408]. Features [361]. Feedback [502]. Feeding [571]. Fentanyl [252]. Ferguson [54]. Fertility [112, 310]. Festival [309]. Fever [26]. Fiber [321]. Fibre [67]. Fibrillation [456, 521]. Fibrin [202]. Field [346, 466, 272, 41, 471]. Filamentous [176]. Film [73, 211]. Final [474, 317]. Finite [568]. Fire [268, 531]. First [351, 63]. Fish [2, 239]. Fisher [313]. Fitness [387, 304]. Fitting [400]. FitzHugh [421, 243]. Five [93]. Five-Year [93]. Fixed [374]. Flow [361, 42, 418, 347, 85, 319]. flowered [148]. Flowering [148]. Flu [141, 315]. Fluctuating [83]. Fluctuations [158, 430]. Fluid [347, 178, 69, 348]. Fluorescence [505]. Flux [19, 506]. Flux-Regulation [506]. Focussed [250]. Forager [534]. Force [134]. Forced [543]. Forecast [408]. Forecasting [489, 431, 493, 366]. Forest [530, 395]. Forest-Based [395]. Form [532, 155, 235]. Formation [131, 274, 265, 559, 362, 550, 6, 146, 567, 207, 554, 549, 434, 58, 85, 416]. Formulations [134]. Four [577]. Fourth [79]. Foxes [259]. Fractional [414, 572, 469]. Frame [446]. Framework [538, 56, 178, 169, 80, 137, 348]. Free [313, 533, 333]. Freshwater [245]. Frogging [562]. Front [426]. Fronted [190]. Fronts [344]. Frozen [95]. Fully [28]. Function [134]. Functions [403]. Fundamental [547]. Future [262].

Game [82, 141, 494]. Game-Theoretic [141]. Game-Theoretical [82]. Games [33, 471]. Gansu [75]. Gaussian [345]. Gelatin [576]. Gene [365, 374, 568, 511, 294]. General [68, 159]. Generalist [424]. Generalized [534, 572, 228, 531]. Generation [556, 11, 373]. Generic [377]. Genes [446, 48]. Genetic [50, 175, 48, 332, 430]. Genetically [38]. Genetics [33]. Genomes [529]. Geographic [370]. Geometric [500, 304]. Geometrical [169]. Geometries [41]. Geometry [559, 149, 413]. Germination [225, 172]. Getting [108, 519]. Giffen [275]. Gilpin [396]. Give [76]. Given [443]. Gland [360, 40]. Glioblastoma [145, 45]. Global [198, 435]. Glossina [249]. Glucose [480]. Glycolytic [376]. Go [28]. Goals [337]. Gonorrhea [87]. Good [175, 517]. Goods [127]. Granular [485]. Granulation [277]. Graph [118]. Graph-Theoretic [118]. Grass [268]. Grassland [90]. Grazer [57]. Grip [108]. Gromov [499]. Group [378, 561]. Growing [58]. Growth [19, 506, 356, 496, 45, 406, 203, 42, 215, 324, 22, 483, 363, 29, 523, 41, 554, 380, 396]. Growth-Migration [380]. GTPase [30, 189]. Guide [379]. Guinea [104].

H1N1 [364]. Habitat [387, 313, 175, 543, 391]. Hajj [162, 359]. Hard [282]. Hare [206]. Harnessing [212]. Harvesting [526, 263, 396, 481]. Healing [518, 25, 202]. Health [325]. Healthcare [454, 54]. Heat [293]. Heat-Sensitive [293]. Hedging [52]. Hematocrit [472]. Hemorrhagic [26]. Hepatitis [82]. Herbaceous [148, 109]. Herbivore [59]. Herbivores [526, 255, 109]. Herbivores-Induced [255]. Herd [272, 392]. Heroin [252].

Heterogeneity [423, 520, 387, 572, 494, 419, 473]. Heterogeneous [490, 132, 311, 171, 433, 216]. Heterosexual [160]. Hierarchical [502]. High [557, 250, 104, 93]. High- [104]. High-Dimensional [557]. Highlights [332]. Highs [336]. Hill [393]. HIP [93]. Hippocampal [531]. HIV [226, 527, 173, 210, 194, 478, 236, 442, 463, 408]. HIV-1 [194]. HIV/AIDS [226, 210, 463, 408]. Hive [534]. Hodge [110]. Homeostasis [480]. Homogeneous [269, 317]. Honey [222]. Honey-Bee [222]. Horizon [52]. Horizontal [462, 245, 352]. Hormone [147, 523, 281]. Host [221, 450, 306, 344, 503, 452, 438, 236, 448, 428, 216, 424]. Hosts [43]. Household [326]. Human [167, 420, 339, 209, 160, 87]. Humans [298, 406]. Hybrid [417, 20, 367, 29, 327, 295, 542]. Hydra [296]. Hydroxypropionic [38]. Hyperbola [5]. Hyperpredation [206]. Hypothesis [333]. Hypothesis-Free [333]. Hypoxia [45, 238]. Hypoxia-Induced [45].

Identifiability [558, 440, 318]. Identifiable [329]. if [69]. II [115, 69]. Imaging [223, 199]. Immersed [348]. Immigrating [398]. Immune [221, 417, 524, 450, 1, 81, 537, 234, 364, 515, 37, 273, 195, 528, 126, 473, 261]. Immunity [26, 546, 503, 272, 392]. Immuno [306, 173, 271]. Immuno-Epidemiological [306, 173, 271]. Immunodeficiency [142, 87]. Immunology [74]. IMO [93]. Impact [162, 359, 417, 249, 112, 520, 527, 262, 54, 326, 22, 191, 272, 419, 51, 134, 65, 452, 87, 373, 279, 43]. Impacts [577, 391]. Imperfect [488, 371, 366]. Implant [553]. Implant-Induced [553]. Implications [194]. Importance [234, 318]. Impossibility [125]. Improves [569]. Impulsive [268, 519, 545]. Inactive [301]. Incentivize [335]. Income [453]. Incompatibility [233]. Incompetent [43]. Incomplete [371]. Incorporating [309]. Increase [530]. Increased [254]. Incubation [96]. Independent [231, 510]. Indices [500]. Indirect [339]. Individual [533, 315]. Individual-Based [533]. Individuals [317, 104, 427]. Induce [336]. Induced [106, 387, 28, 45, 553, 250, 25, 421, 296, 272, 208, 255, 57]. Infants [26]. Infected [397, 310]. Infection [240, 226, 210, 234, 364, 514, 452, 194, 368]. Infections [222, 37]. Infectious [445, 431, 464, 573, 453, 88]. Infectivity [15]. Inference [152, 248, 441, 23]. Inferring [552, 99]. Infiltration [417]. Inflammation [227, 47]. Inflammation-Based [227]. Inflammatory [364, 459]. Inflows [46]. Influence [423, 224, 496, 238, 510]. Influenza [162, 359, 123, 75, 537]. Information [92, 551]. Information-Theoretic [551]. Inheritance [86]. Inhibited [330]. Inhibiting [258]. Inhibition [78, 165, 22, 334]. Inhibitor [10, 421, 416]. Inhibitor-Induced [421]. Inhibitors [223]. Inhibitory [70, 48]. Inhomogeneous [164]. Initial [300]. Initiation [158]. Injured [85]. Innate [537, 234]. Inoculum [37]. Insect [112, 255]. Insight [382]. Insights [566, 342]. Instabilities [376, 567]. instantaneous [389]. Insulin [405]. Integrate [531]. Integrate-and-Fire [531]. Integrated [93, 373]. Integrating [88]. Integrodifference [55, 9]. Intensity [250]. Intensive [8]. Interacting [509]. Interaction [109].

Interactions [20, 287, 508, 144, 255, 350, 439, 279, 284, 473]. Interactive [535]. Interconnected [386]. Interface [192]. Interfacial [533]. Interference [278]. Intermittent [183, 508, 147, 281, 545]. Interneurons [531]. Internship [93]. Interplay [566, 15, 315]. Interruption [465]. Intertaxon [99]. Intervention [414, 398, 463]. Interventions [540, 54, 120]. Intracellular [31, 128, 504, 236, 4]. Intrinsic [183]. introduction [372]. Introductory [89]. intuition [372]. intuition-building [372]. Invading [190]. Invadopodia [576]. Invasion [181, 392, 67, 415, 17, 150]. Invasions [424]. Invasive [206, 168]. Invasiveness [380]. Investigate [191]. Investigating [496]. Investigation [47, 356, 145]. Ion [548]. Ischemia [145]. Ischemia-Driven [145]. Ischemic [227]. Isolating [237]. Isolation [342, 218].

J. [181, 292]. Jab [519]. John [314]. Joint [307]. Journals [102]. Juvenile [66, 263].

Karma [243]. Keloids [539]. Kenya [226, 210]. Kermack [323]. Key [17, 497]. Kidney [287]. Kill [194]. Killer [230]. Kinase [223]. Kind [22]. Kinesin [561]. Kinetic [445, 321, 572, 568, 389, 458]. Kinetics [95, 94, 64, 413, 202, 548, 507]. KPP [313]. Kuznetsov [406].

Labellable [468]. Lag [506]. Lake [245]. Landscape [62]. Landscapes [3]. Large [131, 95]. Large-Scale [131]. Larvae [555, 569]. Latency [194, 116]. Latent [96]. Later [177]. Law [21]. Lck [219]. Leaders [224]. Leadership [224]. Leads [438]. Leaf [5]. Leaky [531]. Leap [562]. Leaping [246]. Learning [538, 345, 121, 174]. Length [2, 239, 152]. Length-Structured [2, 239]. Lengths [125, 152]. Leone [104]. Lesions [36]. Lessons [282]. Levels [83, 407]. Liberia [104]. Life [249, 89]. Lifestyles [56]. Light [225, 172, 537]. Likelihood [512, 24, 12]. Likely [530]. Limit [498]. Limited [331, 121, 101]. Lineages [490, 369, 516]. Linear [329, 128]. Lines [157]. Linking [382]. Lipid [507]. Lipid-Structured [507]. Lipolysis [504]. Living [95]. Local [142, 321, 558, 325, 279, 115, 114, 416]. Locomotion [183]. Logistic [21]. Long [378, 209, 574, 84, 408]. Long-Term [209, 408]. Long-Transient [574]. Longitudinal [427]. Losses [254]. Lotka [486, 560]. Low [232, 453, 104]. Low-Dimensional [232]. Low-Income [453]. Low-risk [104]. Lows [336]. LSD1 [170]. LSD1-Mediated [170]. Luck [569]. Lung [223]. Lungs [209]. Lusaka [477]. Lyme [571]. Lymph [209, 418].

Macaques [234]. Macdonald [49]. Macrophage [459]. Macrophages [150, 507]. Macroscale [247, 60]. Macular [137]. Maintenance [152]. Making [569]. Malaria [44, 465, 34, 161]. Males [397, 127, 310]. MalFGK [94, 64]. Malonyl [38]. Maltose [94, 64]. Management [262, 141, 453, 488]. Manipulation [332]. Mapping [486]. Marine [11, 569]. Markov [336, 561, 456, 118]. Markovian [536, 27]. Mask [180]. Masks [351]. Mass [465]. Master [208]. Match [17]. Material [178]. Maternal [26, 371]. Mathematical [111, 162, 359, 19, 357, 83, 520, 35, 47, 388, 524, 565, 106, 206, 356, 429, 539, 93, 502, 238, 139, 418, 61, 414, 153, 94, 64, 49, 199, 119, 188, 56, 353, 122, 358, 292, 161, 554, 288, 252, 379, 168, 485, 480, 137, 230, 100, 427, 434, 360, 334, 373, 86, 433, 380, 332, 396, 261, 297, 166]. Mathematics [103, 63, 89, 497]. Matrix [518, 51, 67]. Maturation [242]. Mature [158]. Maximal [495]. Maximum [512]. May [508, 344]. McKendrick [323]. MDR [332]. Mean [346, 272, 471, 304]. Mean-Field [346, 272]. Measures [355, 272]. Mechanical [30, 308, 132, 235]. Mechanics [510, 482, 525]. Mechanism [211, 307]. Mechanisms [47, 224, 356, 496, 402, 155]. Mechanistic [145]. Media [260, 88, 212, 247]. Mediated [170, 551, 332]. Medical [326, 331, 101]. Meeting [62]. Meets [19, 144]. Membrane [576, 32]. Memory [273, 126]. Memoryful [27]. Men [160]. Mendelian [33]. Mentoring [353]. Merging [305]. Mesic [549]. Meta [317]. Meta-Population [317]. Metabolic [544, 405]. Metabolism [563, 451, 275]. Metaecoepidemic [90]. Metal [485]. Metalloproteinases [518]. Metapopulations [175, 11]. Metastatic [434]. Meteorological [75]. Methicillin [373]. Methicillin-Resistant [373]. Method [24, 441, 85]. Methods [490, 248, 4]. Methylation [86]. Metric [99]. Metropolitan [577]. MEX [184]. MEX-5 [184]. MEX-5/6 [184]. Micro [178]. Micro-swimmer [178]. Microbes [38]. Microeconomics [275]. Microenvironment [515, 270]. Microglia [227]. microRNAs [434]. Microscale [247, 60]. Microtubule [525]. Microtubules [164]. Microvascular [361, 572]. Midge [491]. Migration [321, 45, 510, 331, 51, 471, 90, 380]. Mimulus [542]. Minimal [123, 558, 460]. Minutiae [259]. Mismatch [162, 359]. Mitigate [527, 299, 435]. Mitigation [463, 163]. Mitochondrial [410, 399]. Mitotic [308]. Mitral [383]. Mix [17]. Mixed [455, 302, 211]. Mixed-Mechanism [211]. Mixing [317, 272]. Mixotroph [284]. Mixture [135]. Mobility [339, 476, 216]. Model [489, 131, 445, 44, 459, 98, 167, 357, 180, 555, 31, 2, 239, 20, 206, 105, 123, 82, 544, 576, 223, 321, 402, 385, 232, 340, 383, 317, 558, 429, 339, 79, 450, 346, 250, 406, 156, 125, 97, 362, 313, 326, 158, 256, 148, 42, 215, 324, 139, 166, 421, 160, 475, 342, 354, 491, 306, 173, 258, 492, 94, 64, 572, 268, 331, 367, 182, 363, 29, 49, 199, 218, 245, 146, 115, 323, 191, 272, 198, 422, 397, 519, 289, 477, 207, 531, 177, 34, 161, 5, 234, 13, 391, 90, 327, 295, 21, 74, 244, 281]. Model [545, 252, 574, 425, 480, 230, 410, 399, 53, 548, 434, 542, 32, 343, 360, 382, 114, 525, 398, 551, 86, 568, 40, 66, 192, 416, 375, 411, 564, 409, 507, 463, 159, 337, 57, 266, 396, 381, 424, 261, 368, 23, 522]. Model-Agnostic [244]. Model-Based [477, 368]. Model-Strengthened [223]. Modeling [282, 162, 359, 227, 566, 527, 557, 103, 285, 467, 240, 203, 365, 260, 155, 483, 147, 8, 75, 188, 353, 210, 271, 358, 104, 170, 515, 311, 197, 315, 202, 270, 379, 80, 194, 109, 136, 100, 427, 528, 348, 334, 278, 373, 236, 91, 58, 85, 171, 384, 442, 126, 473, 101, 110, 435, 307, 309, 226]. Modelling [96, 19, 417, 520, 35, 297, 276, 565, 106, 501, 287, 518, 262, 539, 510, 414, 181, 251, 273, 168, 205, 137, 212, 277, 469, 332]. Models [431, 474, 506, 557, 269, 329, 532, 512, 403, 128, 451, 553, 229, 466, 78, 533, 536, 22, 241, 243, 571, 149, 133, 264, 440, 178, 24, 272, 456, 560, 140, 320, 134, 177, 517, 514, 554, 154, 71, 39, 281, 253, 254, 301, 247, 174, 235, 460, 433, 578, 393, 487, 113, 458, 319, 294, 257]. Modulation [1, 81]. Modules [133]. Moffitt [93]. Molecular [27, 143, 4]. Momentary [427]. Monocarpic [148]. Monocytopenia [106]. Monolayer [70, 258]. Morphoelastic [460]. Morphogenesis [287, 312]. Mortality [386, 54, 438, 263]. Mosaic [242]. Mosquito [555, 213, 397, 310]. Mosquito-Borne [310]. Mosquitoes [371, 310]. Most [12]. Motif [443]. Motility [492]. Motion [562, 327, 295]. Motor [164]. Motors [561, 143]. Mountain [530, 220, 409]. Mouse [571]. Movement [423, 404, 370, 253, 348, 150, 3]. Moving [131, 369, 363, 150]. Moving-Boundary [150]. Multi [321, 538, 148, 115, 289, 267, 308, 255, 154, 488, 382, 159]. Multi-centrosomal [308]. Multi-criterial [488]. Multi-Cue [321]. Multi-flowered [148]. Multi-objective [538]. Multi-Patch [154, 267]. Multi-scale [382]. Multi-species [115, 255, 159]. Multi-strain [289]. Multicellular [533, 40]. Multidimensional [557, 568]. Multigenerational [542]. Multilayer [6]. Multilevel [68, 402, 385]. Multiphase [459, 363]. Multiple [123, 268, 432, 12, 382, 84, 497, 266]. Multiresolution [282]. Multiscale [357, 260, 523, 150, 277, 375]. Multistationarity [341]. Multivariate [431]. Murine [209]. Murray [181, 292]. Muscle [97, 362, 410, 399]. Mussel [53, 254]. Mutant [503]. Mutated [223]. Mutation [369, 578, 294]. Mutations [9, 407]. Mutually [48]. Myeloproliferative [578]. Mysteries [497].

Nagumo [421, 243]. Narrow [335]. Narrow-Spectrum [335]. Nascent [158]. Native [206]. Natural [470, 230]. Nature [129, 16]. Needs [62]. Negative [566, 429]. Neoplasm [578]. Nerve [421]. Network [321, 232, 346, 293, 173, 133, 264, 440, 552, 405, 283, 80, 67, 48]. Network-Based [283, 80]. Networks [98, 509, 426, 361, 461, 556, 176, 341, 470, 502, 72, 365, 468, 6, 231, 355, 395, 462, 130, 118, 246, 253, 568, 4, 433, 401, 305, 352, 511]. Neural [502]. Neurobiology [513]. Neurons [531]. Neuroprotectant [227]. Neutral [142, 175, 407]. Neutropenia [106]. Next [11]. Next-Generation [11]. NF [551]. Niche [434]. NK [566]. Node [418]. Nodes [209]. Noise [46, 383, 336, 208, 267, 57]. Noise-Induced [208, 57]. Noises [481]. Non [223, 321, 317, 54, 33, 149, 115, 430, 120, 114, 416, 333, 487, 389, 407]. Non-adherence [487]. Non-conducting [149]. Non-genetic [430]. Non-homogeneous [317]. Non-instantaneous [389]. Non-Local [321, 115, 114, 416]. Non-neutral [407]. Non-pharmaceutical [54, 120, 333]. Non-small-cell [223]. Nonadherence [299]. Nonequilibrium [27]. Nonlinear [112, 532, 544, 95, 457]. Nonlocal [313, 466]. Nonparametric [148]. Nonreflecting [79]. Nonsmooth [57].

Normal [305]. Note [39]. Novel [550]. Novo [277]. NPIs [54]. Nucleus [510, 536]. Number [530, 317, 516, 244, 116, 350, 439, 372]. Numbers [233, 113]. Numerical [97, 115, 134, 444, 389]. Nutrient [483].

Obesity [139, 166]. objective [538]. Oblique [5]. Observation [39]. Observations [519]. Obstacles [131]. Oceanic [481]. ODE [532, 174]. Off [438]. Olfactory [383]. Oncology [93]. Oncolytic [105, 340, 514, 230]. One [260, 393]. One-Compartmental [393]. One-Dimensional [260]. Only [90]. Onset [464, 176, 444]. Open [286, 237]. Opinion [458]. Opioid [467]. Opportunities [63, 212]. Optima [12]. Optimal [465, 141, 179, 201, 453, 14, 437, 120, 398, 521, 575, 163, 18]. Optimality [495]. Optimization [340, 95, 545]. Optimizing [547, 77]. Optimum [369]. Orchard [556, 352]. Order [572, 578]. Order-of-Mutation [578]. Ordinary [469]. Organisation [419]. Organism [348]. Organs [209]. Orientation [501]. Origin [385]. Oscillations [361, 59, 217, 265, 544, 192, 216, 274]. Oscillatory [276, 215, 324]. Other [259]. Our [61]. Outbreak [431, 412, 141, 320]. Outbreaks [530, 491, 75, 220, 104]. Outcomes [565, 315]. Outperform [28]. Overcoming [10]. Overcompensatory [296]. Overfishing [325]. Overloading [136]. Overt [222]. Oviparous [91]. Oviposition [409]. Own [569]. Oxygen [472, 83]. Oxygenation [472].

P. [465]. Pacemaker [97, 362]. Pairs [516]. Pairwise [232]. Pancreatic [105, 270]. Pandemic [364, 120]. Pandemics [212]. Papillomavirus [160]. Paradox [415]. Parameter [557, 558, 486, 470, 256, 508, 440, 73, 211, 446, 484, 74, 202, 400, 318]. Parameters [532, 178]. Parametric [152]. Parasite [450]. Parasitoid [424]. Parotid [40]. ParS [557]. Parsimonious [12]. Part [69, 92]. Partial [36, 535]. Partially [28]. Particle [85]. Particles [131]. Partitioning [338]. Passive [26, 146]. Pasture [526]. Patch [151, 133, 560, 154, 264, 267]. Patches [331, 425, 204, 200]. Patchy [214, 3]. Pathogen [489, 37]. Pathogens [56]. Paths [329]. Pathway [35, 297]. Pathways [193]. Patient [565, 197, 205]. Patient-Specific [205]. Patients [223, 456, 480]. Pattern [131, 274, 265, 559, 362, 550, 503, 6, 146, 567, 207, 549, 53, 235, 416]. Patterning [138]. Patterns [276, 148, 508, 188, 237, 325, 253, 114, 86]. Pausing [158]. Paying [61]. PBPK [558, 278]. PCR [505]. PDE [385, 256]. Peaks [350, 439]. PEGylated [278]. Perceptions [141]. Percolation [380]. Performance [386, 441]. Period [167]. Periodic [298, 423, 31]. Periods [96]. Peritumoural [67]. Permanently [425]. Persistence [59, 411]. Perspective [128, 141, 25, 494, 272, 488, 318]. Perspectives [539]. Perturbation [286, 328]. Perturbations [390, 470, 159]. Pest [112]. Pest/Vector [112]. Pesticides [213]. Petal [542]. Phage [330, 77]. Phagocytosis [251]. pharmaceutical [54, 120, 333]. Pharmacokinetic [393, 487]. Pharmacology [318]. Phase [167, 404, 506, 466, 394]. Phase-Field [466]. Phase-Type [394]. Phenomenon [457]. Phenotype [511]. Phenotypes [542]. Phenotypic [303, 419, 17, 473]. Philippines [478]. Photoreceptors [575]. Photosynthesis [5]. Phylogenetic [72, 468, 462, 500, 441, 401, 305]. Phylogenetics [153]. Phylogenies [394]. Phylosymmetric [153]. Physiologically [257, 241]. physiology [149]. PIDE [568]. Pine [530, 220, 409]. Placing [538]. Planktos [348]. Plant [96, 19, 59, 201, 320, 255, 109]. Plant-Vector-Virus [320]. Plaque [459, 251, 507]. Plasmodium [161, 357, 34]. Plasticity [303]. Platelets [85]. Plausible [546]. Player [302]. Plus [336]. Pod [256]. Point [208]. Points [47, 5]. Polar [390]. Polarity [184]. Polarization [458]. Policies [437, 384, 333]. Policy [373, 398, 366]. Poliovirus [338]. Pollinator [59]. Polymerization [202]. Polymorphic [33]. Polyploidy [401]. Population [530, 317, 76, 160, 336, 257, 241, 370, 115, 397, 438, 114, 338, 411, 409, 3, 216]. Populations [83, 28, 220, 115, 300, 430, 425, 84]. Pore [447]. Porosity [20]. Porous [260]. Positioning [262]. Possible [401]. Post [553, 384]. Post-vaccination [384]. Potassium [215, 324, 185]. Potency [285]. Potential [504, 392, 332]. Power [21]. Practical [304]. Practice [191]. Pre [87, 434]. Pre-exposure [87]. Pre-Metastatic [434]. Precipitation [508]. Precipitation-Dependent [508]. Preconditioning [197]. predation [214]. Predator [423, 274, 265, 55, 391, 343, 200, 481]. Predators [387]. Predict [406]. Predicting [513, 565, 209, 36, 6, 56, 144]. Prediction [223, 27, 366]. Predictive [502, 548]. Predicts [530]. Preferential [251]. PrEP [463]. Presence [458]. Preservice [65]. Pretty [517]. Prevalence [209, 87, 43]. Prevention [307]. Prey [423, 387, 274, 265, 214, 55, 391, 343, 200]. Prey-Induced [387]. Primary [40]. Priming [475]. Principle [55]. Private [524]. Probabilistic [540, 218]. Probabilities [249]. Probability [403, 320, 516, 338]. Problem [135]. Problems [95, 453]. Process [516, 564, 448, 428, 294]. Processes [376, 345, 39]. Produce [305]. Produced [505, 327, 295]. Producer [57]. Production [254]. Professional [61]. Program [93]. Progression [349, 578]. Progressive [136]. Projection [368]. Prokaryotes [541]. Proliferation [447, 430]. Promote [175, 325]. Promotes [308, 444]. Promotion [402]. Proof [118]. Propagation [426, 157, 502, 130]. propelled [131]. Properties [269, 183, 486, 502, 153, 552, 500]. Prophylaxis [87]. Prostate [147, 281, 545]. Protective [180]. Protein [1, 81, 164, 32, 155]. Protocell [385]. Provides [525]. Province [75]. Prune [570]. Public [524, 141]. Pulses [378]. Pulsing [69, 92]. Pushing [140]. Pyramidal [531].

Qssa [504]. Qualitative [486, 479]. Quality [175]. Quantification [548]. Quantify [278]. Quantifying [505, 573]. Quantitative [167, 65, 129, 318, 478]. Quarantine [326, 355, 49, 218, 398, 309]. Quartet [490, 248]. Quartet-Based [490]. Quartets [99]. Quasi [286, 283]. Quasi-Stationary [283]. Quasi-Steady-State [286].

Rabies [171]. Radiation [469]. Radiotherapy [565, 496]. Radius [372].

Random [98, 347, 247]. Range [28, 508, 316, 280]. Ranked [570]. Rapid [535]. Rare [298]. Rate [45, 14, 469]. Rates [354, 407]. Rather [127]. RdCVFL [575]. Reaction [46, 221, 509, 461, 559, 470, 7, 550, 483, 231, 237, 436, 198, 543, 405, 118, 484, 246, 374, 53, 32, 113]. Reactions [286]. Reactor [165]. Reading [446]. Real [565, 403, 199]. Real-Time [565, 199]. Rearrangement [529]. Receding [190]. Receptor [524]. Reciprocal [350, 439]. Reciprocity [402]. Reconstructing [183, 135]. Recovery [354, 325]. Recruitment [296, 230]. Recurrence [145]. Red [172, 225]. Reduce [54]. Reduced [42, 243, 254]. Reducing [505]. Reduction [246]. Reefs [481]. Reentry [444]. Reform [111]. Reframing [453]. Refunding [335]. Regeneration [20]. Regimens [299]. Regional [7, 156, 325]. Regraft [570]. Regular [519]. Regulated [48]. Regulating [195]. Regulation [506, 365, 137]. Regulatory [293, 365, 525, 568, 48, 511]. Reinforcement [538]. Related [504, 364, 137]. Relation [107, 350, 439]. Relationship [240]. Relative [441]. Release [180]. Releasing [397]. Remodeling [136]. Remote [209]. Renewal [128, 124]. Rényi [346]. Reopening [384]. Reorientation [482]. Repair [13]. Replicate [406]. Replication [116]. Replicator [68, 144]. Reply [302]. Repolarization [390, 444]. Reproduces [542]. Reproduction [317, 244, 263, 113, 372]. Reproductive [378]. Republic [412, 351]. Rescorla [71]. Research [102, 353, 271, 117]. Resembling [336]. Resident [425]. Residual [112]. Resilience [268]. Resistance [464, 10, 191]. Resistant [373]. Resonance [31, 241, 257]. Resorption [136]. Resource [378, 396]. Resource-Based [396]. Resources [326, 331, 101]. Respiratory [290, 577, 522]. Response [417, 28, 429, 193, 1, 81, 214, 537, 234, 364, 37, 288, 195, 126]. Responses [496, 258, 16, 261]. Results [547]. Retinal [449]. Retrieval [446]. Retroactive [82]. Reveals [523]. Reversing [194]. Review [539, 181]. Revisited [52]. Reward [128]. Rham [110]. Rhesus [234]. Rhythm [456]. Ribosome [541]. Rigorous [486]. Ring [176]. Rise [76]. Risk [304, 52, 368, 104]. Risk-Spreading [304]. River [95, 133]. rna [135, 443, 158, 155, 536]. Role [388, 142, 61, 208, 392, 515, 255, 195, 425, 476, 253, 184, 67, 235]. Roles [504, 37]. Roots [323]. Ross [49]. Rotating [376]. Route [38]. Routes [266]. RQMC [246]. Rush [309].

Saliva [40]. Salivary [186, 360]. Salmonella [332]. Sampled [513]. Samples [121]. sampling [23]. SARS [331]. SARS-CoV-2 [331]. SAS [79]. Savanna [268]. Savannas [549]. Scaffold [447]. Scalable [499]. Scale [131, 523, 400, 67, 382]. Scales [328, 147]. Schedules [547]. Scheme [335]. School [93, 448, 428]. Schools [351]. Science [129]. Sciences [89]. Sclerosis [497]. Search [347, 270, 92]. Seasonal [571, 75, 154, 315, 337]. Seasonality [555, 182, 198, 422]. Secondary [240, 443]. Secretion [40]. Sedentary [28]. Seed [201]. Seeds [172, 225]. SEIR [98, 180, 156, 384]. Selection [402, 385, 558, 369, 328, 202]. Self [131, 355, 419]. Self-Distancing [355].

Self-Organisation [419]. Self-propelled [131]. Semi [561]. Semi-Markov [561]. Seminiferous [188]. Senegal [74]. Sensing [321]. Sensitive [293]. Sensitivity [306, 435]. Sensors [27]. Separation [286, 328]. Sequence [125]. Sequences [524, 23]. Sequencing [536]. Sequential [156]. Sequestration [341]. Sequestration-Transmutation [341]. Serial [386]. Series [431, 499]. Serotypes [546]. Severe [577]. Severity [240, 311]. Sex [160, 91]. Sexual [563]. Shape [300, 394, 525]. Sharp [190]. Sharp-Fronted [190]. Shear [85]. Shear-Injured [85]. Shed [537]. Shifting [282, 543]. Shifts [28, 280]. Shock [194]. Shoots [148]. Short [431, 209]. Short-Term [431, 209]. Should [28]. Shrink [313]. Sierra [104]. Sigmoidal [526, 393]. Sign [350, 439]. Signaling [30, 215, 324, 1, 81, 405, 189]. Signalling [19]. Signals [130]. Signature [220]. Silent [338]. Simple [342, 516, 301]. Simulating [258]. Simulation [449, 103, 134, 246, 444]. Simulations [535, 309]. Single [30, 536, 115]. Single- [115]. Single-Nucleus [536]. Singular [286, 470]. Sink [11, 204]. SIR [269, 346, 7, 425, 381]. SIS [232, 283]. Sites [432]. SIV [194]. Size [474, 30, 317]. Skeletal [410, 399]. Slow [274, 265, 340]. Small [348, 223]. Smooth [97, 362]. SMoRe [557]. Social [68, 229, 88, 212]. Societies [61]. Solid [356, 41, 415, 460]. Solutions [517]. Solve [95]. Somatic [86]. Some [547]. Source [11, 204]. Southeast [368]. Southern [409]. sp. [203]. Space [72, 572]. Spaces [557, 500]. SPADE4 [493]. Sparse [71]. Sparsely [513]. Sparsity [493]. Spatial [445, 187, 151, 387, 339, 256, 300, 140, 267, 514, 400, 114, 411, 60]. Spatially [362, 84, 171, 216]. Spatio [520, 390, 36, 345]. Spatio-Temporal [390, 345, 520]. Spatiotemporal [484, 114, 60]. Special [292, 291]. Specialisation [221]. Speciation [394]. Specie [316]. Species [207, 248, 189, 91, 115, 255, 159]. Specific [205, 384, 349]. Specification [443]. Spectral [372]. Spectrum [335]. Speed [123, 45, 55, 280]. Speeds [344]. Spherical [418, 149]. Spheroid [70, 258, 199]. Spheroids [554, 460]. Spinal [79]. Spindle [308]. Spine [185]. Spiral [376, 421]. Split [344, 233]. Spot [542]. spp [249]. Spread [187, 151, 47, 339, 179, 256, 344, 220, 140, 411]. Spreading [426, 269, 55, 304]. Spring [309]. Stability [423, 41, 32, 189]. Stable [302, 268]. Stage [455, 411]. Stage-Structured [411]. Stages [249]. Stagnation [170]. Stalemate [528]. Staphyloccus [373]. State [426, 286, 5, 74]. States [268, 255, 574, 577]. Static [98]. Stationary [152, 283, 53]. Statistical [183, 482, 427, 487]. Statistically [23]. Statistics [158, 169]. Stay [28]. Steady [286, 5]. Steady-State [5]. Steer [224]. Stem [415, 334]. Step [167]. Sterile [112, 310]. Sterility [438]. Stewardship [373]. Stigma [226, 210]. Stimuli [288]. Stochastic [46, 509, 285, 461, 286, 179, 156, 158, 492, 503, 146, 320, 13, 391, 21, 154, 430, 246, 525, 66, 159, 380, 424]. Stochasticity [367, 208]. Stoichiometric [475, 109, 284, 57]. Stoichiometry [228]. Strain [144, 235, 289]. Strains [221, 162, 359, 420, 123]. Strand [13]. Strategies [227, 160, 414, 8, 289, 244, 384, 163, 337, 408, 435]. Strategy [68, 302, 124, 194, 381]. Stratified [138]. Streams [560]. Strengthened [223]. Streptococcus [203]. Stress [235]. Stressors [534]. Stretch [501, 482]. Stripe [58]. Stroke [227]. Strong [543]. Structural [361]. Structure [455, 176, 272]. Structured [2, 239, 450, 510, 257, 241, 245, 84, 411, 507, 463, 337, 435]. Structures [443, 135, 348, 377]. Study [162, 359, 167, 520, 388, 467, 390, 386, 486, 238, 561, 1, 81, 391, 74, 244, 379, 485, 444, 442, 401, 337, 389, 396, 368]. Subcellular [523]. Subflattenings [441]. Subinhibitory [407]. Subset [558, 202]. Substrate [563, 501]. Subtree [570]. Success [569]. Successive [344]. Sufficient [133]. Suggests [234, 448, 428]. Super [232]. Supplying [127]. Support [24]. Supporting [206]. Suppression [397, 545]. Surface [32]. Surfing [5]. Surgery [553]. Surrogate [557]. Survival [223]. Susceptible [104]. Suspended [165]. Sustainable [262]. Swarm [224]. swimmer [178]. Swimming [569]. Switch [45]. Switches [189]. Switching [128]. Symbiotic [203]. Synchronized [216]. Syncytial [290]. Syndrome [577, 444]. Synergy [330]. Synthesis [451]. Synthetic [80]. Syphilis [442]. System [35, 265, 486, 7, 203, 268, 271, 55, 267, 364, 371, 380, 297, 274]. Systematic [295, 327]. Systemic [111]. Systems [46, 513, 378, 59, 217, 128, 559, 498, 95, 233, 237, 436, 374, 318, 91, 196, 377, 204, 200, 533].

T [417, 197, 379, 334]. T-Cell [417, 197]. Tackling [303]. Tailfins [58]. Taking [180]. Target [37]. Targeting [258]. Targets [270]. Tau [246]. Tau-Leaping [246]. Taxa [233, 516]. taxon [23]. tBregs [388]. Teach [181]. Teachers [65, 100]. Teaching [322]. Tear [73, 211]. Technique [112, 512]. techniques [511]. Tell [5]. Telomere [152]. Temperate [330]. Temperature [91, 368]. Temperature-Dependent [91]. Temporal [540, 390, 36, 345, 503, 520]. Temporary [546]. Tennessee [252]. Tension [30]. Tenth [412]. Term [431, 209, 469, 408]. Termination [158]. Tessera [50]. Test [519, 381]. Test-and-Trace [381]. Testing [342, 488]. TGF [334]. TGF- [334]. Their [5, 300]. Theorems [461]. Theoretic [141, 77, 118, 551]. Theoretical [82, 5]. Theory [494, 446, 92]. Therapeutic [77]. Therapy [303, 31, 340, 147, 197, 281, 545, 334, 294]. There [69]. Thermodynamic [78, 165, 136]. Thermotherapy [250]. Thinning [73, 211]. Third [373]. Third-Generation [373]. Three [449, 264, 207]. Three-patch [264]. Three-Species [207]. Threshold [180, 76, 182, 74]. Throughout [13]. Tick [571, 452, 216]. Tick-Borne [452]. Tilapia [245]. Time [242, 431, 565, 328, 354, 147, 572, 199, 499, 477, 121, 374, 400, 52, 48, 411]. Time-Delayed [48]. Time-Distributed [354]. Time-Scale [400]. Time-Series [431]. Times [516, 394]. Timescale [286]. Timescales [251]. Timing [176, 77, 16]. Tipping [208, 267]. Tissue [472, 447, 363, 523, 67]. Tissue-Engineering [447]. Tissue-Scale [523, 67]. Tissues [19, 132, 247, 235]. TKF91 [125, 23]. Tolerance [438]. Tool [153]. tools [294]. Topological [176]. Topology [133, 264, 413]. Torque [308]. Trace [381]. Trade [438]. Trade-Off [438]. Tradeoffs [18]. Training [548]. Trait [316]. Transacylation [504]. Transcription [479]. Transcriptional [158].

Transgressive [542]. Transient [530, 574]. Transients [378, 84]. Transit [572]. Transition [303, 5, 255]. Transitional [508]. Transitions [325, 57]. Translation [479]. Transmission [357, 465, 474, 412, 339, 414, 573, 370, 331, 245, 198, 577, 392, 349, 371, 382, 171, 60, 163, 448, 428, 266, 101, 435, 309]. Transmutation [341]. Transport [19, 128, 451, 365, 572, 146, 523, 164, 186, 247, 389, 319]. Transporter [64, 94]. Travel [309]. Traveling [123, 457, 115, 543]. Travelling [423, 190, 363]. Travelling-Wave [363]. Treated [223, 334]. Treatment [527, 340, 258, 547, 358, 433, 469]. Tree [556, 72, 268, 516, 248, 401]. Tree-Based [72]. Tree-Child [401]. Trees [269, 99, 352]. Trending [88]. Trends [282]. Tributes [291]. Trigger [361]. Triple [566, 429]. Triple-Negative [566, 429]. Triplet [248]. Tryptophan [451]. Tsetse [249]. Tuberculosis [349, 80, 478, 337]. Tubules [188]. Tug [564]. Tug-of-War [564]. Tumbling [389]. Tumor [303, 466, 41, 170, 515, 415, 528, 334, 473]. Tumor-Immune [515, 528, 473]. Tumors [340, 415, 473]. Tumour [496, 25, 22, 199, 554, 150, 460, 417]. Tumour-Induced [25]. Tumours [356]. Turing [376, 559, 550, 146, 138, 567, 542, 196, 377]. Tutorial [460]. Two [167, 151, 63, 68, 386, 486, 215, 324, 344, 331, 133, 214, 560, 288, 425, 109, 48, 200, 487, 380]. Two-Compartment [487]. Two-Dimensional [215, 324, 380]. Two-Patch [133]. Two-Patches [425]. Two-Patchy [214]. Two-Step [167]. Two-Strategy [68]. Two-Variable [486]. Type [466, 394, 564]. Tyrosine [223].

UCLA [322]. UK [191]. Ultrasound [250]. Ultrasound-Induced [250]. Uncertainty [505, 573, 548]. Uncovering [176]. Undergoing [456]. Undergraduate [62, 102, 353, 65]. Undergraduates [322]. Understand [491]. Understanding [566, 392]. Uniform [250]. United [577]. Units [8]. Univariate [431]. Unlocking [497]. Unreported [75]. Unstable [426]. Uranium [260]. USA [442]. Usage [201, 446]. Use [8, 427]. User [302]. Uses [401]. Using [465, 557, 532, 565, 576, 36, 414, 80, 69, 92, 91, 4, 384, 328, 484, 373].

Vaccination [454, 82, 160, 14, 577, 124, 315, 437, 495, 384, 366, 448, 428, 309]. Vaccine [162, 359]. Validated [475, 463]. Validity [286]. Values [486]. Variability [108, 154]. Variable [486]. Variance [246, 316]. Variants [577]. Variation [571, 438]. Variational [312]. Varying [264, 37]. Vasodilatory [130]. Vector [96, 242, 112, 306, 320, 15, 400, 301]. Vector-Borne [96, 400, 301]. Vector-Host [306]. Vectors [301]. Vegetation [508]. Veillonella [203]. Ventricle [79]. Verification [558]. Vertical [245, 471, 569]. Vesicles [143]. Via [263, 112, 385, 488, 171, 463, 99]. Viral [96, 420, 56, 37]. Virotherapy [105, 230]. Virulence [450, 18]. Virus [31, 503, 290, 182, 245, 330, 198, 537, 104, 320, 87, 163]. Viruses [142, 514, 522]. Vision [92]. VisualPDE [535]. Vitro [285, 258]. Vitro-Calibrated [258]. vivax [161, 357, 465, 34]. Vivo [537]. Volatile [255]. Volcano [389]. Volterra [486, 560]. Volume [568].

Wagner [71]. Walk [247]. War [412, 564]. Warburg [275]. Warming [280]. Wasserstein [499]. Water [186]. Wave [351, 123, 502, 363, 457]. Waves [423, 123, 376, 190, 421, 492, 432, 115, 543, 382]. Wavetrains [421]. Way [301]. Weak [532, 328, 569]. Weak-Form [532]. Weed [168]. Weighted [313, 248, 99]. Weights [490]. WENDy [532]. Which [155]. Who [160, 373, 337]. Within [509, 538, 42, 503, 67, 438, 236, 448, 428]. Within-Host [236, 428]. Within-Population [438]. without [55]. Wolbachia [213, 397, 371]. Wolbachia-Infected [397]. Workplace [573]. Wound [518, 25, 202]. Wounds [139, 166]. Wuhan [326, 307].

Xeniid [69, 92]. Xylella [489, 187].

Year [93]. Years [63]. Yeast [544]. Yield [508]. Yule [516].

Zambia [477]. Zebrafish [58]. Zika [182, 198, 15, 266, 101, 368]. Zombie [177]. Zonal [327, 295].

References

Leander:2014:MCR

[1] R. Leander and A. Friedman. Modulation of the cAMP response by Gα_{i} and Gβγ: a computational study of G protein signaling in immune cells. Bulletin of Mathematical Biology, 76(6):1352–1375, June 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

     link.springer.com/article/10.1007/s11538-014-9964-4;   http:
     //link.springer.com/content/pdf/10.1007/s11538-014-9964-4.pdf.

See correction [81].

Callahan:2019:ALS

[2] Jason Callahan, Eric Eager, Richard Rebarber, Eva Strawbridge, and Shenglan Yuan. Analysis of a length-structured density-dependent model for fish. Bulletin of Mathematical Biology, 81(10):3732–3753, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

     tronic). URL http://link.springer.com/article/10.1007/s11538-
     019-00648-3;   http://link.springer.com/content/pdf/10.1007/
     s11538-019-00648-3.pdf. See correction [239].

Zaker:2020:EMB

[3] Nazanin Zaker, Laurence Ketchemen, and Frithjof Lutscher. The effect of movement behavior on population density in patchy landscapes.

Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-019-00680-3;
 http://link.springer.com/content/pdf/10.1007/s11538-019-00680-
3.pdf.

Vieira:2020:CIM

[4] Luis Sordo Vieira, Reinhard C. Laubenbacher, and David Murrugarra. Control of intracellular molecular networks using algebraic methods. Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00679-w;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00679-w.pdf.

Miranda-Apodaca:2020:SHE

[5] Jon Miranda-Apodaca, Emilio L. Marcos-Barbero, Rosa Morcuende, and Juan B. Arellano. Surfing the hyperbola equations of the steady-state Farquhar–von Caemmerer–Berry C_{3} leaf photosynthesis model: What can a theoretical analysis of their oblique asymptotes and transition points tell us? Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00676-z;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00676-z.pdf.

Hayes:2020:PPF

[6] Sean M. Hayes and Kurt E. Anderson. Predicting pattern formation in multilayer networks. Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00682-1;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00682-1.pdf.

ElAlamiLaaroussi:2020:RCR

[7] Adil El Alami Laaroussi and Mostafa Rachik. On the regional control of a reaction–diffusion system SIR. Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-019-00673-2;   http://link.springer.com/content/
pdf/10.1007/s11538-019-00673-2.pdf.

Huo:2020:MAU

[8] Xi Huo. Modeling antibiotic use strategies in intensive care units: Comparing de-escalation and continuation. Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-019-00686-x;   http://link.springer.com/content/
pdf/10.1007/s11538-019-00686-x.pdf.

Marculis:2020:IDI

[9] Nathan G. Marculis and Mark A. Lewis. Inside dynamics of integrodifference equations with mutations. Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-019-00683-0;   http://link.springer.com/content/
pdf/10.1007/s11538-019-00683-0.pdf.

Friedman:2020:ODR

[10] Avner Friedman and Nourridine Siewe. Overcoming drug resistance to BRAF inhibitor. Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00691-0;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00691-0.pdf.

Harrington:2020:NGA

[11] Peter D. Harrington and Mark A. Lewis. A next-generation approach to calculate source–sink dynamics in marine metapopulations. Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-019-00674-1;
 http://link.springer.com/content/pdf/10.1007/s11538-019-00674-
1.pdf.

Matsieva:2020:MPL

[12] Julia Matsieva and Katherine St.John. Most parsimonious likelihood exhibits multiple optima for compatible characters. Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-019-00689-8;
 http://link.springer.com/content/pdf/10.1007/s11538-019-00689-
8.pdf.

Mohseni-Salehi:2020:SMD

[13] Fazeleh S. Mohseni-Salehi, Fatemeh Zare-Mirakabad, Mehdi Sadeghi, and Soudeh Ghafouri-Fard. A stochastic model of DNA double-strand breaks repair throughout the cell cycle. Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-019-00692-z;   http://link.springer.com/content/
pdf/10.1007/s11538-019-00692-z.pdf.

Maier:2020:OAV

[14] Sandra B. Maier, Eduardo Massad, Marcos Amaku, Marcelo N. Burattini, and David Greenhalgh. The optimal age of vaccination against dengue with an age-dependent biting rate with application to Brazil. Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00690-1;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00690-1.pdf.

Olawoyin:2020:CAV

[15] Omomayowa Olawoyin and Christopher Kribs. Coinfection, altered vector infectivity, and antibody-dependent enhancement: The dengue– Zika interplay. Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00681-2;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00681-2.pdf.

Tyson:2020:TNB

[16] Rebecca C. Tyson, Stephanie D. Hamilton, Aboubakr S. Lo, Bert O. Baumgaertner, and Stephen M. Krone. The timing and nature of behavioural responses affect the course of an epidemic. Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-019-00684-z;
 http://link.springer.com/content/pdf/10.1007/s11538-019-00684-
z.pdf.

Strobl:2020:MMP

[17] Maximilian A. R. Strobl, Andrew L. Krause, Mehdi Damaghi, Robert Gillies, Alexander R. A. Anderson, and Philip K. Maini. Mix and match: Phenotypic coexistence as a key facilitator of cancer invasion. Bulletin of Mathematical Biology, 82(1):??, January 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00675-0;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00675-0.pdf.

daSilva:2020:OVD

[18] Esdras Jafet Aristides da Silva and César Castilho. Optimal virulence, diffusion and tradeoffs. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00688-9;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00688-9.pdf.

Allen:2020:MMA

[19] Henry R. Allen and Mariya Ptashnyk. Mathematical modelling of auxin transport in plant tissues: Flux meets signalling and growth. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-019-00685-y;
 http://link.springer.com/content/pdf/10.1007/s11538-019-00685-
y.pdf.

Cassani:2020:HMC

[20] Simone Cassani and Sarah D. Olson. A hybrid model of cartilage regeneration capturing the interactions between cellular dynamics and porosity. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00695-1;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00695-1.pdf.

Naasell:2020:ACS

[21] Ingemar Nåsell. Approximations of cumulants of the stochastic power law logistic model. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00687-w;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00687-w.pdf.

Grimes:2020:CEC

[22] David Robert Grimes and Alexander G. Fletcher. Close encounters of the cell kind: The impact of contact inhibition on tumour growth and cancer models. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00677-y;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00677-y.pdf.

Fan:2020:SCC

[23] Wai-Tong Fan and Sebastien Roch. Statistically consistent and computationally efficient inference of ancestral DNA sequences in the TKF91 model under dense taxon sampling. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00693-3;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00693-3.pdf.

Laudanno:2020:AAS

[24] Giovanni Laudanno, Bart Haegeman, and Rampal S. Etienne. Additional analytical support for a new method to compute the likelihood of diversification models. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00698-y;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00698-y.pdf.

Flegg:2020:CPW

[25] Jennifer A. Flegg, Shakti N. Menon, Helen M. Byrne, and D. L. Sean McElwain. A current perspective on wound healing and tumour-induced angiogenesis. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00696-0;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00696-0.pdf.

Adimy:2020:MPI

[26] Mostafa Adimy, Paulo F. A. Mancera, Diego S. Rodrigues, Fernando L. P. Santos, and Cláudia P. Ferreira. Maternal passive immunity and dengue hemorrhagic fever in infants. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00699-x;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00699-x.pdf.

Marzen:2020:PDN

[27] Sarah E. Marzen and James P. Crutchfield. Prediction and dissipation in nonequilibrium molecular sensors: Conditionally Markovian channels driven by memoryful environments. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00694-2;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00694-2.pdf.

Cobbold:2020:SSS

[28] Christina A. Cobbold and Remus Stana. Should I stay or should I go: Partially sedentary populations can outperform fully dispersing populations in response to climate-induced range shifts. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00700-7;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00700-7.pdf.

Jin:2020:HMB

[29] Xing Jin, Jeffrey S. Marshall, and Matthew J. Wargo. Hybrid model of bacterial biofilm growth. Bulletin of Mathematical Biology, 82(2): ??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00701-6;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00701-6.pdf.

Buttenschon:2020:CSM

[30] Andreas Buttenschön, Yue Liu, and Leah Edelstein-Keshet. Cell size, mechanical tension, and GTPase signaling in the single cell. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00702-5;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00702-
5.pdf.

Browne:2020:RPC

[31] Cameron J. Browne, Xuejun Pan, Hongying Shu, and Xiang-Sheng Wang. Resonance of periodic combination antiviral therapy and intracellular delays in virus model. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00704-3;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00704-3.pdf.

Stolerman:2020:SAB

[32] Lucas M. Stolerman, Michael Getz, Stefan G. Llewellyn Smith, Michael Holst, and Padmini Rangamani. Stability analysis of a bulk–surface reaction model for membrane protein clustering. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00703-4;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00703-
4.pdf.

Fishman:2020:PEG

[33] Michael A. Fishman. Polymorphic evolutionary games and non– Mendelian genetics. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00705-2;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00705-2.pdf.

Mehra:2020:ACM

[34] Somya Mehra, James M. McCaw, Mark B. Flegg, Peter G. Taylor, and Jennifer A. Flegg. An activation–clearance model for Plasmodium vivax malaria. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00706-1;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00706-1.pdf.

Bakshi:2020:MMA

[35] Suruchi Bakshi, Fraser Cunningham, and Piet H. van der Graaf. Mathematical modelling of alternative pathway of complement system. Bulletin of Mathematical Biology, 82(2):??, February 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/article/10.1007/s11538-020-00708-z. See correction

[297].

Dobreva:2020:TPS

[36] Atanaska Dobreva, Ralf Paus, and N. G. Cogan. Toward predicting the spatio–temporal dynamics of Alopecia Areata lesions using partial differential equation analysis. Bulletin of Mathematical Biology, 82(3):??, March 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00707-0;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00707-0.pdf.

Moore:2020:VID

[37] James R. Moore, Hasan Ahmed, Balaji Manicassamy, Adolfo Garcia-Sastre, Andreas Handel, and Rustom Antia. Varying inoculum dose to assess the roles of the immune response and target cell depletion by the pathogen in control of acute viral infections. Bulletin of Mathematical Biology, 82(3):??, March 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00711-4;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00711-4.pdf.

Dalwadi:2020:AAM

[38] Mohit P. Dalwadi and John R. King. An asymptotic analysis of the Malonyl–CoA route to 3-hydroxypropionic acid in genetically engineered microbes. Bulletin of Mathematical Biology, 82(3):??, March 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00714-1;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00714-1.pdf.

Park:2020:NOP

[39] Sang Woo Park and Benjamin M. Bolker. A note on observation processes in epidemic models. Bulletin of Mathematical Biology, 82(3):??, March 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00713-2;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00713-2.pdf.

Vera-Siguenza:2020:MMP

[40] Elías Vera-Sigüenza, Nathan Pages, John Rugis, David I. Yule, and James Sneyd. A multicellular model of primary saliva secretion in the parotid gland. Bulletin of Mathematical Biology, 82(3):??, March 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00712-3;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00712-3.pdf.

Lu:2020:CFF

[41] Min-Jhe Lu, Chun Liu, John Lowengrub, and Shuwang Li. Complex far-field geometries determine the stability of solid tumor growth with chemotaxis. Bulletin of Mathematical Biology, 82(3):??, March 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

020-00716-z;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00716-z.pdf.

Ford:2020:DRM

[42] Noah Ford and David Chopp. A dimensionally reduced model of biofilm growth within a flow cell. Bulletin of Mathematical Biology, 82(3):??, March 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00715-0;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00715-0.pdf.

Zahid:2020:DDI

[43] Mondal Hasan Zahid and Christopher M. Kribs. Decoys and dilution: The impact of incompetent hosts on prevalence of Chagas disease. Bulletin of Mathematical Biology, 82(3):??, March 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00710-5;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00710-5.pdf.

Aguilar:2020:EMM

[44] Jacob B. Aguilar and Juan B. Gutierrez. An epidemiological model of malaria accounting for asymptomatic carriers. Bulletin of Mathematical Biology, 82(3):??, March 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-020-00717-y;   http://link.springer.com/
content/pdf/10.1007/s11538-020-00717-y.pdf.

Curtin:2020:SSG

[45] Lee Curtin, Andrea Hawkins-Daarud, Kristoffer G. van der Zee, Kristin R. Swanson, and Markus R. Owen. Speed switch in glioblastoma growth rate due to enhanced hypoxia-induced migration. Bulletin of Mathematical Biology, 82(3):??, March 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00718-x;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00718-x.pdf.

Adamer:2020:CNS

[46] Michael F. Adamer, Heather A. Harrington, Eamonn A. Gaffney, and Thomas E. Woolley. Coloured noise from stochastic inflows in reaction– diffusion systems. Bulletin of Mathematical Biology, 82(4):??, April 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

020-00719-w;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00719-w.pdf.

Bayani:2020:MPC

[47] A. Bayani, J. L. Dunster, J. J. Crofts, and M. R. Nelson. Mechanisms and points of control in the spread of inflammation: A mathematical investigation. Bulletin of Mathematical Biology, 82(4):??, April 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00709-y;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00709-y.pdf.

Wang:2020:DAT

[48] Guiyuan Wang, Zhuoqin Yang, and Marc Turcotte. Dynamic analysis of the time-delayed genetic regulatory network between two auto-regulated and mutually inhibitory genes. Bulletin of Mathematical Biology, 82(4):??, April 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00722-1;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00722-1.pdf.

Jin:2020:MAR

[49] Xiulei Jin, Shuwan Jin, and Daozhou Gao. Mathematical analysis of the Ross–Macdonald model with quarantine. Bulletin of Mathematical Biology, 82(4):??, April 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00723-0;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00723-0.pdf.

Fimmel:2020:CTC

[50] Elena Fimmel, Martin Starman, and Lutz Strüngmann. Circular tessera codes in the evolution of the genetic code. Bulletin of Mathematical Biology, 82(4):??, April 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00724-z;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00724-z.pdf.

Malik:2020:IED

[51] A. A. Malik, B. Wennberg, and P. Gerlee. The impact of elastic deformations of the extracellular matrix on cell migration. Bulletin of Mathematical Biology, 82(4):??, April 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-020-00721-2;   http://link.springer.com/
content/pdf/10.1007/s11538-020-00721-2.pdf.

Tal:2020:ABH

[52] Omri Tal and Tat Dat Tran. Adaptive bet-hedging revisited: Considerations of risk and time horizon. Bulletin of Mathematical Biology, 82(4):??, April 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00729-8;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00729-8.pdf.

Shen:2020:SPR

[53] Zuolin Shen and Junjie Wei. Stationary pattern of a reaction–diffusion mussel–algae model. Bulletin of Mathematical Biology, 82(4):??, April 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00727-w;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00727-w.pdf.

Eubank:2020:CFA

[54] S. Eubank, I. Eckstrand, B. Lewis, S. Venkatramanan, M. Marathe, and C. L. Barrett. Commentary on Ferguson, et al., “Impact of Nonpharmaceutical Interventions (NPIs) to Reduce COVID-19 Mortality and Healthcare Demand”. Bulletin of Mathematical Biology, 82(4):??, April 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00726-x;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00726-x.pdf.

Lin:2020:SSI

[55] Guo Lin, Yibin Niu, Shuxia Pan, and Shigui Ruan. Spreading speed in an integrodifference predator–prey system without comparison principle. Bulletin of Mathematical Biology, 82(5):??, May 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00725-y;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00725-y.pdf.

Lange:2020:MFP

[56] Alexander Lange. A mathematical framework for predicting lifestyles of viral pathogens. Bulletin of Mathematical Biology, 82(5):??, May 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

020-00730-1;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00730-1.pdf.

Yuan:2020:NIT

[57] Sanling Yuan, Dongmei Wu, Guijie Lan, and Hao Wang. Noiseinduced transitions in a nonsmooth producer–grazer model with stoichiometric constraints. Bulletin of Mathematical Biology, 82(5):??, May 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00733-y;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00733-y.pdf.

Volkening:2020:MSF

[58] A. Volkening, M. R. Abbott, N. Chandra, B. Dubois, F. Lim, D. Sexton, and B. Sandstede. Modeling stripe formation on growing zebrafish tailfins. Bulletin of Mathematical Biology, 82(5):??, May 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00731-0;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00731-0.pdf.

Chen:2020:POP

[59] Mingshu Chen, Hong Wu, and Yuanshi Wang. Persistence and oscillations of plant–pollinator–herbivore systems. Bulletin of Mathematical Biology, 82(5):??, May 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00735-w;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00735-w.pdf.

Xiao:2020:CMM

[60] Yanni Xiao, Changcheng Xiang, Robert A. Cheke, and Sanyi Tang. Coupling the macroscale to the microscale in a spatiotemporal context to examine effects of spatial diffusion on disease transmission. Bulletin of Mathematical Biology, 82(5):??, May 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00736-9;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00736-
9.pdf.

Greer:2020:POD

[61] Meredith L. Greer, Olcay Akman, Timothy D. Comar, Daniel Hrozencik, and Jonathan E. Rubin. Paying our dues: The role of professional societies in the evolution of mathematical biology education. Bulletin of Mathematical Biology, 82(5):??, May 2020. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00728-9;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00728-
9.pdf.

Aikens:2020:MNC

[62] Melissa L. Aikens. Meeting the needs of a changing landscape: Advances and challenges in undergraduate biology education. Bulletin of Mathematical Biology, 82(5):??, May 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-020-00739-6;   http://link.springer.com/
content/pdf/10.1007/s11538-020-00739-6.pdf.

Bressoud:2020:OCF

[63] David M. Bressoud. Opportunities for change in the first two years of college mathematics. Bulletin of Mathematical Biology, 82(5):??, May 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00738-7;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00738-7.pdf.

Hiller:2020:MMK

[64] Rebecca M. Hiller, Julius von Kügelgen, Huan Bao, Franck Duong Van Hoa, and Eric N. Cytrynbaum. A mathematical model for the kinetics of the MalFGK_{2} maltose transporter. Bulletin of Mathematical Biology, 82(5):??, May 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00737-8;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00737-8.pdf. See correction [94].

Mayes:2020:UQB

[65] Robert Mayes, Tammy Long, Lacey Huffling, Aaron Reedy, and Brad Williamson. Undergraduate quantitative biology impact on biology preservice teachers. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00740-z;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00740-z.pdf.

Wang:2020:SMB

[66] Xueying Wang, Chadi M. Saad-Roy, and P. van den Driessche. Stochastic model of Bovine Babesiosis with juvenile and adult cattle. Bulletin of Mathematical Biology, 82(6):??, June 2020. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00734-x;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00734-
x.pdf.

Shuttleworth:2020:CSD

[67] Robyn Shuttleworth and Dumitru Trucu. Cell-scale degradation of peritumoural extracellular matrix fibre network and its role within tissuescale cancer invasion. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00732-z;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00732-z.pdf.

Cooney:2020:AMR

[68] Daniel B. Cooney. Analysis of multilevel replicator dynamics for general two-strategy social dilemma. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00742-x;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00742-x.pdf.

Samson:2020:CPXa

[69] Julia E. Samson and Laura A. Miller. Collective pulsing in Xeniid corals: Part II — using computational fluid dynamics to determine if there are benefits to coordinated pulsing. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00741-y;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00741-y.pdf.

Berrouet:2020:CDI

[70] Catherine Berrouet, Naika Dorilas, and Necibe Tuncer. Comparison of drug inhibitory effects (IC_{50}) in monolayer and spheroid cultures. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-020-00746-7.

Nishimura:2020:RWM

[71] Joel Nishimura and Amy L. Cochran. Rescorla–Wagner models with sparse dynamic attention. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
020-00743-w.

Fischer:2020:STB

[72] Mareike Fischer and Andrew Francis. The space of tree-based phylogenetic networks. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-020-00744-9.

Luke:2020:PEE

[73] Rayanne A. Luke, Richard J. Braun, and Deborah Awisi-Gyau. Parameter estimation for evaporation-driven tear film thinning. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-020-00745-8.

Ogola:2020:PSE

[74] Beda O. Ogola, Woldegebriel A. Woldegerima, and E. O. Omondi. Parameter and state estimation in a cholera model with threshold immunology: a case study of Senegal. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
020-00755-6.

Jing:2020:MEM

[75] Shuang-Lin Jing, Hai-Feng Huo, and Hong Xiang. Modeling the effects of meteorological factors and unreported cases on seasonal influenza outbreaks in Gansu Province, China. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-020-00747-6.

Fadai:2020:PDT

[76] Nabil T. Fadai and Matthew J. Simpson. Population dynamics with threshold effects give rise to a diverse family of Allee effects. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-020-00756-5.

Li:2020:OTC

[77] Guanlin Li, Chung Yin Leung, and Joshua S. Weitz. Optimizing the timing and composition of therapeutic phage cocktails: A control-theoretic approach. Bulletin of Mathematical Biology, 82(6):??, June 2020. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-020-00751-w.

Gaebler:2020:TIC

[78] Harry J. Gaebler and Hermann J. Eberl. Thermodynamic inhibition in chemostat models. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-020-00758-3.

Donatelli:2020:NBC

[79] Donatella Donatelli and Licia Romagnoli. Nonreflecting boundary conditions for a CSF model of fourth ventricle: Spinal SAS dynamics. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-020-00749-4.

Renardy:2020:FNB

[80] Marissa Renardy and Denise E. Kirschner. A framework for networkbased epidemiological modeling of tuberculosis dynamics using synthetic datasets. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-020-00752-9.

Leander:2020:CMC

[81] R. Leander and A. Friedman. Correction to: Modulation of the cAMP Response by Gα_{i} and Gβγ: a Computational Study of G Protein Signaling in Immune Cells. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
020-00753-8. See [1].

Chouhan:2020:GTM

[82] Ali Chouhan, Sohail Maiwand, and Dewey Taylor. Game-theoretical model of retroactive hepatitis B vaccination in China. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-020-00748-5.

Ardaseva:2020:MDA

[83] Aleksandra Ardaševa, Robert A. Gatenby, and Tommaso Lorenzi. A mathematical dissection of the adaptation of cell populations to fluctuating oxygen levels. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-020-00754-7.

Vortkamp:2020:MAL

[84] Irina Vortkamp, Sebastian J. Schreiber, and Frank M. Hilker. Multiple attractors and long transients in spatially structured populations with an Allee effect. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-020-00750-x.

Wang:2020:MCF

[85] Liwei Wang, Zengsheng Chen, Jiafeng Zhang, Xiwen Zhang, and Zhongjun J. Wu. Modeling clot formation of shear-injured platelets in flow by a dissipative particle dynamics method. Bulletin of Mathematical Biology, 82(7):??, July 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00760-9;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00760-9.pdf.

Utsey:2020:MMI

[86] Kiersten Utsey and James P. Keener. A mathematical model for inheritance of DNA methylation patterns in somatic cells. Bulletin of Mathematical Biology, 82(7):??, July 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-020-00765-4;   http://link.springer.com/
content/pdf/10.1007/s11538-020-00765-4.pdf.

Pharaon:2020:IPE

[87] Joe Pharaon and Chris T. Bauch. The impact of pre-exposure prophylaxis for human immunodeficiency virus on gonorrhea prevalence. Bulletin of Mathematical Biology, 82(7):??, July 2020. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00762-7;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00762-
7.pdf.

Sooknanan:2020:TSM

[88] J. Sooknanan and D. M. G. Comissiong. Trending on social media: Integrating social media into infectious disease dynamics. Bulletin of Mathematical Biology, 82(7):??, July 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00757-4;

 http://link.springer.com/content/pdf/10.1007/s11538-020-00757-
4.pdf.

Cozzens:2020:ICM

[89] Margaret Cozzens and Fred S. Roberts. Introductory College mathematics for the life sciences: Has anything changed? Bulletin of Mathematical Biology, 82(7):??, July 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00761-8;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00761-8.pdf.

Moulin:2020:MMG

[90] Thibault Moulin, Antoine Perasso, and Ezio Venturino. A metaecoepidemic model of grassland ecosystem with only consumers’ migration. Bulletin of Mathematical Biology, 82(7):??, July 2020. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00764-5;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00764-
5.pdf.

Verma:2020:MTD

[91] Nitu Verma, Babita K. Verma, and S. Pushpavanam. Modeling temperature-dependent sex determination in oviparous species using a dynamical systems approach. Bulletin of Mathematical Biology, 82(7):??, July 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00763-6;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00763-6.pdf.

Samson:2020:CPXb

[92] Julia E. Samson, Dylan D. Ray, Maurizio Porfiri, Laura A. Miller, and Simon Garnier. Collective pulsing in Xeniid corals: Part I — using computer vision and information theory to search for coordination. Bulletin of Mathematical Biology, 82(7):??, July 2020. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00759-2;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00759-
2.pdf.

Enderling:2020:HSI

[93] Heiko Enderling, Philipp M. Altrock, Noemi Andor, David Basanta, Joel S. Brown, Robert A. Gatenby, Andriy Marusyk, Katarzyna A. Rejniak, Ariosto Silva, and Alexander R. A. Anderson. High School Internship Program in Integrated Mathematical Oncology (HIP IMO): Fiveyear experience at Moffitt Cancer Center. Bulletin of Mathematical Biology, 82(7):??, July 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00768-1;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00768-1.pdf.

Hiller:2020:CMM

[94] Rebecca M. Hiller, Julius von Kügelgen, Huan Bao, Franck Duong Van Hoa, and Eric N. Cytrynbaum. Correction to: A Mathematical Model for the Kinetics of the MalFGK_{2} Maltose Transporter. Bulletin of Mathematical Biology, 82(7):??, July 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00772-5;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00772-

5.pdf. See [64].

Filippov:2020:LRE

[95] Alexander Filippov, Alexander Kovalev, and Stanislav Gorb. Large river effect or frozen kinetics: How complex nonlinear living systems solve optimization problems. Bulletin of Mathematical Biology, 82(7):??, July 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00766-3;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00766-3.pdf.

AlBasir:2020:MEI

[96] Fahad Al Basir, Sagar Adhurya, Malay Banerjee, Ezio Venturino, and Santanu Ray. Modelling the effect of incubation and latent periods on the dynamics of vector-borne plant viral diseases. Bulletin of Mathematical Biology, 82(7):??, July 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00767-2;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00767-2.pdf.

Fatoyinbo:2020:NBA

[97] H. O. Fatoyinbo, R. G. Brown, D. J. W. Simpson, and B. van Brunt. Numerical bifurcation analysis of pacemaker dynamics in a model of smooth muscle cells. Bulletin of Mathematical Biology, 82(7):??, July 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00771-6;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00771-6.pdf.

Alota:2020:EBM

[98] Cherrylyn P. Alota, Carlene P. C. Pilar-Arceo, and Aurelio A. de los Reyes V. An edge-based model of SEIR epidemics on static random networks. Bulletin of Mathematical Biology, 82(7):??, July 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00769-0;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00769-0.pdf.

Yourdkhani:2020:IMT

[99] Samaneh Yourdkhani and John A. Rhodes. Inferring metric trees from weighted quartets via an intertaxon distance. Bulletin of Mathematical Biology, 82(7):??, July 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00773-4;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00773-4.pdf.

Seshaiyer:2020:CTT

[100] Padmanabhan Seshaiyer and Suzanne Lenhart. Connecting with teachers through modeling in mathematical biology. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00774-3;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00774-3.pdf.

Zhao:2020:MDA

[101] Hongyong Zhao, Liping Wang, Sergio Muniz Oliva, and Huaiping Zhu. Modeling and dynamics analysis of Zika transmission with limited medical resources. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00776-1;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00776-1.pdf.

Bendinskas:2020:CUR

[102] Kestutis G. Bendinskas, Lester Caudill, and Luis A. Melara, Jr. The case for undergraduate research journals. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-020-00775-2.

Bodine:2020:ABM

[103] Erin N. Bodine, Robert M. Panoff, and Anton E. Weisstein. Agent-based modeling and simulation in mathematics and biology education. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-020-00778-z.

Lin:2020:MEV

[104] Qianying Lin, Salihu S. Musa, and Daihai He. Modeling the 2014–2015 Ebola virus disease outbreaks in Sierra Leone, Guinea, and Liberia with effect of high- and low-risk susceptible individuals. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-020-00779-y.

Chen:2020:CAM

[105] J. Chen, D. Weihs, and F. J. Vermolen. A cellular automata model of oncolytic virotherapy in pancreatic cancer. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-020-00780-5.

Cassidy:2020:CCI

[106] Tyler Cassidy, Antony R. Humphries, and Michael C. Mackey. Characterizing chemotherapy-induced neutropenia and monocytopenia through mathematical modelling. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
020-00777-0.

Fimmel:2020:RBC

[107] Elena Fimmel, Christian J. Michel, and Lutz Strüngmann. The relation between k-circularity and circularity of codes. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-020-00770-7.

Lehrer:2020:GGV

[108] Richard Lehrer, Leona Schauble, and Panchompoo Wisittanawat. Getting a grip on variability. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
020-00782-3.

Rong:2020:SMA

[109] Xinmiao Rong, Yazhou Sun, and Hao Wang. Stoichiometric modeling of aboveground–belowground interaction of herbaceous plant and two herbivores. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-020-00784-1.

Zhao:2020:RHA

[110] Rundong Zhao, Menglun Wang, and Guo-Wei Wei. The de Rham–Hodge analysis and modeling of biomolecules. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-020-00783-2.

Akman:2020:BCB

[111] Olcay Akman, Carrie Diaz Eaton, and Katerina V. Thompson. Building community-based approaches to systemic reform in mathematical biology education. Bulletin of Mathematical Biology, 82(8):??, August 2020. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-020-00781-4.

Aronna:2020:NPV

[112] M. Soledad Aronna and Yves Dumont. On nonlinear pest/vector control via the sterile insect technique: Impact of residual fertility. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-020-00790-3.

Yang:2020:BRN

[113] Chayu Yang and Jin Wang. Basic reproduction numbers for a class of reaction–diffusion epidemic models. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-020-00788-x.

Taylor:2020:NLC

[114] Nick P. Taylor, Hyunyeon Kim, and Robert A. Van Gorder. A non-local cross-diffusion model of population dynamics I: Emergent spatial and spatiotemporal patterns. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
020-00786-z.

Krause:2020:NLC

[115] Andrew L. Krause and Robert A. Van Gorder. A non-local cross-diffusion model of population dynamics II: Exact, approximate, and numerical traveling waves in single- and multi-species populations. Bulletin of Mathematical Biology, 82(8):??, August 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-020-00787-y.

Sadun:2020:ELE

[116] Lorenzo Sadun. Effects of latency on estimates of the COVID-19 replication number. Bulletin of Mathematical Biology, 82(9):??, September 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00791-2;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00791-2.pdf.

Macauley:2020:CAB

[117] Matthew Macauley and Nora Youngs. The case for algebraic biology: from research to education. Bulletin of Mathematical Biology, 82(9):??, September 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00789-w;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00789-w.pdf.

Muller:2020:DBC

[118] Stefan Müller and Badal Joshi. Detailed balance = complex balance + cycle balance: a graph-theoretic proof for reaction networks and Markov chains. Bulletin of Mathematical Biology, 82(9):??, September 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00792-1;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00792-1.pdf.

Jungck:2020:MBE

[119] John R. Jungck, Raina Robeva, and Louis J. Gross. Mathematical biology education: Changes, communities, connections, and challenges. Bulletin of Mathematical Biology, 82(9):??, September 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00793-0;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00793-0.pdf.

Perkins:2020:OCC

[120] T. Alex Perkins and Guido España. Optimal control of the COVID-19 pandemic with non-pharmaceutical interventions. Bulletin of Mathematical Biology, 82(9):??, September 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-020-00795-y;   http://link.springer.com/
content/pdf/10.1007/s11538-020-00795-y.pdf.

Nardini:2020:LEB

[121] John T. Nardini, John H. Lagergren, Andrea Hawkins-Daarud, Lee Curtin, Bethan Morris, Erica M. Rutter, Kristin R. Swanson, and Kevin B. Flores. Learning equations from biological data with limited time samples. Bulletin of Mathematical Biology, 82(9):??, September 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00794-z;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00794-z.pdf.

Lee:2020:MBE

[122] Shernita Lee and LaShania Clinedinst. Mathematical biology: Expand, expose, and educate! Bulletin of Mathematical Biology, 82(9):??, September 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00796-x;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00796-x.pdf.

Chen:2020:TWE

[123] Guoting Chen, Xinchu Fu, and Mengfeng Sun. Traveling waves and estimation of minimal wave speed for a diffusive influenza model with multiple strains. Bulletin of Mathematical Biology, 82(9):??, September 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00799-8;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00799-8.pdf.

Nzokem:2020:EDA

[124] Aubain Nzokem and Neal Madras. Epidemic dynamics and adaptive vaccination strategy: Renewal equation approach. Bulletin of Mathematical Biology, 82(9):??, September 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-020-00802-2;   http://link.springer.com/
content/pdf/10.1007/s11538-020-00802-2.pdf.

Fan:2020:ICD

[125] Wai-Tong Louis Fan, Brandon Legried, and Sebastien Roch. Impossibility of consistent distance estimation from sequence lengths under the TKF91 model. Bulletin of Mathematical Biology, 82(9):??, September 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00801-3;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00801-3.pdf.

Wyatt:2020:MEM

[126] Asia Wyatt and Doron Levy. Modeling the effect of memory in the adaptive immune response. Bulletin of Mathematical Biology, 82(9): ??, September 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00798-9;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00798-9.pdf.

Loo:2020:WMC

[127] Sara L. Loo, Danya Rose, Michael Weight, Kristen Hawkes, and Peter S. Kim. Why males compete rather than care, with an application to supplying collective goods. Bulletin of Mathematical Biology, 82(10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00800-4;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00800-4.pdf.

Ciocanel:2020:RRP

[128] Maria-Veronica Ciocanel, John Fricks, Peter R. Kramer, and Scott A. McKinley. Renewal reward perspective on linear switching diffusion systems in models of intracellular transport. Bulletin of Mathematical Biology, 82(10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00797-w;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00797-w.pdf.

Robeva:2020:CNQ

[129] Raina S. Robeva, John R. Jungck, and Louis J. Gross. Changing the nature of quantitative biology education: Data science as a driver. Bulletin of Mathematical Biology, 82(10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00785-0;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00785-0.pdf.

Lee:2020:EPV

[130] Pilhwa Lee. Electrical propagation of vasodilatory signals in capillary networks. Bulletin of Mathematical Biology, 82(10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00806-y;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00806-y.pdf.

Aceves-Sanchez:2020:LSD

[131] P. Aceves-Sanchez, P. Degond, E. E. Keaveny, A. Manhart, S. Merino-Aceituno, and D. Peurichard. Large-scale dynamics of self-propelled particles moving through obstacles: Model derivation and pattern formation. Bulletin of Mathematical Biology, 82(10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00805-z;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00805-z.pdf.

Murphy:2020:MCC

[132] Ryan J. Murphy, Pascal R. Buenzli, Ruth E. Baker, and Matthew J. Simpson. Mechanical cell competition in heterogeneous epithelial tissues. Bulletin of Mathematical Biology, 82(10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00807-x;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00807-x.pdf.

Jiang:2020:TPM

[133] Hongyan Jiang, King-Yeung Lam, and Yuan Lou. Are two-patch models sufficient? The evolution of dispersal and topology of river network modules. Bulletin of Mathematical Biology, 82(10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00803-1;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00803-1.pdf.

Mathias:2020:IFF

[134] Sonja Mathias, Adrien Coulier, Anass Bouchnita, and Andreas Hellander. Impact of force function formulations on the numerical simulation of centre-based models. Bulletin of Mathematical Biology, 82(10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

020-00810-2;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00810-2.pdf.

Greenwood:2020:PRM

[135] Torin Greenwood and Christine E. Heitsch. On the problem of reconstructing a mixture of rna structures. Bulletin of Mathematical Biology, 82(10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00804-0;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00804-0.pdf.

Sego:2020:MPD

[136] T. J. Sego, Yung-Ting Hsu, Tien-Min Chu, and Andres Tovar. Modeling progressive damage accumulation in bone remodeling explains the thermodynamic basis of bone resorption by overloading. Bulletin of Mathematical Biology, 82(10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00808-w;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00808-
w.pdf.

Scheepers:2020:CRA

[137] Ronél Scheepers, Graeme J. Pettet, Peter van Heijster, and Robyn P. Araujo. Cholesterol regulation in age-related macular degeneration: a framework for mathematical modelling of Drusen biogenesis. Bulletin of Mathematical Biology, 82(10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00812-0;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00812-
0.pdf.

Krause:2020:TPS

[138] Andrew L. Krause, Václav Klika, Jacob Halatek, Paul K. Grant, Thomas E. Woolley, Neil Dalchau, and Eamonn A. Gaffney. Turing patterning in stratified domains. Bulletin of Mathematical Biology, 82 (10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00809-9;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00809-9.pdf.

Friedman:2020:MMC

[139] Avner Friedman and Nourridine Siewe. Mathematical model of chronic dermal wounds in diabetes and obesity. Bulletin of Mathematical Biology, 82(10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00815-x;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00815-x.pdf. See correction [166].

Lutscher:2020:PBM

[140] Frithjof Lutscher, Justus Fink, and Yingjie Zhu. Pushing the boundaries: Models for the spatial spread of ecosystem engineers. Bulletin of Mathematical Biology, 82(10):??, October 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00818-8;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00818-
8.pdf.

Deka:2020:OMP

[141] Aniruddha Deka, Buddhi Pantha, and Samit Bhattacharyya. Optimal management of public perceptions during a flu outbreak: a game-theoretic perspective. Bulletin of Mathematical Biology, 82(11):??, November 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00817-9;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00817-9.pdf.

Bunimovich:2020:LIR

[142] Leonid Bunimovich and Longmei Shu. Local immunodeficiency: Role of neutral viruses. Bulletin of Mathematical Biology, 82(11):??, November 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00813-z;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00813-z.pdf.

Park:2020:DVD

[143] Youngmin Park and Thomas G. Fai. Dynamics of vesicles driven into closed constrictions by molecular motors. Bulletin of Mathematical Biology, 82(11):??, November 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-020-00820-0;   http://link.springer.com/
content/pdf/10.1007/s11538-020-00820-0.pdf.

Madec:2020:PSC

[144] Sten Madec and Erida Gjini. Predicting N-strain coexistence from cocolonization interactions: Epidemiology meets ecology and the replicator equation. Bulletin of Mathematical Biology, 82(11):??, November 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00816-w;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00816-w.pdf.

Curtin:2020:MII

[145] Lee Curtin, Andrea Hawkins-Daarud, Alyx B. Porter, Kristoffer G. van der Zee, Markus R. Owen, and Kristin R. Swanson. A mechanistic investigation into ischemia-driven distal recurrence of glioblastoma. Bulletin of Mathematical Biology, 82(11):??, November 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00814-y;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00814-y.pdf.

Kim:2020:STP

[146] Hyunjoong Kim and Paul C. Bressloff. Stochastic Turing pattern formation in a model with active and passive transport. Bulletin of Mathematical Biology, 82(11):??, November 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00822-y;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00822-
y.pdf.

Higgins:2020:TSA

[147] Raegan Higgins, Casey J. Mills, and Angela Peace. A time scales approach for modeling intermittent hormone therapy for prostate cancer. Bulletin of Mathematical Biology, 82(11):??, November 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00821-z;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00821-z.pdf.

Fomin:2020:NMA

[148] Eduard Fomin and Tatyana Fomina. A nonparametric model for analysis of flowering patterns of herbaceous multi-flowered monocarpic shoots. Bulletin of Mathematical Biology, 82(12):??, December 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00824-w;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00824-w.pdf.

Jiang:2020:EPM

[149] Jiamu Jiang, Paul Smith, and Mark C. W. van Rossum. Electrophysiology models of cells with spherical geometry with non-conducting center. Bulletin of Mathematical Biology, 82(12):??, December 2020. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00828-6;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00828-
6.pdf.

Suveges:2020:DMM

[150] Szabolcs Suveges, Raluca Eftimie, and Dumitru Trucu. Directionality of macrophages movement in tumour invasion: a multiscale movingboundary approach. Bulletin of Mathematical Biology, 82(12):??, December 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00819-7;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00819-7.pdf.

Arroyo-Esquivel:2020:SDS

[151] Jorge Arroyo-Esquivel and Alan Hastings. Spatial dynamics and spread of ecosystem engineers: Two patch analysis. Bulletin of Mathematical Biology, 82(12):??, December 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-020-00833-9;   http://link.springer.com/
content/pdf/10.1007/s11538-020-00833-9.pdf.

Lee:2020:SDT

[152] Kyung Hyun Lee and Marek Kimmel. Stationary distribution of telomere lengths in cells with telomere length maintenance and its parametric inference. Bulletin of Mathematical Biology, 82(12):??, December 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00811-1;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00811-1.pdf.

Hendriksen:2020:PAM

[153] Michael Hendriksen and Julia A. Shore. Phylosymmetric algebras: Mathematical properties of a new tool in phylogenetics. Bulletin of Mathematical Biology, 82(12):??, December 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00832-w;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00832-
w.pdf.

Nipa:2020:DEM

[154] Kaniz Fatema Nipa and Linda J. S. Allen. Disease emergence in multi-patch stochastic epidemic models with demographic and seasonal variability. Bulletin of Mathematical Biology, 82(12):??, December 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00831-x;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00831-x.pdf.

Gasior:2020:MMW

[155] K. Gasior, M. G. Forest, A. S. Gladfelter, and J. M. Newby. Modeling the mechanisms by which coexisting biomolecular RNA–Protein condensates form. Bulletin of Mathematical Biology, 82(12):??, December 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00823-x;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00823-x.pdf.

Engbert:2021:SDA

[156] Ralf Engbert, Maximilian M. Rabe, Reinhold Kliegl, and Sebastian Reich. Sequential data assimilation of the stochastic SEIR epidemic model for regional COVID-19 dynamics. Bulletin of Mathematical Biology, 83(1):??, January 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00834-8;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00834-8.pdf.

Berestycki:2021:PEA

[157] Henri Berestycki, Jean-Michel Roquejoffre, and Luca Rossi. Propagation of epidemics along lines with fast diffusion. Bulletin of Mathematical Biology, 83(1):??, January 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00826-8;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00826-8.pdf.

Filatova:2021:SNM

[158] Tatiana Filatova, Nikola Popovic, and Ramon Grima. Statistics of nascent and mature RNA fluctuations in a stochastic model of transcriptional initiation, elongation, pausing, and termination. Bulletin of Mathematical Biology, 83(1):??, January 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00827-7;

 http://link.springer.com/content/pdf/10.1007/s11538-020-00827-
7.pdf.

Xu:2021:CEG

[159] Chaoqun Xu, Sanling Yuan, and Tonghua Zhang. Competitive exclusion in a general multi-species chemostat model with stochastic perturbations. Bulletin of Mathematical Biology, 83(1):??, January 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00843-7;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00843-7.pdf.

Gao:2021:DMA

[160] Shasha Gao, Maia Martcheva, Hongyu Miao, and Libin Rong. A dynamic model to assess human papillomavirus vaccination strategies in a heterosexual population combined with men who have sex with men. Bulletin of Mathematical Biology, 83(1):??, January 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00830-y;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00830-y.pdf.

Mehra:2021:ADP

[161] Somya Mehra, James M. McCaw, Mark B. Flegg, Peter G. Taylor, and Jennifer A. Flegg. Antibody dynamics for plasmodium vivax malaria: a mathematical model. Bulletin of Mathematical Biology, 83(1):??, January 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00837-5;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00837-5.pdf.

Alharbi:2021:MMS

[162] Mohammed H. Alharbi and Christopher M. Kribs. A mathematical modeling study: Assessing impact of mismatch between influenza vaccine strains and circulating strains in Hajj. Bulletin of Mathematical Biology, 83(1):??, January 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00836-6;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00836-6.pdf. See correction [359].

Xue:2021:OCM

[163] Ling Xue, Xue Ren, Felicia Magpantay, Wei Sun, and Huaiping Zhu. Optimal control of mitigation strategies for dengue virus transmission. Bulletin of Mathematical Biology, 83(2):??, February 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00839-3;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00839-3.pdf.

Ryan:2021:MPT

[164] S. D. Ryan, Z. McCarthy, and M. Potomkin. Motor protein transport along inhomogeneous microtubules. Bulletin of Mathematical Biology, 83(2):??, February 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00838-4;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00838-4.pdf.

Gaebler:2021:TIB

[165] Harry J. Gaebler, Jack M. Hughes, and Hermann J. Eberl. Thermodynamic inhibition in a biofilm reactor with suspended bacteria. Bulletin of Mathematical Biology, 83(2):??, February 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00840-w;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00840-
w.pdf.

Friedman:2021:CMM

[166] Avner Friedman and Nourridine Siewe. Correction to: Mathematical Model of Chronic Dermal Wounds in Diabetes and Obesity. Bulletin of Mathematical Biology, 83(2):??, February 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00835-7;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00835-

7.pdf. See [139].

An:2021:TSM

[167] Zheming An, Nathaniel J. Merrill, Kwangwon Lee, Rémi Robin, Amaury Hayat, Olivia Zapfe, and Benedetto Piccoli. A two-step model of human entrainment: a quantitative study of circadian period and phase of entrainment. Bulletin of Mathematical Biology, 83(2):??, February 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00829-5;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00829-5.pdf.

Ramaj:2021:MMC

[168] Tedi Ramaj. On the mathematical modelling of competitive invasive weed dynamics. Bulletin of Mathematical Biology, 83(2):??, February 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00825-9;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00825-9.pdf.

Oteo-Garcia:2021:GFS

[169] Gonzalo Oteo-García and José-Angel Oteo. A geometrical framework for f-statistics. Bulletin of Mathematical Biology, 83(2):??, February 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00850-8;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00850-8.pdf.

Milzman:2021:MLM

[170] Jesse Milzman, Wanqiang Sheng, and Doron Levy. Modeling LSD1mediated tumor stagnation. Bulletin of Mathematical Biology, 83(2): ??, February 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00842-8;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00842-8.pdf.

Wang:2021:MRT

[171] Xiunan Wang, Hao Wang, and Michael Y. Li. Modeling rabies transmission in spatially heterogeneous environments via θdiffusion. Bulletin of Mathematical Biology, 83(2):??, February 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00857-1;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00857-1.pdf.

FitzGerald:2021:RLD

[172] Cody FitzGerald and James Keener. Red light and the dormancygermination decision in Arabidopsis seeds. Bulletin of Mathematical Biology, 83(3):??, March 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00849-1;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00849-1.pdf. See correction [225].

Gupta:2021:NIE

[173] Churni Gupta, Necibe Tuncer, and Maia Martcheva. A network immunoepidemiological HIV model. Bulletin of Mathematical Biology, 83(3):??, March 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00855-3;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00855-3.pdf.

Su:2021:DLB

[174] Wei-Hung Su, Ching-Shan Chou, and Dongbin Xiu. Deep learning of biological models from data: Applications to ODE models. Bulletin of Mathematical Biology, 83(3):??, March 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00851-7;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00851-
7.pdf.

Garnier:2021:DGH

[175] Jimmy Garnier and Pierre Lafontaine. Dispersal and good habitat quality promote neutral genetic diversity in metapopulations. Bulletin of Mathematical Biology, 83(3):??, March 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00853-5;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00853-
5.pdf.

Ciocanel:2021:TDA

[176] Maria-Veronica Ciocanel, Riley Juenemann, Adriana T. Dawes, and Scott A. McKinley. Topological data analysis approaches to uncovering the timing of ring structure onset in filamentous networks. Bulletin of Mathematical Biology, 83(3):??, March 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00847-3;
 http://link.springer.com/content/pdf/10.1007/s11538-020-00847-
3.pdf.

McGahan:2021:MLM

[177] Ian McGahan, James Powell, and Elizabeth Spencer. 28 models later: Model competition and the Zombie Apocalypse. Bulletin of Mathematical Biology, 83(3):??, March 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-020-00845-5;

 http://link.springer.com/content/pdf/10.1007/s11538-020-00845-
5.pdf.

Larson:2021:BFE

[178] Karen Larson, Sarah D. Olson, and Anastasios Matzavinos. A Bayesian framework to estimate fluid and material parameters in microswimmer models. Bulletin of Mathematical Biology, 83(3):??, March 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00852-6;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00852-6.pdf.

Enciso:2021:SCD

[179] German Enciso, Christine Sütterlin, Ming Tan, and Frederic Y. M. Wan. Stochastic Chlamydia dynamics and optimal spread. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-020-00846-4;   http://link.springer.com/
content/pdf/10.1007/s11538-020-00846-4.pdf.

Auger:2021:TRC

[180] Pierre Auger and Ali Moussaoui. On the threshold of release of confinement in an epidemic SEIR model taking into account the protective effect of mask. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00858-8;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00858-8.pdf.

Kulesa:2021:MCI

[181] Paul M. Kulesa, Jennifer C. Kasemeier-Kulesa, Jason A. Morrison, Rebecca McLennan, Mary Cathleen McKinney, and Caleb Bailey. Modelling cell invasion: a review of what J. D. Murray and the embryo can teach us. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00859-7;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00859-7.pdf.

Ibrahim:2021:TDM

[182] Mahmoud A. Ibrahim and Attila Dénes. Threshold dynamics in a model for Zika virus disease with seasonality. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-020-00844-6;   http://link.springer.com/content/
pdf/10.1007/s11538-020-00844-6.pdf.

Cocconi:2021:RIS

[183] Luca Cocconi, Alexander Kuhn-Régnier, Malte Neuss, Ana B. Sendova-Franks, and Kim Christensen. Reconstructing the intrinsic statistical properties of intermittent locomotion through corrections for boundary effects. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00848-2;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00848-2.pdf.

Seirin-Lee:2021:RCM

[184] Sungrim Seirin-Lee. The role of cytoplasmic MEX-5/6 polarity in asymmetric cell division. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00860-0;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00860-0.pdf.

Lee:2021:ECA

[185] Pilhwa Lee. Electrodiffusion with calcium-activated potassium channels in dendritic spine. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00854-4;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00854-4.pdf.

Sneyd:2021:CDW

[186] James Sneyd, Elias Vera-Sigüenza, John Rugis, Nathan Pages, and David I. Yule. Calcium dynamics and water transport in salivary acinar cells. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
020-00841-9;   http://link.springer.com/content/pdf/10.1007/
s11538-020-00841-9.pdf.

Anita:2021:CSS

[187] Sebastian Aniţa, Vincenzo Capasso, and Simone Scacchi. Controlling the spatial spread of a Xylella epidemic. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-021-00861-z;   http://link.springer.com/content/
pdf/10.1007/s11538-021-00861-z.pdf.

Kawamura:2021:MMD

[188] Mari Kawamura, Kei Sugihara, Hisako Takigawa-Imamura, Toshiyuki Ogawa, and Takashi Miura. Mathematical modeling of dynamic cellular association patterns in seminiferous tubules. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-021-00863-x;   http://link.springer.com/content/
pdf/10.1007/s11538-021-00863-x.pdf.

Stolerman:2021:SAS

[189] Lucas M. Stolerman, Pradipta Ghosh, and Padmini Rangamani. Stability analysis of a signaling circuit with dual species of GT-Pase switches. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00864-w;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00864-w.pdf.

El-Hachem:2021:IRS

[190] Maud El-Hachem, Scott W. McCue, and Matthew J. Simpson. Invading and receding sharp-fronted travelling waves. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-021-00862-y;   http://link.springer.com/content/
pdf/10.1007/s11538-021-00862-y.pdf.

Lanyon:2021:MII

[191] Christopher W. Lanyon, John R. King, Dov J. Stekel, and Rachel L. Gomes. A model to investigate the impact of farm practice on antimicrobial resistance in UK dairy farms. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-021-00865-9;   http://link.springer.com/content/
pdf/10.1007/s11538-021-00865-9.pdf.

Wang:2021:DIM

[192] Xiaoqiang Wang and Liyong Zhu. Diffusive interface model for actomyosin driven cell oscillations. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00866-8;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00866-8.pdf.

Fedak:2021:AAA

[193] Elizabeth A. Fedak, Frederick R. Adler, Lisa M. Abegglen, and Joshua D. Schiffman. ATM and ATR activation through crosstalk between DNA damage response pathways. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00868-6;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00868-6.pdf.

Roda:2021:MEL

[194] Weston C. Roda, Suli Liu, Christopher Power, and Michael Y. Li. Modeling the effects of latency reversing drugs during HIV-1 and SIV brain infection with implications for the “shock and kill” strategy. Bulletin of Mathematical Biology, 83(4):??, April 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00875-7;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00875-7.pdf.

Pooladvand:2021:RAC

[195] Pantea Pooladvand, Peter S. Kim, and Barbara Fazekas de St Groth. The role of antigen-competitive dynamics in regulating the immune response. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00867-7;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00867-7.pdf.

Woolley:2021:BTS

[196] Thomas E. Woolley, Andrew L. Krause, and Eamonn A. Gaffney. Bespoke Turing systems. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00870-y;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00870-y.pdf.

Owens:2021:MCC

[197] Katherine Owens and Ivana Bozic. Modeling CAR T-cell therapy with patient preconditioning. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00869-5;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00869-5.pdf.

Li:2021:GDR

[198] Fuxiang Li and Xiao-Qiang Zhao. Global dynamics of a reaction–diffusion model of Zika virus transmission with seasonality. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-021-00879-3;   http://link.springer.com/
content/pdf/10.1007/s11538-021-00879-3.pdf.

Jin:2021:MMT

[199] Wang Jin, Loredana Spoerri, Nikolas K. Haass, and Matthew J. Simpson. Mathematical model of tumour spheroid experiments with realtime cell cycle imaging. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00878-4;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00878-4.pdf.

Xiao:2021:EPD

[200] Siheng Xiao, Yuanshi Wang, and Shikun Wang. Effects of prey’s diffusion on predator–prey systems with two patches. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-021-00884-6;   http://link.springer.com/content/
pdf/10.1007/s11538-021-00884-6.pdf.

Hamelin:2021:OCP

[201] F. M. Hamelin, B. Bowen, P. Bernhard, and V. A. Bokil. Optimal control of plant disease epidemics with clean seed usage. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-021-00872-w;   http://link.springer.com/content/
pdf/10.1007/s11538-021-00872-w.pdf.

Pearce:2021:MPS

[202] Katherine J. Pearce, Kimberly Nellenbach, Ralph C. Smith, Ashley C. Brown, and Mansoor A. Haider. Modeling and parameter subset selection for fibrin polymerization kinetics with applications to wound healing. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00876-6;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00876-6.pdf.

Feng:2021:MSB

[203] Dianlei Feng, Insa Neuweiler, Regina Nogueira, and Udo Nackenhorst. Modeling of symbiotic bacterial biofilm growth with an example of the Streptococcus–Veillonella sp. system. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-021-00888-2;   http://link.springer.com/content/
pdf/10.1007/s11538-021-00888-2.pdf.

Wu:2021:DCS

[204] Hong Wu and Yuanshi Wang. Dynamics of competitive systems with diffusion between source–sink patches. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-021-00885-5;   http://link.springer.com/content/
pdf/10.1007/s11538-021-00885-5.pdf.

Ratto:2021:PSM

[205] N. Ratto, A. Bouchnita, P. Chelle, M. Marion, M. Panteleev, D. Nechipurenko, B. Tardy-Poncet, and V. Volpert. Patient-specific modelling of blood coagulation. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00890-8;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00890-8.pdf.

Caudera:2021:MMS

[206] Elisa Caudera, Simona Viale, Sandro Bertolino, Jacopo Cerri, and Ezio Venturino. A mathematical model supporting a hyperpredation effect in the apparent competition between invasive eastern cottontail and native European hare. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
021-00873-9;   http://link.springer.com/content/pdf/10.1007/
s11538-021-00873-9.pdf.

Manna:2021:PFT

[207] Kalyan Manna, Vitaly Volpert, and Malay Banerjee. Pattern formation in a three-species cyclic competition model. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-021-00886-4;   http://link.springer.com/content/
pdf/10.1007/s11538-021-00886-4.pdf.

Mallela:2021:RSN

[208] Abhishek Mallela and Alan Hastings. The role of stochasticity in noise-induced tipping point cascades: a master equation approach. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-021-00889-1;
 http://link.springer.com/content/pdf/10.1007/s11538-021-00889-
1.pdf.

DeStefano:2021:PLT

[209] Alisa DeStefano, Clyde Martin, and Dorothy Wallace. Predicting longterm asbestos prevalence in human lungs, lymph nodes, and remote organs from short-term murine experiments. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00882-8.

Levy:2021:MEH

[210] Ben Levy, Hannah E. Correia, and K. A. Jane White. Modeling the effect of HIV/AIDS stigma on HIV infection dynamics in Kenya. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00891-7. See [226].

Luke:2021:PEM

[211] Rayanne A. Luke, Richard J. Braun, and Carolyn G. Begley. Parameter estimation for mixed-mechanism tear film thinning. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00871-x.

Sooknanan:2021:HSM

[212] Joanna Sooknanan and Nicholas Mays. Harnessing social media in the modelling of pandemics — challenges and opportunities. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00895-3.

Hu:2021:MCB

[213] Linchao Hu, Cui Yang, and Jianshe Yu. Mosquito control based on pesticides and endosymbiotic bacterium Wolbachia. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00881-9.

Li:2021:EAA

[214] Ao Li and Xingfu Zou. Evolution and adaptation of anti-predation response of prey in a two-patchy environment. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00893-5.

Ford:2021:TDM

[215] Noah Ford, Garth Fisher, and David Chopp. A two-dimensional model of potassium signaling and oscillatory growth in a biofilm. Bulletin of Mathematical Biology, 83(5):??, May 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00887-3. See correction [324].

Zhang:2021:STP

[216] Xue Zhang and Jianhong Wu. Synchronized tick population oscillations driven by host mobility and spatially heterogeneous developmental delays combined. Bulletin of Mathematical Biology, 83(6):??, June 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-021-00874-8.

Chen:2021:COC

[217] Kuan-Wei Chen and Chih-Wen Shih. Collective oscillations in coupled-cell systems. Bulletin of Mathematical Biology, 83(6):??, June 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-021-00883-7.

Kalbaugh:2021:PMC

[218] David V. Kalbaugh. Probabilistic model for control of an epidemic by isolation and quarantine. Bulletin of Mathematical Biology, 83(6):??, June 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00897-1.

Kreusser:2021:ADA

[219] Lisa Maria Kreusser and Alan D. Rendall. Autophosphorylation and the dynamics of the activation of Lck. Bulletin of Mathematical Biology, 83(6):??, June 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00900-9.

Koch:2021:SEP

[220] Dean Koch, Mark A. Lewis, and Subhash Lele. The signature of endemic populations in the spread of mountain pine beetle outbreaks. Bulletin of Mathematical Biology, 83(6):??, June 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00899-z.

Adams:2021:HSI

[221] Ben Adams, Katharine S. Walter, and Maria A. Diuk-Wasser. Host specialisation, immune cross-reaction and the composition of communities of co-circulating Borrelia strains. Bulletin of Mathematical Biology, 83(6):??, June 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00896-2.

Britton:2021:ECO

[222] Nicholas F. Britton and K. A. Jane White. The effect of covert and overt infections on disease dynamics in honey-bee colonies. Bulletin of Mathematical Biology, 83(6):??, June 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00892-6.

Collin:2021:MSI

[223] Annabelle Collin, Vladimir Groza, and Olivier Saut. A modelstrengthened imaging biomarker for survival prediction in EGFR-mutated non-small-cell lung carcinoma patients treated with tyrosine kinase inhibitors. Bulletin of Mathematical Biology, 83(6):??, June 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-021-00902-7.

Bernardi:2021:LTI

[224] Sara Bernardi, Raluca Eftimie, and Kevin J. Painter. Leadership through influence: What mechanisms allow leaders to steer a swarm? Bulletin of Mathematical Biology, 83(6):??, June 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00901-8.

FitzGerald:2021:CRL

[225] Cody FitzGerald and James Keener. Correction to: Red Light and the Dormancy–Germination Decision in Arabidopsis Seeds. Bulletin of Mathematical Biology, 83(6):??, June 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00894-4. See [172].

Levy:2021:CME

[226] Ben Levy, Hannah E. Correia, and K. A. Jane White. Correction to: Modeling the Effect of HIV/AIDS Stigma on HIV Infection Dynamics in Kenya. Bulletin of Mathematical Biology, 83(6):??, June 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00904-5. See [210].

Amato:2021:MMA

[227] Sara Amato and Andrea Arnold. Modeling microglia activation and inflammation-based neuroprotectant strategies during ischemic stroke. Bulletin of Mathematical Biology, 83(6):??, June 2021. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-021-00905-4.

Kempes:2021:GSB

[228] Christopher P. Kempes, Michael J. Follows, and Simon A. Levin. Generalized stoichiometry and biogeochemistry for astrobiological applications. Bulletin of Mathematical Biology, 83(7):??, July 2021. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-021-00877-5.

Ferrari:2021:CEM

[229] Carlo Giambiagi Ferrari, Juan Pablo Pinasco, and Nicolas Saintier. Coupling epidemiological models with social dynamics. Bulletin of Mathematical Biology, 83(7):??, July 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00910-7.

Senekal:2021:NKC

[230] Noma Susan Senekal, Khaphetsi Joseph Mahasa, and Rachid Ouifki. Natural killer cells recruitment in oncolytic virotherapy: a mathematical model. Bulletin of Mathematical Biology, 83(7):??, July 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-021-00903-6.

Hernandez:2021:IDC

[231] Bryan S. Hernandez and Ralph John L. De la Cruz. Independent decompositions of chemical reaction networks. Bulletin of Mathematical Biology, 83(7):??, July 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00906-3.

Corcoran:2021:LDN

[232] Carl Corcoran and Alan Hastings. A low-dimensional network model for an SIS epidemic: Analysis of the super compact pairwise model. Bulletin of Mathematical Biology, 83(7):??, July 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00907-2.

Hendriksen:2021:CIS

[233] Michael Hendriksen and Nils Kapust. On the comparison of incompatibility of split systems across different numbers of taxa. Bulletin of Mathematical Biology, 83(7):??, July 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00911-6.

Mochan:2021:CMS

[234] Ericka Mochan, T. J. Sego, and G. Bard Ermentrout. Compartmental model suggests importance of innate immune response to COVID-19 infection in Rhesus macaques. Bulletin of Mathematical Biology, 83(7):??, July 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00909-0.

Villa:2021:MMP

[235] Chiara Villa, Mark A. J. Chaplain, and Tommaso Lorenzi. Mechanical models of pattern and form in biological tissues: The role of stress– strain constitutive equations. Bulletin of Mathematical Biology, 83(7):??, July 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00912-5.

Vaidya:2021:MID

[236] Naveen K. Vaidya and Michael Peter. Modeling intracellular delay in within-host HIV dynamics under conditioning of drugs of abuse. Bulletin of Mathematical Biology, 83(7):??, July 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00908-1.

Krause:2021:IPO

[237] Andrew L. Krause, Václav Klika, and Eamonn A. Gaffney. Isolating patterns in open reaction–diffusion systems. Bulletin of Mathematical Biology, 83(7):??, July 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00913-4.

Fiandaca:2021:MSI

[238] Giada Fiandaca, Marcello Delitala, and Tommaso Lorenzi. A mathematical study of the influence of hypoxia and acidity on the evolutionary dynamics of cancer. Bulletin of Mathematical Biology, 83(7):??, July 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-021-00914-3.

Callahan:2021:CAL

[239] Jason Callahan, Eric Eager, and Shenglan Yuan. Correction to: Analysis of a Length-Structured Density-Dependent Model for Fish. Bulletin of Mathematical Biology, 83(8):??, August 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00916-1. See [2].

Camargo:2021:MRB

[240] Felipe de A. Camargo, Mostafa Adimy, and Cláudia P. Ferreira. Modeling the relationship between antibody-dependent enhancement and disease severity in secondary dengue infection. Bulletin of Mathematical Biology, 83(8):??, August 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00919-y.

Gross:2021:RPS

[241] Kevin Gross and André M. de Roos. Resonance in physiologically structured population models. Bulletin of Mathematical Biology, 83(8):??, August 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-

021-00915-2. See correction [257].

AlBasir:2021:EVM

[242] F. Al Basir, Y. N. Kyrychko, and S. Ray. Effects of vector maturation time on the dynamics of cassava mosaic disease. Bulletin of Mathematical Biology, 83(8):??, August 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00921-4.

Herrero:2021:RMC

[243] Maria Elena Gonzalez Herrero, Christian Kuehn, and Krasimira Tsaneva-Atanasova. Reduced models of cardiomyocytes excitability: Comparing Karma and FitzHugh–Nagumo. Bulletin of Mathematical Biology, 83(8):??, August 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00898-0.

Paez:2021:ASE

[244] Gustavo Nicolás Páez, Juan Felipe Cerón, and Carlos Castañeda. Alternative strategies for the estimation of a disease’s basic reproduction number: a model-agnostic study. Bulletin of Mathematical Biology, 83(8):??, August 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00922-3.

Kenne:2021:ASM

[245] Cyrille Kenne, René Dorville, and Pascal Zongo. An age-structured model for tilapia lake virus transmission in freshwater with vertical and horizontal transmission. Bulletin of Mathematical Biology, 83(8):??, August 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-021-00923-2.

Puchhammer:2021:VRA

[246] Florian Puchhammer, Amal Ben Abdellah, and Pierre L’Ecuyer. Variance reduction with array-RQMC for tau-leaping simulation of stochastic biological and chemical reaction networks. Bulletin of Mathematical Biology, 83(8):??, August 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00920-5.

Stotsky:2021:RWA

[247] Jay A. Stotsky, Jia Gou, and Hans G. Othmer. A random walk approach to transport in tissues and complex media: From microscale descriptions to macroscale models. Bulletin of Mathematical Biology, 83(9):??, September 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00917-0.

Richards:2021:BWT

[248] Andrew Richards and Laura Kubatko. Bayesian-weighted triplet and quartet methods for species tree inference. Bulletin of Mathematical Biology, 83(9):??, September 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00918-z.

Are:2021:DCD

[249] Elisha B. Are, John W. Hargrove, and Jonathan Dushoff. Does counting different life stages impact estimates for extinction probabilities for tsetse (Glossina spp)? Bulletin of Mathematical Biology, 83(9):??, September 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00924-1.

Efendiev:2021:DEU

[250] M. A. Efendiev, J. Murley, and S. Sivaloganathan. Dimension estimate of uniform attractor for a model of high intensity focussed ultrasound-induced thermotherapy. Bulletin of Mathematical Biology, 83 (9):??, September 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00928-x.

Lui:2021:MPP

[251] Gigi Lui and Mary R. Myerscough. Modelling preferential phagocytosis in atherosclerosis: Delineating timescales in plaque development. Bulletin of Mathematical Biology, 83(9):??, September 2021. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-021-00926-z.

Phillips:2021:DDM

[252] Tricia Phillips, Suzanne Lenhart, and W. Christopher Strickland. A datadriven mathematical model of the heroin and fentanyl epidemic in Tennessee. Bulletin of Mathematical Biology, 83(10):??, October 2021. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-021-00925-0.

Ruiz-Herrera:2021:RMP

[253] Alfonso Ruiz-Herrera and Pedro J. Torres. The role of movement patterns in epidemic models on complex networks. Bulletin of Mathematical Biology, 83(10):??, October 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00929-w.

Sherratt:2021:CRL

[254] Jonathan A. Sherratt, Quan-Xing Liu, and Johan van de Koppel. A comparison of the “Reduced Losses” and “Increased Production” models for mussel bed dynamics. Bulletin of Mathematical Biology, 83(10):??, October 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00932-1.

Mondal:2021:BTA

[255] Ritwika Mondal, Suman Saha, and Debasis Mukherjee. Basin transition and alternative states: Role of multi-species herbivores-induced volatile in plant–insect interactions. Bulletin of Mathematical Biology, 83(10):??, October 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00930-3.

Fomba:2021:PEP

[256] C. G. Nembot Fomba, G. M. ten Hoopen, and P. Takam Soh. Parameter estimation in a PDE model for the spatial spread of cocoa black pod disease. Bulletin of Mathematical Biology, 83(10):??, October 2021. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-021-00934-z.

Gross:2021:CRP

[257] Kevin Gross and André M. de Roos. Correction to: Resonance in Physiologically Structured Population Models. Bulletin of Mathematical Biology, 83(10):??, October 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00931-2. See [241].

Hamis:2021:TCD

[258] Sara Hamis, James Yates, and Gibin G. Powathil. Targeting cellular DNA damage responses in cancer: an in vitro-calibrated agent-based model simulating monolayer and spheroid treatment responses to ATR-inhibiting drugs. Bulletin of Mathematical Biology, 83(10):??, October 2021. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-021-00935-y.

Fowler:2021:AFO

[259] A. C. Fowler. Atto-foxes and other minutiae. Bulletin of Mathematical Biology, 83(10):??, October 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00936-x.

Gaebler:2021:MMU

[260] Harry J. Gaebler and Hermann J. Eberl. Multiscale modeling of uranium bioreduction in porous media by one-dimensional biofilms. Bulletin of Mathematical Biology, 83(10):??, October 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00938-9.

Zheng:2021:MMD

[261] Collin Y. Zheng and Peter S. Kim. Mathematical model for delayed responses in immune checkpoint blockades. Bulletin of Mathematical Biology, 83(10):??, October 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00933-0.

DeAngelis:2021:TBS

[262] Donald L. DeAngelis, Daniel Franco, and Rebecca C. Tyson. Towards building a sustainable future: Positioning ecological modelling for impact in ecosystems management. Bulletin of Mathematical Biology, 83 (10):??, October 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00927-y.

Xue:2021:EHA

[263] Shuyang Xue, Meili Li, and Jia Li. The effect of harvesting adults on the evolution of reproduction age via density-dependent juvenile mortality. Bulletin of Mathematical Biology, 83(10):??, October 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-021-00940-1.

Jiang:2021:TPM

[264] Hongyan Jiang, King-Yeung Lam, and Yuan Lou. Three-patch models for the evolution of dispersal in advective environments: Varying drift and network topology. Bulletin of Mathematical Biology, 83(10):??, October 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00939-8.

Chowdhury:2021:OPF

[265] Pranali Roy Chowdhury, Sergei Petrovskii, and Malay Banerjee. Oscillations and pattern formation in a slow–fast prey–predator system. Bulletin of Mathematical Biology, 83(11):??, November 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/article/10.1007/s11538-021-00941-0. See correction

[274].

Yuan:2021:ZEM

[266] Xiaoyan Yuan, Yijun Lou, and Daozhou Gao. A Zika endemic model for the contribution of multiple transmission routes. Bulletin of Mathematical Biology, 83(11):??, November 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00945-w.

Mallela:2021:TCM

[267] Abhishek Mallela and Alan Hastings. Tipping cascades in a multi-patch system with noise and spatial coupling. Bulletin of Mathematical Biology, 83(11):??, November 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00943-y.

Hoyer-Leitzel:2021:IFD

[268] Alanna Hoyer-Leitzel and Sarah Iams. Impulsive fire disturbance in a savanna model: Tree–grass coexistence states, multiple stable system states, and resilience. Bulletin of Mathematical Biology, 83(11):??, November 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00944-x.

Besse:2021:SPS

[269] Christophe Besse and Grégory Faye. Spreading properties for SIR models on homogeneous trees. Bulletin of Mathematical Biology, 83(11): ??, November 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00948-7.

Plaugher:2021:MPC

[270] Daniel Plaugher and David Murrugarra. Modeling the pancreatic cancer microenvironment in search of control targets. Bulletin of Mathematical Biology, 83(11):??, November 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00937-w.

Li:2021:MRI

[271] Xue-Zhi Li, Shasha Gao, and Maia Martcheva. Modeling and research on an immuno-epidemiological coupled system with coinfection. Bulletin of Mathematical Biology, 83(11):??, November 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/article/10.1007/s11538-021-00946-9.

Lauro:2021:ICS

[272] Francesco Di Lauro, Luc Berthouze, and István Z. Kiss. The impact of contact structure and mixing on control measures and disease-induced herd immunity in epidemic models: a mean-field model perspective. Bulletin of Mathematical Biology, 83(11):??, November 2021. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-021-00947-8.

Pascucci:2021:MIM

[273] Eleonora Pascucci and Andrea Pugliese. Modelling immune memory development. Bulletin of Mathematical Biology, 83(12):??, December 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-021-00949-6.

Chowdhury:2021:COP

[274] Pranali Roy Chowdhury, Sergei Petrovskii, and Malay Banerjee. Correction to: Oscillations and Pattern Formation in a Slow–Fast Prey– Predator System. Bulletin of Mathematical Biology, 83(12):??, December 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00954-9. See [265].

Yamagishi:2021:MMW

[275] Jumpei F. Yamagishi and Tetsuhiro S. Hatakeyama. Microeconomics of metabolism: The Warburg Effect as Giffen behaviour. Bulletin of Mathematical Biology, 83(12):??, December 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00952-x.

Berg:2021:MOP

[276] Mascha Berg, Julia Plöntzke, and Susanna Röblitz. Modelling oscillatory patterns in the bovine estrous cycle with Boolean delay equations. Bulletin of Mathematical Biology, 83(12):??, December 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-021-00942-z.

Tenore:2021:MMN

[277] A. Tenore, F. Russo, and L. Frunzo. Multiscale modelling of De Novo anaerobic granulation. Bulletin of Mathematical Biology, 83(12): ??, December 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00951-y.

Talkington:2021:EDP

[278] Anne M. Talkington, Timothy Wessler, and M. Gregory Forest. Experimental data and PBPK modeling quantify antibody interference in PEGylated drug carrier delivery. Bulletin of Mathematical Biology, 83 (12):??, December 2021. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00950-z.

Wren:2021:HLI

[279] Lydia Wren and Alex Best. How local interactions impact the dynamics of an epidemic. Bulletin of Mathematical Biology, 83(12):??, December 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-021-00961-w.

Zhou:2022:RSU

[280] Ying Zhou. Range shifts under constant-speed and accelerated climate warming. Bulletin of Mathematical Biology, 84(1):??, January 2022. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-021-00963-8.

Pasetto:2022:IHT

[281] S. Pasetto, H. Enderling, and R. Brady-Nicholls. Intermittent hormone therapy models analysis and Bayesian model comparison for prostate cancer. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-021-00953-w.

Akman:2022:HLS

[282] Olcay Akman, Sudipa Chauhan, and Jai Prakash Tripathi. The hard lessons and shifting modeling trends of COVID-19 dynamics: Multiresolution modeling approach. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00959-4.

Overton:2022:AQS

[283] Christopher E. Overton, Robert R. Wilkinson, and Kieran J. Sharkey. Approximating quasi-stationary behaviour in network-based SIS dynamics. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-021-00964-7.

Yan:2022:DSA

[284] Yawen Yan, Jimin Zhang, and Hao Wang. Dynamics of stoichiometric autotroph–mixotroph–bacteria interactions in the epilimnion. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00962-9.

Bogdanov:2022:SMV

[285] Anita Bogdanov, Péter Kevei, and Dezső Virok. Stochastic modeling of in vitro bactericidal potency. Bulletin of Mathematical Biology, 84 (1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00967-4.

Eilertsen:2022:VSQ

[286] Justin Eilertsen and Santiago Schnell. On the validity of the stochastic quasi-steady-state approximation in open enzyme catalyzed reactions: Timescale Separation or Singular Perturbation? Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00966-5.

Cook:2022:MCI

[287] Blake Cook, Alex Combes, and James M. Osborne. Modelling cellular interactions and dynamics during kidney morphogenesis. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00968-3.

Park:2022:MDB

[288] Jeungeun Park and Zahra Aminzare. A mathematical description of bacterial chemotaxis in response to two stimuli. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00965-6.

Lou:2022:CSM

[289] Yuan Lou and Rachidi B. Salako. Control strategies for a multi-strain epidemic model. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-021-00957-6.

Heumann:2022:DDR

[290] Ryan Heumann, Carly Duncan, and Scott Kaschner. Dynamical differences in respiratory syncytial virus. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00971-8.

Murray:2022:ST

[291] James D. Murray. Special tributes. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00956-7.

Maini:2022:SCC

[292] Philip K. Maini, Mark A. J. Chaplain, and Jonathan A. Sherratt. Special collection: Celebrating J. D. Murray’s contributions to mathematical biology. Bulletin of Mathematical Biology, 84(1):??, January 2022. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-021-00955-8.

FitzGerald:2022:BAH

[293] Cody FitzGerald and James P. Keener. Bifurcation analysis of a heatsensitive epigenetic regulatory network. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00960-x.

Stutz:2022:CTA

[294] Timothy C. Stutz, Janet S. Sinsheimer, and Jason Xu. Computational tools for assessing gene therapy under branching process models of mutation. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-021-00969-2.

Mudaliar:2022:SAE

[295] Rajnesh K. Mudaliar, Andrei V. Zvezdin, and Timothy M. Schaerf. Systematic analysis of emergent collective motion produced by a 3D hybrid zonal model. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-

021-00977-2. See correction [327].

Jaramillo:2022:DIH

[296] J. M. Jaramillo, Junling Ma, and Abdul-Aziz Yakubu. Disease-induced hydra effect with overcompensatory recruitment. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00975-4.

Bakshi:2022:CMM

[297] Suruchi Bakshi, Fraser Cunningham, and Piet H. van der Graaf. Correction to: Mathematical Modelling of Alternative Pathway of Complement System. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00981-6. See [35].

Adimy:2022:WPE

[298] Mostafa Adimy, Louis Babin, and Laurent Pujo-Menjouet. Why are periodic erythrocytic diseases so rare in humans? Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00973-6.

Counterman:2022:DDR

[299] Elijah D. Counterman and Sean D. Lawley. Designing drug regimens that mitigate nonadherence. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00976-3.

Li:2022:EBP

[300] Yifei Li, Stuart T. Johnston, and Matthew J. Simpson. Extinction of bistable populations is affected by the shape of their initial spatial distribution. Bulletin of Mathematical Biology, 84(1):??, January 2022. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-021-00974-5.

Simoy:2022:VBD

[301] Mario Ignacio Simoy and Juan Pablo Aparicio. Vector-borne disease models with active and inactive vectors: a simple way to consider biting behavior. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-021-00972-7.

Garay:2022:BRP

[302] József Garay and Tamás F. Móri. Best reply player against mixed evolutionarily stable strategy user. Bulletin of Mathematical Biology, 84 (1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00980-7.

Aguade-Gorgorio:2022:TTT

[303] Guim Aguadé-Gorgorió, Stuart Kauffman, and Ricard Solé. Transition therapy: Tackling the ecology of tumor phenotypic plasticity. Bulletin of Mathematical Biology, 84(1):??, January 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00970-9.

Okabe:2022:PBG

[304] Takuya Okabe and Jin Yoshimura. Practical basis of the geometric mean fitness and its application to risk-spreading behavior. Bulletin of Mathematical Biology, 84(2):??, February 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00984-3.

Willson:2022:MAP

[305] Stephen J. Willson. Merging arcs to produce acyclic phylogenetic networks and normal networks. Bulletin of Mathematical Biology, 84(2):??, February 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00986-1.

Gulbudak:2022:SAI

[306] Hayriye Gulbudak, Zhuolin Qu, and Necibe Tuncer. Sensitivity analysis in an immuno-epidemiological vector-host model. Bulletin of Mathematical Biology, 84(2):??, February 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-021-00979-0.

Zhou:2022:MEJ

[307] Linhua Zhou, Xinmiao Rong, and Huaiping Zhu. Modeling and evaluation of the joint prevention and control mechanism for curbing COVID-19 in Wuhan. Bulletin of Mathematical Biology, 84(2):??, February 2022. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-021-00983-4.

Miles:2022:MTP

[308] Christopher E. Miles, Jie Zhu, and Alex Mogilner. Mechanical torque promotes bipolarity of the mitotic spindle through multi-centrosomal clustering. Bulletin of Mathematical Biology, 84(2):??, February 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-021-00985-2.

Zou:2022:VQE

[309] Yukun Zou, Wei Yang, and Wei Lin. Vaccination and quarantine effect on COVID-19 transmission dynamics incorporating Chinese-spring-festival travel rush: Modeling and simulations. Bulletin of Mathematical Biology, 84(2):??, February 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-021-00958-5.

Sun:2022:ESM

[310] Xiaoli Sun, Shengqiang Liu, and Yongzhen Pei. Effects of sterile males and fertility of infected mosquitoes on mosquito-borne disease dynamics. Bulletin of Mathematical Biology, 84(2):??, February 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-022-00991-y.

Musa:2022:HSC

[311] Salihu Sabiu Musa, Xueying Wang, and Daihai He. The heterogeneous severity of COVID-19 in African countries: a modeling approach. Bulletin of Mathematical Biology, 84(3):??, March 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-00992-x.

Cortes-Poza:2022:VAM

[312] Yuriria Cortés-Poza and Pablo Padilla-Longoria. A variational approach to morphogenesis. Bulletin of Mathematical Biology, 84(3):??, March 2022.

CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-00993-w.

Feng:2022:FKM

[313] Chunxi Feng, Mark A. Lewis, and Hao Wang. A Fisher–KPP model with a nonlocal weighted free boundary: Analysis of how habitat boundaries expand, balance or shrink. Bulletin of Mathematical Biology, 84(3):??, March 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
022-00995-8.

Maini:2022:EJC

[314] Philip K. Maini, Peter J. Hunter, and Nic P. Smith. Edmund John Crampin 1973–2021. Bulletin of Mathematical Biology, 84(3):??, March 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
021-00987-0.

Papst:2022:MIB

[315] Irena Papst, Kevin P. O’Keeffe, and Steven H. Strogatz. Modeling the interplay between seasonal flu outcomes and individual vaccination decisions. Bulletin of Mathematical Biology, 84(3):??, March 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-021-00988-z.

Shirani:2022:CTV

[316] Farshad Shirani and Judith R. Miller. Competition, trait variance dynamics, and the evolution of a species’ range. Bulletin of Mathematical Biology, 84(3):??, March 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-00990-z.

Cui:2022:ENH

[317] Jingan Cui, Yucui Wu, and Songbai Guo. Effect of non-homogeneous mixing and asymptomatic individuals on final epidemic size and basic reproduction number in a meta-population model. Bulletin of Mathematical Biology, 84(3):??, March 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-00996-7.

Sher:2022:QSP

[318] Anna Sher, Steven A. Niederer, and Denis Noble. A quantitative systems pharmacology perspective on the importance of parameter identifiability. Bulletin of Mathematical Biology, 84(3):??, March 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-021-00982-5.

Zmurchok:2022:BMP

[319] Cole Zmurchok and William R. Holmes. Biophysical models of PAR cluster transport by cortical flow in C. elegans early embryogenesis. Bulletin of Mathematical Biology, 84(3):??, March 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-00997-6.

Maity:2022:CDS

[320] Sunil Maity and Partha Sarathi Mandal. A comparison of deterministic and stochastic plant-vector-virus models based on probability of disease extinction and outbreak. Bulletin of Mathematical Biology, 84(3):??, March 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
022-01001-x.

Conte:2022:MCK

[321] Martina Conte and Nadia Loy. Multi-cue kinetic model with non-local sensing for cell migration on a fiber network with chemotaxis. Bulletin of Mathematical Biology, 84(3):??, March 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-021-00978-1.

Garfinkel:2022:TDB

[322] Alan Garfinkel, Steve Bennoun, and Blaire Van Valkenburgh. Teaching dynamics to biology undergraduates: the UCLA experience. Bulletin of Mathematical Biology, 84(3):??, March 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-00999-4.

Kreck:2022:BRD

[323] Matthias Kreck and Erhard Scholz. Back to the roots: a discrete Kermack–McKendrick model adapted to Covid-19. Bulletin of Mathematical Biology, 84(4):??, April 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-00994-9.

Ford:2022:CTD

[324] Noah Ford, Garth Fisher, and David Chopp. Correction to: A Two-Dimensional Model of Potassium Signaling and Oscillatory Growth in a Biofilm. Bulletin of Mathematical Biology, 84(4):??, April 2022. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01003-9.

See [215].

Milne:2022:LOP

[325] Russell Milne, Chris T. Bauch, and Madhur Anand. Local overfishing patterns have regional effects on health of coral, and economic transitions can promote its recovery. Bulletin of Mathematical Biology, 84(4):??, April 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
022-01000-y.

Feng:2022:IQM

[326] Shanshan Feng, Juping Zhang, and Zhen Jin. The impact of quarantine and medical resources on the control of COVID-19 in Wuhan based on a household model. Bulletin of Mathematical Biology, 84(4):??, April 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-021-00989-y.

Mudaliar:2022:CSA

[327] Rajnesh K. Mudaliar, Andrei V. Zvezdin, and Timothy M. Schaerf. Correction to: Systematic analysis of emergent collective motion produced by a 3D hybrid zonal model. Bulletin of Mathematical Biology, 84(5):??, May 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
022-01008-4. See [295].

Gerlee:2022:WSS

[328] Philip Gerlee. Weak selection and the separation of eco-evo time scales using perturbation analysis. Bulletin of Mathematical Biology, 84(5):??, May 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
022-01009-3.

Bortner:2022:IPC

[329] Cashous Bortner and Nicolette Meshkat. Identifiable paths and cycles in linear compartmental models. Bulletin of Mathematical Biology, 84(5):??, May 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
022-01007-5.

Landa:2022:PAS

[330] Kylie J. Landa, Lauren M. Mossman, and Sara M. Clifton. Phage– antibiotic synergy inhibited by temperate and chronic virus competition.

Bulletin of Mathematical Biology, 84(5):??, May 2022. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-022-01006-6.

Hu:2022:EML

[331] Lin Hu, Shengfu Wang, and Zhidong Teng. The effects of migration and limited medical resources of the transmission of SARS-CoV-2 model with two patches. Bulletin of Mathematical Biology, 84(5):??, May 2022. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01010-w.

Youlden:2022:MMH

[332] George Youlden, Helen E. McNeil, and John R. King. Mathematical modelling highlights the potential for genetic manipulation as an adjuvant to counter efflux-mediated MDR in Salmonella. Bulletin of Mathematical Biology, 84(5):??, May 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01011-9.

Wang:2022:HFB

[333] Xiunan Wang, Hao Wang, and Mark Lewis. A hypothesis-free bridging of disease dynamics and non-pharmaceutical policies. Bulletin of Mathematical Biology, 84(5):??, May 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-01012-8.

Swanson:2022:MMT

[334] Ellen R. Swanson, Emek Köse, and Samantha L. Elliott. Mathematical modeling of tumor and cancer stem cells treated with CAR-T therapy and inhibition of TGF-β. Bulletin of Mathematical Biology, 84(6):??, June 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
022-01015-5.

Bottcher:2022:RSI

[335] Lucas Böttcher and Hans Gersbach. A refunding scheme to incentivize narrow-spectrum antibiotic development. Bulletin of Mathematical Biology, 84(6):??, June 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01013-7.

Gordillo:2022:AEP

[336] Luis F. Gordillo and Priscilla E. Greenwood. Allee effects plus noise induce population dynamics resembling binary Markov highs and lows. Bulletin of Mathematical Biology, 84(6):??, June 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01017-3.

Xue:2022:EST

[337] Ling Xue, Shuanglin Jing, and Hao Wang. Evaluating strategies for tuberculosis to achieve the goals of WHO in China: A seasonal age-structured model study. Bulletin of Mathematical Biology, 84(6):??, June 2022. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01019-1.

Vallejo:2022:EPP

[338] Celeste Vallejo, Carl A. B. Pearson, and Thomas J. Hladish. Effect of population partitioning on the probability of silent circulation of poliovirus. Bulletin of Mathematical Biology, 84(6):??, June 2022. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-022-01014-6.

David:2022:EHM

[339] Jummy F. David and Sarafa A. Iyaniwura. Effect of human mobility on the spatial spread of airborne diseases: An epidemic model with indirect transmission. Bulletin of Mathematical Biology, 84(6):??, June 2022. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01020-8.

Cordelier:2022:SFM

[340] Pierre Cordelier, Manon Costa, and Jérôme Fehrenbach. Slow–fast model and therapy optimization for oncolytic treatment of tumors. Bulletin of Mathematical Biology, 84(6):??, June 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01025-3.

Craciun:2022:MCS

[341] Gheorghe Craciun, Badal Joshi, and Ike Tan. Multistationarity in cyclic sequestration-transmutation networks. Bulletin of Mathematical Biology, 84(6):??, June 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01021-7.

Gharouni:2022:TIE

[342] Ali Gharouni, Fred M. Abdelmalek, and Benjamin M. Bolker. Testing and isolation efficacy: Insights from a simple epidemic model. Bulletin of Mathematical Biology, 84(6):??, June 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01018-2.

Streipert:2022:DAD

[343] Sabrina H. Streipert, Gail S. K. Wolkowicz, and Martin Bohner. Derivation and analysis of a discrete predator–prey model. Bulletin of Mathematical Biology, 84(7):??, July 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-01016-4.

Hamelin:2022:HDM

[344] F. M. Hamelin, Y. Mammeri, and M. A. Lewis. Host diversification may split epidemic spread into two successive fronts advancing at different speeds. Bulletin of Mathematical Biology, 84(7):??, July 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-022-01023-5.

Han:2022:LBD

[345] Lifeng Han, Changhan He, and Yang Kuang. Learning biological dynamics from spatio-temporal data by Gaussian processes. Bulletin of Mathematical Biology, 84(7):??, July 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-01022-6.

Duron:2022:MFA

[346] Christina Durón and Alex Farrell. A mean-field approximation of SIR epidemics on an Erdős–Rényi network model. Bulletin of Mathematical Biology, 84(7):??, July 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01026-2.

Gong:2022:RSF

[347] Yishu Gong, Siming He, and Alexander Kiselev. Random search in fluid flow aided by chemotaxis. Bulletin of Mathematical Biology, 84(7):??, July 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
022-01024-4.

Strickland:2022:EPA

[348] W. C. Strickland, N. A. Battista, and L. A. Miller. Planktos: an agentbased modeling framework for small organism movement and dispersal in a fluid environment with immersed structures. Bulletin of Mathematical Biology, 84(7):??, July 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01027-1.

Mu:2022:TDT

[349] Yu Mu, Tsz-Lik Chan, and Wing-Cheong Lo. Transmission dynamics of tuberculosis with age-specific disease progression. Bulletin of Mathematical Biology, 84(7):??, July 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-01032-4.

Saona:2022:RBN

[350] Raimundo Saona, Fyodor A. Kondrashov, and Ksenia A. Khudiakova. Relation between the number of peaks and the number of reciprocal sign epistatic interactions. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-

022-01029-z. See correction [439].

Berec:2022:DME

[351] Luděk Berec, Jan Smyčka, and Petra Vidnerová. Delays, masks, the elderly, and schools: First Covid-19 wave in the Czech Republic. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01031-5.

vanIersel:2022:ONT

[352] Leo van Iersel, Remie Janssen, and Yukihiro Murakami. Orchard networks are trees with additional horizontal arcs. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01037-z.

Ledder:2022:MUR

[353] Glenn Ledder. Mentoring undergraduate research in mathematical modeling. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-022-01040-4.

Ghosh:2022:EMT

[354] Samiran Ghosh, Vitaly Volpert, and Malay Banerjee. An epidemic model with time-distributed recovery and death rates. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-01028-0.

Horstmeyer:2022:BQS

[355] Leonhard Horstmeyer, Christian Kuehn, and Stefan Thurner. Balancing quarantine and self-distancing measures in adaptive epidemic networks. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-022-01033-3.

Colson:2022:CMG

[356] Chloé Colson, Helen M. Byrne, and Philip K. Maini. Combining mechanisms of growth arrest in solid tumours: a mathematical investigation. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-022-01034-2.

Anwar:2022:MMM

[357] Md Nurul Anwar, Roslyn I. Hickson, and Jennifer A. Flegg. A multiscale mathematical model of Plasmodium Vivax transmission. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01036-0.

Liao:2022:MMA

[358] Kang-Ling Liao and Kenton D. Watt. Mathematical modeling and analysis of CD200–CD200R in cancer treatment. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01039-x.

Alharbi:2022:CMM

[359] Mohammed H. Alharbi and Christopher M. Kribs. Correction to: A mathematical modeling study: Assessing impact of mismatch between influenza vaccine strains and circulating strains in Hajj. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-01038-y. See [162].

Su:2022:MMS

[360] Shan Su, John Rugis, and James Sneyd. A mathematical model of salivary gland duct cells. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01041-3.

Ben-Ami:2022:SFM

[361] Y. Ben-Ami, G. W. Atkinson, and H. M. Byrne. Structural features of microvascular networks trigger blood flow oscillations. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-01046-y.

Fatoyinbo:2022:PFS

[362] H. O. Fatoyinbo, R. G. Brown, and B. van Brunt. Pattern formation in a spatially extended model of pacemaker dynamics in smooth muscle cells. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-022-01043-1.

Jepson:2022:TWA

[363] Jacob M. Jepson, Nabil T. Fadai, and Reuben D. O’Dea. Travelling-wave and asymptotic analysis of a multiphase moving boundary model for engineered tissue growth. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
022-01044-0.

Mochan:2022:ARC

[364] Ericka Mochan, T. J. Sego, and Bard Ermentrout. Age-related changes to the immune system exacerbate the inflammatory response to pandemic H1N1 infection. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01045-z.

Fox:2022:MTR

[365] Erika Fox, Bree Cummins, and Tomáš Gedeon. Modeling transport regulation in gene regulatory networks. Bulletin of Mathematical Biology, 84(8):??, August 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01035-1.

Wang:2022:PPF

[366] Xiunan Wang, Hao Wang, and Mark Lewis. From policy to prediction: Forecasting COVID-19 dynamics under imperfect vaccination. Bulletin of Mathematical Biology, 84(9):??, September 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/article/10.1007/s11538-022-01047-x.

Hwang:2022:HEM

[367] Karen K. L. Hwang, Christina J. Edholm, and Nika Shakiba. A hybrid epidemic model to explore stochasticity in COVID-19 dynamics. Bulletin of Mathematical Biology, 84(9):??, September 2022. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-022-01030-6.

Zhu:2022:MBP

[368] Guanghu Zhu, Yuanyuan Shi, and Qihuai Liu. Model-based projection of Zika infection risk with temperature effect: a case study in Southeast Asia. Bulletin of Mathematical Biology, 84(9):??, September 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-022-01049-9.

Forien:2022:ALM

[369] Raphaël Forien, Jimmy Garnier, and Florian Patout. Ancestral lineages in mutation selection equilibria with moving optimum. Bulletin of Mathematical Biology, 84(9):??, September 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01048-w.

Hu:2022:AGT

[370] Yi Hu, Kaifa Wang, and Wendi Wang. Analysis of the geographic transmission differences of COVID-19 in China caused by population movement and population density. Bulletin of Mathematical Biology, 84 (9):??, September 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01050-2.

Su:2022:EWD

[371] Ying Su, Bo Zheng, and Xingfu Zou. Wolbachia dynamics in mosquitoes with incomplete CI and imperfect maternal transmission by a DDE system. Bulletin of Mathematical Biology, 84(9):??, September 2022. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01042-2.

Brouwer:2022:WSR

[372] Andrew F. Brouwer. Why the Spectral Radius? An intuition-building introduction to the basic reproduction number. Bulletin of Mathematical Biology, 84(9):??, September 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-01057-9.

Terefe:2022:IWI

[373] Y. A. Terefe, S. M. Kassa, and J. B. H. Njagarah. Impact of the WHO integrated stewardship policy on the control of methicillin-resistant Staphyloccus aureus and third-generation cephalosporin-resistant Escherichia coli: Using a Mathematical Modeling Approach. Bulletin of Mathematical Biology, 84(9):??, September 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-01051-1.

Sargood:2022:FDG

[374] Alec Sargood, Eamonn A. Gaffney, and Andrew L. Krause. Fixed and distributed gene expression time delays in reaction–diffusion systems. Bulletin of Mathematical Biology, 84(9):??, September 2022. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-022-01052-0.

Wang:2022:MMC

[375] Xueying Wang, Sunpeng Wang, and Libin Rong. A multiscale model of COVID-19 dynamics. Bulletin of Mathematical Biology, 84(9):??, September 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
022-01058-8.

Cisneros-Ake:2022:TIR

[376] Luis A. Cisneros-Ake, Juan C. Gonzalez-Rodriguez, and Laura R. González-Ramírez. Turing instabilities and rotating spiral waves in glycolytic processes. Bulletin of Mathematical Biology, 84(9):??, September 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
022-01060-0.

Woolley:2022:BCC

[377] Thomas E. Woolley. Boundary conditions cause different generic bifurcation structures in Turing systems. Bulletin of Mathematical Biology, 84(9):??, September 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01055-x.

Arroyo-Esquivel:2022:CLT

[378] Jorge Arroyo-Esquivel, Alan Hastings, and Marissa L. Baskett. Characterizing long transients in consumer–resource systems with group defense and discrete reproductive pulses. Bulletin of Mathematical Biology, 84(9):??, September 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01059-7.

Rajakaruna:2022:MMG

[379] Harshana Rajakaruna and Vitaly V. Ganusov. Mathematical modeling to guide experimental design: T cell clustering as a case study. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01063-x.

Yang:2022:IGM

[380] Renlong Yang, Chongming Jiang, and Yuanzhi Shao. Invasiveness of a growth-migration system in a two-dimensional percolation cluster: a stochastic mathematical approach. Bulletin of Mathematical Biology, 84 (10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01056-w.

Zhang:2022:AET

[381] Dongni Zhang and Tom Britton. Analysing the effect of test-and-trace strategy in an SIR epidemic model. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01065-9.

Tang:2022:CMC

[382] Biao Tang, Weike Zhou, and Yanni Xiao. Controlling multiple COVID-19 epidemic waves: an insight from a multi-scale model linking the behaviour change dynamics to the disease transmission dynamics. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01061-z.

Craft:2022:EBN

[383] Michelle Craft and Cheng Ly. The effects of background noise on a biophysical model of olfactory bulb mitral cells. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-01066-8.

Wang:2022:AAS

[384] Xia Wang, Hulin Wu, and Sanyi Tang. Assessing age-specific vaccination strategies and post-vaccination reopening policies for COVID-19 control using SEIR modeling approach. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01064-w.

Cooney:2022:PMP

[385] Daniel B. Cooney, Fernando W. Rossine, and Simon A. Levin. A PDE model for protocell evolution and the origin of chromosomes via multilevel selection. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01062-y.

Dali-Youcef:2022:PST

[386] Manel Dali-Youcef, Alain Rapaport, and Tewfik Sari. Performance study of two serial interconnected chemostats with mortality. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01068-6.

Chang:2022:FPP

[387] Youngseok Chang, Wonhyung Choi, and Inkyung Ahn. On the fitness of predators with prey-induced dispersal in a habitat with spatial heterogeneity. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01069-5.

Bitsouni:2022:MSR

[388] Vasiliki Bitsouni, Nikolaos Gialelis, and Vasilis Tsilidis. A mathematical study of the role of tBregs in breast cancer. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01054-y.

Yasuda:2022:NSV

[389] Shugo Yasuda. Numerical study of the volcano effect in chemotactic aggregation based on a kinetic transport equation with non-instantaneous tumbling. Bulletin of Mathematical Biology, 84(10):??, October 2022. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01071-x.

Buttenschon:2022:CRB

[390] Andreas Buttenschön and Leah Edelstein-Keshet. Cell repolarization: a bifurcation study of spatio-temporal perturbations of polar cells. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-022-01053-z.

Mondal:2022:CSP

[391] Sudeshna Mondal, Guruprasad Samanta, and Manuel De la Sen. A comparison study of predator–prey model in deterministic and stochastic environments with the impacts of fear and habitat complexity. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01067-7.

Masud:2022:URE

[392] M. A. Masud, Md. Hamidul Islam, and Byul Nim Kim. Understanding the role of environmental transmission on COVID-19 herd immunity and invasion potential. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01070-y.

Wu:2022:AAO

[393] Xiaotian Wu, Hao Zhang, and Jun Li. An analytical approach of onecompartmental pharmacokinetic models with sigmoidal hill elimination. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-022-01078-4.

Soewongsono:2022:SPU

[394] Albert Ch. Soewongsono, Barbara R. Holland, and Małgorzata M. O’Reilly. The shape of phylogenies under phase-type distributed times to speciation and extinction. Bulletin of Mathematical Biology, 84 (10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01072-w.

Huber:2022:FBN

[395] K. T. Huber, V. Moulton, and G. E. Scholz. Forest-based networks. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-022-01081-9.

Zahan:2022:MSR

[396] Ishrat Zahan, Md. Kamrujjaman, and Saleh Tanveer. Mathematical study of a resource-based diffusion model with Gilpin–Ayala growth and harvesting. Bulletin of Mathematical Biology, 84(10):??, October 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-022-01074-8.

Liu:2022:MPS

[397] Yunfeng Liu, Jianshe Yu, and Jia Li. A mosquito population suppression model by releasing Wolbachia-infected males. Bulletin of Mathematical Biology, 84(11):??, November 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-01073-9.

Treesatayapun:2022:MDO

[398] Chidentree Treesatayapun. Model dynamics and optimal control for intervention policy of COVID-19 epidemic with quarantine and immigrating disturbances. Bulletin of Mathematical Biology, 84(11):??, November 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
022-01080-w.

Senneff:2022:CME

[399] Sageanne Senneff and Madeleine M. Lowery. Computational model of the effect of mitochondrial dysfunction on excitation–contraction coupling in skeletal muscle. Bulletin of Mathematical Biology, 84(11):??, November 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-

022-01079-3. See correction [410].

Sartori:2022:TSA

[400] Larissa Sartori, Marcone Pereira, and Sergio Oliva. Time-scale analysis and parameter fitting for vector-borne diseases with spatial dynamics. Bulletin of Mathematical Biology, 84(11):??, November 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-022-01083-7.

Willson:2022:DCT

[401] Stephen J. Willson. Distinct-cluster tree-child phylogenetic networks and possible uses to study polyploidy. Bulletin of Mathematical Biology, 84 (11):??, November 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01084-6.

Cooney:2022:ARM

[402] Daniel B. Cooney. Assortment and reciprocity mechanisms for promotion of cooperation in a model of multilevel selection. Bulletin of Mathematical Biology, 84(11):??, November 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01082-8.

Catano-Lopez:2022:DME

[403] Alexandra Catano-Lopez, Daniel Rojas-Diaz, and María Eugenia Puerta Yepes. Discrete models in epidemiology: New contagion probability functions based on real data behavior. Bulletin of Mathematical Biology, 84 (11):??, November 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01076-6.

Arino:2022:EME

[404] Julien Arino and Evan Milliken. Effect of movement on the early phase of an epidemic. Bulletin of Mathematical Biology, 84(11):??, November 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01077-5.

Lubenia:2022:RNA

[405] Patrick Vincent N. Lubenia, Eduardo R. Mendoza, and Angelyn R. Lao. Reaction network analysis of metabolic insulin signaling. Bulletin of Mathematical Biology, 84(11):??, November 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01087-3.

ElWajeh:2022:CKM

[406] Mohammad El Wajeh, Falco Jung, and Jakob Nikolas Kather. Can the Kuznetsov model replicate and predict cancer growth in humans? Bulletin of Mathematical Biology, 84(11):??, November 2022. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-022-01075-7.

Zheng:2022:ERN

[407] Qi Zheng. Estimation of rates of non-neutral mutations when bacteria are exposed to subinhibitory levels of antibiotics. Bulletin of Mathematical Biology, 84(11):??, November 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-01085-5.

Xue:2022:LTF

[408] Ling Xue, Kai Zhang, and Hao Wang. Long-term forecast of HIV/AIDS epidemic in China with fear effect and 90–90–90 strategies. Bulletin of Mathematical Biology, 84(11):??, November 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/article/10.1007/s11538-022-01091-7.

Wangen:2022:OMS

[409] Catherine E. Wangen, James A. Powell, and Barbara J. Bentz. Oviposition model for a southern population of mountain pine beetle. Bulletin of Mathematical Biology, 84(11):??, November 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01089-1.

Senneff:2022:CCM

[410] Sageanne Senneff and Madeleine M. Lowery. Correction to: Computational model of the effect of mitochondrial dysfunction on excitation– contraction coupling in skeletal muscle. Bulletin of Mathematical Biology, 84(11):??, November 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01093-5. See [399].

Wang:2022:SSC

[411] Minglong Wang, Yuxiang Zhang, and Qihua Huang. A stage-structured continuous-/discrete-time population model: Persistence and spatial spread. Bulletin of Mathematical Biology, 84(11):??, November 2022. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01090-8.

Chapwanya:2022:AWC

[412] Michael Chapwanya, Jean Lubuma, Yibeltal Terefe, and Berge Tsanou. Analysis of war and conflict effect on the transmission dynamics of the tenth Ebola outbreak in the Democratic Republic of Congo. Bulletin of Mathematical Biology, 84(12):??, December 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/article/10.1007/s11538-022-01094-4.

Marsh:2022:AGT

[413] Lewis Marsh, Emilie Dufresne, Helen M. Byrne, and Heather A. Harrington. Algebra, geometry and topology of ERK kinetics. Bulletin of Mathematical Biology, 84(12):??, December 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01088-2.

Hamdan:2022:MMD

[414] Nur ’Izzati Hamdan and Adem Kilicman. Mathematical modelling of dengue transmission with intervention strategies using fractional derivatives. Bulletin of Mathematical Biology, 84(12):??, December 2022. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01096-2.

Shyntar:2022:TIP

[415] Alexandra Shyntar, Ashna Patel, Meghan Rhodes, Heiko Enderling, and Thomas Hillen. The tumor invasion paradox in cancer stem cell-driven solid tumors. Bulletin of Mathematical Biology, 84(12):??, December 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01086-4.

Wang:2022:BPF

[416] Xiaoli Wang, Junping Shi, and Guohong Zhang. Bifurcation and pattern formation in an activator–inhibitor model with non-local dispersal. Bulletin of Mathematical Biology, 84(12):??, December 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-022-01098-0.

Almeida:2022:HDC

[417] Luis Almeida, Chloe Audebert, Emma Leschiera, and Tommaso Lorenzi. A hybrid discrete–continuum modelling approach to explore the impact of T-cell infiltration on anti-tumour immune response. Bulletin of Mathematical Biology, 84(12):??, December 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01095-3.

Giantesio:2022:MDF

[418] Giulia Giantesio, Alberto Girelli, and Alessandro Musesti. A mathematical description of the flow in a spherical lymph node. Bulletin of Mathematical Biology, 84(12):??, December 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01103-6.

Macfarlane:2022:IPH

[419] Fiona R. Macfarlane, Tommaso Lorenzi, and Kevin J. Painter. The impact of phenotypic heterogeneity on chemotactic self-organisation. Bulletin of Mathematical Biology, 84(12):??, December 2022. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-022-01099-z.

Azizi:2022:EHB

[420] Asma Azizi, Caner Kazanci, Natalia L. Komarova, and Dominik Wodarz. Effect of human behavior on the evolution of viral strains during an epidemic. Bulletin of Mathematical Biology, 84(12):??, December 2022. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01102-7.

Gani:2022:IIW

[421] M. Osman Gani, M. Humayun Kabir, and Toshiyuki Ogawa. Inhibitorinduced wavetrains and spiral waves in an extended FitzHugh–Nagumo model of nerve cell dynamics. Bulletin of Mathematical Biology, 84 (12):??, December 2022. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01100-9.

Li:2022:ACE

[422] Zhimin Li and Tailei Zhang. Analysis of a COVID-19 epidemic model with seasonality. Bulletin of Mathematical Biology, 84(12):??, December 2022. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01105-4.

Andrade:2023:HBM

[423] Renato Andrade and Christina A. Cobbold. Heterogeneity in behaviour and movement can influence the stability of predator–prey periodic travelling waves. Bulletin of Mathematical Biology, 85(1):??, January 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01101-8.

Zhang:2023:CBI

[424] Shengqiang Zhang, Xichao Duan, Tonghua Zhang, and Sanling Yuan. Controlling biological invasions: a stochastic host–generalist parasitoid model. Bulletin of Mathematical Biology, 85(1):??, January 2023. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01106-3.

Rapaport:2023:RPR

[425] Alain Rapaport and Ismail Mimouni. The role of permanently resident populations in the two-patches SIR model with commuters. Bulletin of Mathematical Biology, 85(1):??, January 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01111-6.

Armbruster:2023:ESC

[426] Ashley Armbruster, Matt Holzer, Noah Roselli, and Lena Underwood. Epidemic spreading on complex networks as front propagation into an unstable state. Bulletin of Mathematical Biology, 85(1):??, January 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01110-7.

Shao:2023:CMS

[427] Sijing Shao, Judith E. Canner, Rebecca A. Everett, Kidist Bekele-Maxwell, Alexis Kuerbis, Lyric Stephenson, Jennifer Menda, Jon Morgenstern, and H. T. Banks. A comparison of mathematical and statistical modeling with longitudinal data: an application to ecological momentary assessment of behavior change in individuals with alcohol use disorder. Bulletin of Mathematical Biology, 85(1):??, January 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-022-01097-1.

Xue:2023:CWH

[428] Yuyi Xue, Daipeng Chen, Stacey R. Smith, Xiaoe Ruan, and Sanyi Tang. Coupling the within-host process and between-host transmission of COVID-19 suggests vaccination and school closures are critical. Bulletin of Mathematical Biology, 85(1):??, January 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01104-5. See correction [448].

Davenport:2023:MMT

[429] Angelica A. Davenport, Yun Lu, Carlos A. Gallegos, Adriana V. F. Massicano, Katherine A. Heinzman, Patrick N. Song, Anna G. Sorace, and N. G. Cogan. Mathematical model of triple-negative breast cancer in response to combination chemotherapies. Bulletin of Mathematical Biology, 85(1):??, January 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01108-1.

Ortega-Sabater:2023:SFD

[430] Carmen Ortega-Sabater, Gabriel F. Calvo, Jelena Dinić, Ana Podolski, Milica Pesic, and Víctor Pérez-García. Stochastic fluctuations drive nongenetic evolution of proliferation in clonal cancer cell populations. Bulletin of Mathematical Biology, 85(1):??, January 2023. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-022-01113-4.

Assad:2023:CST

[431] Daniel Bouzon Nagem Assad, Javier Cara, and Miguel Ortega-Mier. Comparing short-term univariate and multivariate time-series forecasting models in infectious disease outbreak. Bulletin of Mathematical Biology, 85(1):??, January 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01112-5.

Kazmierczak:2023:EBM

[432] Bogdan Kazmierczak, James Sneyd, and Je-Chiang Tsai. Effect of buffers with multiple binding sites on calcium waves. Bulletin of Mathematical Biology, 85(1):??, January 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-022-01109-0.

Wang:2023:MAE

[433] Yi Wang, Jinde Cao, Changfeng Xue, and Li Li. Mathematical analysis of epidemic models with treatment in heterogeneous networks. Bulletin of Mathematical Biology, 85(2):??, February 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01116-1.

Siewe:2023:BCE

[434] Nourridine Siewe and Avner Friedman. Breast cancer exosomal microR-NAs facilitate pre-metastatic niche formation in the bone: a mathematical model. Bulletin of Mathematical Biology, 85(2):??, February 2023. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-022-01117-0.

Zhao:2023:MGS

[435] Lihong Zhao, Fabian Santiago, Erica M. Rutter, Shilpa Khatri, and Suzanne S. Sindi. Modeling and global sensitivity analysis of strategies to mitigate Covid-19 transmission on a structured college campus. Bulletin of Mathematical Biology, 85(2):??, February 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-022-01107-2.

Krause:2023:CDD

[436] Andrew L. Krause, Eamonn A. Gaffney, and Benjamin J. Walker. Concentration-dependent domain evolution in reaction–diffusion systems. Bulletin of Mathematical Biology, 85(2):??, February 2023. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-022-01115-2.

Penn:2023:AAO

[437] Matthew J. Penn and Christl A. Donnelly. Asymptotic analysis of optimal vaccination policies. Bulletin of Mathematical Biology, 85(3):??, March 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
022-01114-3.

Singh:2023:SMT

[438] Prerna Singh and Alex Best. A sterility–mortality tolerance trade-off leads to within-population variation in host tolerance. Bulletin of Mathematical Biology, 85(3):??, March 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01119-6.

Saona:2023:CRB

[439] Raimundo Saona, Fyodor A. Kondrashov, and Ksenia A. Khudiakova. Correction to: Relation between the number of peaks and the number of reciprocal sign epistatic interactions. Bulletin of Mathematical Biology, 85(3):??, March 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-022-01118-z. See [350].

Kiss:2023:PIN

[440] István Z. Kiss and Péter L. Simon. On parameter identifiability in network–based epidemic models. Bulletin of Mathematical Biology, 85 (3):??, March 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01121-y.

Stevenson:2023:ERP

[441] Joshua Stevenson, Barbara Holland, Michael Charleston, and Jeremy Sumner. Evaluation of the relative performance of the subflattenings method for phylogenetic inference. Bulletin of Mathematical Biology, 85(3):??, March 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01120-z.

Wang:2023:MSH

[442] Cheng-Long Wang, Shasha Gao, Xue-Zhi Li, and Maia Martcheva. Modeling syphilis and HIV coinfection: a case study in the USA. Bulletin of Mathematical Biology, 85(3):??, March 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01123-w.

Chen:2023:RSS

[443] Ricky X. F. Chen, Christian M. Reidys, and Michael S. Waterman. RNA secondary structures with given motif specification: Combinatorics and algorithms. Bulletin of Mathematical Biology, 85(3):??, March 2023. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01128-5.

Scacchi:2023:EDR

[444] Simone Scacchi, Piero Colli Franzone, Luca F. Pavarino, Vincenzo Gionti, and Cesare Storti. Epicardial dispersion of repolarization promotes the onset of reentry in Brugada Syndrome: a numerical simulation study. Bulletin of Mathematical Biology, 85(3):??, March 2023. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-023-01124-9.

Agnelli:2023:SKM

[445] Juan Pablo Agnelli, Bruno Buffa, Damián Knopoff, and Germán Torres. A spatial kinetic model of crowd evacuation dynamics with infectious disease contagion. Bulletin of Mathematical Biology, 85(4):??, April 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01127-6.

Michel:2023:RFR

[446] Christian J. Michel and Jean-Sébastien Sereni. Reading frame retrieval of genes: a new parameter of codon usage based on the circular code theory. Bulletin of Mathematical Biology, 85(4):??, April 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-023-01129-4.

Annunziata:2023:EEC

[447] Carlyn Annunziata, Haniyeh Fattahpour, Daniel Fong, Michael Hadjiargyrou, and Pejman Sanaei. Effects of elasticity on cell proliferation in a tissue-engineering scaffold pore. Bulletin of Mathematical Biology, 85(4):??, April 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01134-7.

Xue:2023:CCW

[448] Yuyi Xue, Daipeng Chen, Stacey R. Smith, Xiaoe Ruan, and Sanyi Tang. Correction to: Coupling the within–host process and between–host transmission of COVID-19 suggests vaccination and school closures are critical. Bulletin of Mathematical Biology, 85(4):??, April 2023. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-023-01132-9. See [428].

Alberding:2023:SAT

[449] Jonathan P. Alberding and Timothy W. Secomb. Simulation of angiogenesis in three dimensions: Development of the retinal circulation. Bulletin of Mathematical Biology, 85(4):??, April 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01126-7.

Duan:2023:CDH

[450] Xi-Chao Duan, Jiangyue Zhao, and Maia Martcheva. Coevolutionary dynamics of host immune and parasite virulence based on an age-structured epidemic model. Bulletin of Mathematical Biology, 85(4):??, April 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01131-w.

Deal:2023:BMT

[451] Isadora Deal, Matthew Macauley, and Robin Davies. Boolean models of the transport, synthesis, and metabolism of tryptophan in Escherichia coli. Bulletin of Mathematical Biology, 85(4):??, April 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-023-01122-x.

ONeill:2023:IHA

[452] Xander O’Neill, Andy White, Christian Gortázar, and Francisco Ruiz-Fons. The impact of host abundance on the epidemiology of tick-borne infection. Bulletin of Mathematical Biology, 85(4):??, April 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-023-01133-8.

Igoe:2023:ROC

[453] Morganne Igoe, Renato Casagrandi, Marino Gatto, Christopher M. Hoover, Lorenzo Mari, Calistus N. Ngonghala, Justin V. Remais, James N. Sanchirico, Susanne H. Sokolow, Suzanne Lenhart, and Giulio de Leo. Reframing optimal control problems for infectious disease management in low-income countries. Bulletin of Mathematical Biology, 85(4):??, April 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01137-4.

Betti:2023:CVH

[454] Matthew I. Betti, Amira Hassan Abouleish, Victoria Spofford, Cory Peddigrew, Alan Diener, and Jane M. Heffernan. COVID-19 vaccination and healthcare demand. Bulletin of Mathematical Biology, 85(5):??, May 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01130-x.

Bardon:2023:ESS

[455] Gaël Bardon and Frédéric Barraquand. Effects of stage structure on coexistence: Mixed benefits. Bulletin of Mathematical Biology, 85(5):??, May 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01135-6.

Lee:2023:MCM

[456] Tae Jin Lee, Adam E. Berman, and Arni S. R. Srinivasa Rao. Markov chain models for cardiac rhythm dynamics in patients undergoing catheter ablation of atrial fibrillation. Bulletin of Mathematical Biology, 85(5):??, May 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01125-8.

Kim:2023:NDB

[457] Yong-Jung Kim, Masayasu Mimura, and Changwook Yoon. Nonlinear diffusion for bacterial traveling wave phenomenon. Bulletin of Mathematical Biology, 85(5):??, May 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01138-3.

Zanella:2023:KME

[458] Mattia Zanella. Kinetic models for epidemic dynamics in the presence of opinion polarization. Bulletin of Mathematical Biology, 85(5):??, May 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01147-2.

Ahmed:2023:MAI

[459] Ishraq U. Ahmed, Helen M. Byrne, and Mary R. Myerscough. Macrophage anti-inflammatory behaviour in a multiphase model of atherosclerotic plaque development. Bulletin of Mathematical Biology, 85(5):??, May 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01142-7.

Walker:2023:MMM

[460] Benjamin J. Walker, Giulia L. Celora, Alain Goriely, Derek E. Moulton, and Helen M. Byrne. Minimal morphoelastic models of solid tumour spheroids: a tutorial. Bulletin of Mathematical Biology, 85(5):??, May 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01141-8.

Campos:2023:CTS

[461] Felipe A. Campos, Simone Bruno, Yi Fu, Domitilla Del Vecchio, and Ruth J. Williams. Comparison theorems for stochastic chemical reaction networks. Bulletin of Mathematical Biology, 85(5):??, May 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-023-01136-5.

Huber:2023:AAP

[462] K. T. Huber and L. J. Maher. Autopolyploidy, allopolyploidy, and phylogenetic networks with horizontal arcs. Bulletin of Mathematical Biology, 85(5):??, May 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01140-9.

Wu:2023:PIM

[463] Peng Wu, Shohel Ahmed, Xiunan Wang, and Hao Wang. PrEP intervention in the mitigation of HIV/AIDS epidemics in China via a data–validated age-structured model. Bulletin of Mathematical Biology, 85(5):??, May 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01145-4.

Buckingham:2023:EAO

[464] Lydia J. Buckingham and Ben Ashby. The evolution of the age of onset of resistance to infectious disease. Bulletin of Mathematical Biology, 85(5):??, May 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01144-5.

Anwar:2023:OIE

[465] Md Nurul Anwar, Roslyn I. Hickson, Somya Mehra, David J. Price, James M. McCaw, Mark B. Flegg, and Jennifer A. Flegg. Optimal interruption of P. vivax malaria transmission using mass drug administration. Bulletin of Mathematical Biology, 85(6):??, June 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-023-01153-4.

Fritz:2023:TEM

[466] Marvin Fritz. Tumor evolution models of phase-field type with nonlocal effects and angiogenesis. Bulletin of Mathematical Biology, 85(6):??, June 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01151-6.

Butler:2023:MOA

[467] Cole Butler and Peter Stechlinski. Modeling opioid abuse: a case study of the opioid crisis in New England. Bulletin of Mathematical Biology, 85(6):??, June 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01148-1.

Francis:2023:LPN

[468] Andrew Francis and Mike Steel. Labellable phylogenetic networks. Bulletin of Mathematical Biology, 85(6):??, June 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/article/10.1007/s11538-023-01157-0.

Wilson:2023:MRC

[469] Nicole Wilson, Corina S. Drapaca, Heiko Enderling, Jimmy J. Caudell, and Kathleen P. Wilkie. Modelling radiation cancer treatment with a death-rate term in ordinary and fractional differential equations. Bulletin of Mathematical Biology, 85(6):??, June 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01139-2.

Eilertsen:2023:NPC

[470] Justin Eilertsen, Santiago Schnell, and Sebastian Walcher. Natural parameter conditions for singular perturbations of chemical and biochemical reaction networks. Bulletin of Mathematical Biology, 85(6):??, June 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01150-7.

Mazuryn:2023:MFG

[471] Maksim Mazuryn and Uffe Høgsbro Thygesen. Mean field games for diel vertical migration with diffusion. Bulletin of Mathematical Biology, 85(6):??, June 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01154-3.

Amiri:2023:TOA

[472] Farhad A. Amiri and Junfeng Zhang. Tissue oxygenation around capillaries: Effects of hematocrit and arteriole oxygen condition. Bulletin of Mathematical Biology, 85(6):??, June 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01155-2.

Zhang:2023:CMD

[473] Yuyuan Zhang, Kaiqun Wang, Yaoyao Du, Huiyuan Yang, Guanjie Jia, Di Huang, Weiyi Chen, and Yanhu Shan. Computational modeling to determine the effect of phenotypic heterogeneity in tumors on the collective tumor-immune interactions. Bulletin of Mathematical Biology, 85(6):??, June 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01158-z.

Barril:2023:FSE

[474] Carles Barril, Pierre-Alexandre Bliman, and Sílvia Cuadrado. Final size for epidemic models with asymptomatic transmission. Bulletin of Mathematical Biology, 85(6):??, June 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01159-y.

Garcia:2023:DVS

[475] Pablo Venegas Garcia and Hao Wang. A data-validated stoichiometric model for the priming effect. Bulletin of Mathematical Biology, 85(6):??, June 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01160-5.

Rapti:2023:RMD

[476] Z. Rapti, J. Cuevas-Maraver, E. Kontou, S. Liu, Y. Drossinos, P. G. Kevrekidis, M. Barmann, Q.-Y. Chen, and G. A. Kevrekidis. The role of mobility in the dynamics of the COVID-19 epidemic in Andalusia. Bulletin of Mathematical Biology, 85(6):??, June 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01152-5.

Maity:2023:MBE

[477] Biplab Maity, Bapi Saha, Indrajit Ghosh, and Joydev Chattopadhyay. Model-based estimation of expected time to cholera extinction in Lusaka, Zambia. Bulletin of Mathematical Biology, 85(7):??, July 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-023-01149-0.

Torres:2023:QAD

[478] Monica Torres, Jerrold Tubay, and Aurelio de losReyes V. Quantitative assessment of a dual epidemic caused by tuberculosis and HIV in the Philippines. Bulletin of Mathematical Biology, 85(7):??, July 2023. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01156-1.

Murray:2023:ATT

[479] Philip J. Murray. Autoregulation of transcription and translation: a qualitative analysis. Bulletin of Mathematical Biology, 85(7):??, July 2023.

CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01143-6.

Salentine:2023:MMD

[480] Noah Salentine, Jonathan Doria, Chinh Nguyen, Gabriella Pinter, Shizhen Emily Wang, and Peter Hinow. A mathematical model of the disruption of glucose homeostasis in cancer patients. Bulletin of Mathematical Biology, 85(7):??, July 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01146-3.

Zhao:2023:EEP

[481] Xin Zhao, Lidan Liu, Hao Wang, and Meng Fan. Ecological effects of predator harvesting and environmental noises on oceanic coral reefs. Bulletin of Mathematical Biology, 85(7):??, July 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/article/10.1007/s11538-023-01166-z.

Loy:2023:SMA

[482] N. Loy and L. Preziosi. A statistical mechanics approach to describe cell reorientation under stretch. Bulletin of Mathematical Biology, 85(7):??, July 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01161-4.

He:2023:RDM

[483] Changhan He, Lifeng Han, Duane C. Harris, Samat Bayakhmetov, Xiao Wang, and Yang Kuang. Reaction–diffusion modeling of E. coli colony growth based on nutrient distribution and agar dehydration. Bulletin of Mathematical Biology, 85(7):??, July 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01163-2.

Nguyen:2023:EPD

[484] Kyle Nguyen, Erica M. Rutter, and Kevin B. Flores. Estimation of parameter distributions for reaction–diffusion equations with competition using aggregate spatiotemporal data. Bulletin of Mathematical Biology, 85(7):??, July 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01162-3.

Russo:2023:MSM

[485] Fabiana Russo, Alberto Tenore, Maria Rosaria Mattei, and Luigi Frunzo. A mathematical study of metal biosorption on algal–bacterial granular biofilms. Bulletin of Mathematical Biology, 85(7):??, July 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-023-01168-x.

Duan:2023:RMD

[486] Xiaoyu Duan, Jonathan E. Rubin, and David Swigon. Rigorous mapping of data to qualitative properties of parameter values and dynamics: a case study on a two-variable Lotka–Volterra system. Bulletin of Mathematical Biology, 85(7):??, July 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01165-0.

Yan:2023:SAT

[487] Dingding Yan, Xiaotian Wu, Jun Li, and Sanyi Tang. Statistical analysis of two-compartment pharmacokinetic models with drug non-adherence. Bulletin of Mathematical Biology, 85(7):??, July 2023. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-023-01173-0.

Palma:2023:EMI

[488] Giuseppe Palma, Damiano Caprioli, and Lorenzo Mari. Epidemic management via imperfect testing: a multi-criterial perspective. Bulletin of Mathematical Biology, 85(7):??, July 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01172-1.

Abboud:2023:FPD

[489] Candy Abboud, Eric Parent, Olivier Bonnefon, and Samuel Soubeyrand. Forecasting pathogen dynamics with Bayesian model–averaging: Application to Xylella fastidiosa. Bulletin of Mathematical Biology, 85(7):??, July 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01169-w.

Casanellas:2023:DWQ

[490] Marta Casanellas, Jesús Fernández-Sánchez, Marina Garrote-López, and Marc Sabaté-Vidales. Designing weights for quartet-based methods when data are heterogeneous across lineages. Bulletin of Mathematical Biology, 85(7):??, July 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01167-y.

Gladson:2023:ABM

[491] Shane L. Gladson and Tracy L. Stepien. An agent-based model of biting midge dynamics to understand bluetongue outbreaks. Bulletin of Mathematical Biology, 85(7):??, July 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01177-w.

Hamster:2023:WSC

[492] Christian Hamster and Peter van Heijster. Waves in a stochastic cell motility model. Bulletin of Mathematical Biology, 85(8):??, August 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01164-1.

Saha:2023:SSD

[493] Esha Saha, Lam Si Tung Ho, and Giang Tran. SPADE4: Sparsity and delay embedding based forecasting of epidemics. Bulletin of Mathematical Biology, 85(8):??, August 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01174-z.

Laruelle:2023:EHC

[494] Annick Laruelle, André Rocha, Claudia Manini, José I. López, and Elena Inarra. Effects of heterogeneity on cancer: A game theory perspective. Bulletin of Mathematical Biology, 85(8):??, August 2023. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-023-01178-9.

Penn:2023:OME

[495] Matthew J. Penn and Christl A. Donnelly. Optimality of maximal–effort vaccination. Bulletin of Mathematical Biology, 85(8):??, August 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01179-8.

Colson:2023:IIG

[496] Chloé Colson, Philip K. Maini, and Helen M. Byrne. Investigating the influence of growth arrest mechanisms on tumour responses to radiotherapy. Bulletin of Mathematical Biology, 85(8):??, August 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-023-01171-2.

Weatherley:2023:CMK

[497] Georgia Weatherley, Robyn P. Araujo, Samantha J. Dando, and Adrianne L. Jenner. Could mathematics be the key to unlocking the mysteries of multiple sclerosis? Bulletin of Mathematical Biology, 85(8):??, August 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01181-0.

Erban:2023:CSL

[498] Radek Erban and Hye-Won Kang. Chemical systems with limit cycles. Bulletin of Mathematical Biology, 85(8):??, August 2023. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-023-01170-3.

Kravtsova:2023:SGW

[499] Natalia Kravtsova, Reginald L. McGee II, and Adriana T. Dawes. Scalable Gromov–Wasserstein based comparison of biological time series. Bulletin of Mathematical Biology, 85(8):??, August 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01175-y.

Manson:2023:SPD

[500] Kerry Manson and Mike Steel. Spaces of phylogenetic diversity indices: Combinatorial and geometric properties. Bulletin of Mathematical Biology, 85(8):??, August 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01183-y.

Colombi:2023:MCO

[501] Annachiara Colombi, Luigi Preziosi, and Marco Scianna. Modelling cell orientation under stretch: The effect of substrate elasticity. Bulletin of Mathematical Biology, 85(9):??, September 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/article/10.1007/s11538-023-01180-1.

Faye:2023:MDW

[502] Grégory Faye, Guilhem Fouilhé, and Rufin VanRullen. Mathematical derivation of wave propagation properties in hierarchical neural networks with predictive coding feedback dynamics. Bulletin of Mathematical Biology, 85(9):??, September 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01186-9.

Hayashi:2023:TPE

[503] Rena Hayashi and Yoh Iwasa. Temporal pattern of the emergence of a mutant virus escaping cross-immunity and stochastic extinction within a host. Bulletin of Mathematical Biology, 85(9):??, September 2023. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01184-x.

Elias:2023:PRT

[504] Ján Eliaš, Klemens Fellner, Peter Hofer, Monika Oberer, Renate Schreiber, and Rudolf Zechner. The potential roles of transacylation in intracellular lipolysis and related qssa approximations. Bulletin of Mathematical Biology, 85(9):??, September 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01188-7.

DeJaco:2023:RBQ

[505] Robert F. DeJaco, Matthew J. Roberts, Erica L. Romsos, Peter M. Vallone, and Anthony J. Kearsley. Reducing bias and quantifying uncertainty in fluorescence produced by PCR. Bulletin of Mathematical Biology, 85 (9):??, September 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01182-z.

Bate:2023:ELP

[506] Fiona Bate, Yumechris Amekan, Dmitri O. Pushkin, James P. J. Chong, and Martin Bees. Emergent lag phase in flux-regulation models of bacterial growth. Bulletin of Mathematical Biology, 85(9):??, September 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01189-6.

Watson:2023:LSM

[507] Michael G. Watson, Keith L. Chambers, and Mary R. Myerscough. A lipid-structured model of atherosclerotic plaque macrophages with lipid– dependent kinetics. Bulletin of Mathematical Biology, 85(9):??, September 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01193-w.

Gordillo:2023:IPD

[508] Luis F. Gordillo and Priscilla E. Greenwood. Intermittent precipitationdependent interactions, encompassing Allee effect, may yield vegetation patterns in a transitional parameter range. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01191-y.

Anderson:2023:SRN

[509] David F. Anderson and Aidan S. Howells. Stochastic reaction networks within interacting compartments. Bulletin of Mathematical Biology, 85 (10):??, October 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01185-w.

Giverso:2023:INM

[510] Chiara Giverso, Gaspard Jankowiak, Luigi Preziosi, and Christian Schmeiser. The influence of nucleus mechanics in modelling adhesionindependent cell migration in structured and confined environments. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-023-01187-8.

Plaugher:2023:PCT

[511] Daniel Plaugher and David Murrugarra. Phenotype control techniques for Boolean gene regulatory networks. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01197-6.

Cassidy:2023:CTM

[512] Tyler Cassidy. A continuation technique for maximum likelihood estimators in biological models. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01200-0.

Armstrong:2023:PBS

[513] Eve Armstrong. Predicting the behavior of sparsely–sampled systems across neurobiology and epidemiology. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01176-x.

Morselli:2023:ABC

[514] David Morselli, Marcello Edoardo Delitala, and Federico Frascoli. Agentbased and continuum models for spatial dynamics of infection by oncolytic viruses. Bulletin of Mathematical Biology, 85(10):??, October 2023. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01192-x.

Moffett:2023:MRI

[515] Alexander S. Moffett, Youyuan Deng, and Herbert Levine. Modeling the role of immune cell conversion in the tumor-immune microenvironment. Bulletin of Mathematical Biology, 85(10):??, October 2023. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01201-z.

Mulder:2023:PDT

[516] Willem H. Mulder. Probability distribution of tree age for the simple birth–death process, with applications to distributions of number of ancestral lineages and divergence times for pairs of taxa in a Yule tree. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-023-01196-7.

Montealegre-Mora:2023:PDG

[517] Felipe Montealegre-Mora, Marcus Lapeyrolerie, Melissa Chapman, Abigail G. Keller, and Carl Boettiger. Pretty darn good control: When are approximate solutions better than approximate models. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01198-5.

Dari:2023:MEM

[518] Sonia Dari, Nabil T. Fadai, and Reuben D. O’Dea. Modelling the effect of matrix metalloproteinases in dermal wound healing. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01195-8.

Liu:2023:GJR

[519] Kaihui Liu, Zhenguo Bai, Daihai He, and Yijun Lou. Getting jab or regular test: Observations from an impulsive epidemic COVID-19 model. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-023-01202-y.

Ayensa-Jimenez:2023:MMS

[520] Jacobo Ayensa-Jiménez, Mohamed H. Doweidar, Manuel Doblaré, and Eamonn A. Gaffney. A mathematical modelling study of chemotactic dynamics in cell cultures: The impact of spatio-temporal heterogeneity. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-023-01194-9.

Wang:2023:TOC

[521] Mengshou Wang, Gao Li, Liangrong Peng, and Liu Hong. Towards optimal control of amyloid fibrillation. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01205-9.

Vlad:2023:CRV

[522] A. I. Vlad, A. A. Romanyukha, and T. E. Sannikova. Circulation of respiratory viruses in the city: Towards an agent-based ecosystem model. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-023-01203-x.

Kiradjiev:2023:MAA

[523] K. B. Kiradjiev and L. R. Band. Multiscale asymptotic analysis reveals how cell growth and subcellular compartments affect tissue-scale hormone transport. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01199-4.

Bottcher:2023:MCP

[524] Lucas Böttcher, Sascha Wald, and Tom Chou. Mathematical characterization of private and public immune receptor sequences. Bulletin of Mathematical Biology, 85(10):??, October 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01190-z.

Tian:2023:SMC

[525] Tim Y. Y. Tian, Colin B. Macdonald, and Eric N. Cytrynbaum. A stochastic model of cortical microtubule anchoring and mechanics provides regulatory control of microtubule shape. Bulletin of Mathematical Biology, 85(11):??, November 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01211-x.

Bergland:2023:DPH

[526] Harald Bergland, Evgenii Burlakov, and John Wyller. The dynamics of pasture–herbivores–carnivores with sigmoidal density dependent harvesting. Bulletin of Mathematical Biology, 85(11):??, November 2023. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01210-y.

Barker:2023:MAT

[527] Colin T. Barker, Feng-Bin Wang, and Naveen K. Vaidya. Modeling antiretrovial treatment to mitigate HIV in the brain: Impact of the blood– brain barrier. Bulletin of Mathematical Biology, 85(11):??, November 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01204-w.

Simmons:2023:MDC

[528] Tyler Simmons and Doron Levy. Modeling the development of cellular exhaustion and tumor-immune stalemate. Bulletin of Mathematical Biology, 85(11):??, November 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01207-7.

Stevenson:2023:REC

[529] Joshua Stevenson, Venta Terauds, and Jeremy Sumner. Rearrangement events on circular genomes. Bulletin of Mathematical Biology, 85 (11):??, November 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01209-5.

Brush:2023:CMP

[530] Micah Brush and Mark A. Lewis. Coupling mountain pine beetle and forest population dynamics predicts transient outbreaks that are likely to increase in number with climate change. Bulletin of Mathematical Biology, 85(11):??, November 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01215-7.

Marasco:2023:AGL

[531] Addolorata Marasco, Emiliano Spera, Vittorio De Falco, Annalisa Iuorio, Carmen Alina Lupascu, Sergio Solinas, and Michele Migliore. An adaptive generalized leaky integrate-and-fire model for hippocampal CA1 pyramidal neurons and interneurons. Bulletin of Mathematical Biology, 85(11):??, November 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01206-8.

Bortz:2023:DEP

[532] David M. Bortz, Daniel A. Messenger, and Vanja Dukic. Direct estimation of parameters in ODE models using WENDy: Weak-form estimation of nonlinear dynamics. Bulletin of Mathematical Biology, 85(11):??, November 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01208-6.

Germano:2023:FIB

[533] Domenic P. J. Germano, Adriana Zanca, Stuart T. Johnston, Jennifer A. Flegg, and James M. Osborne. Free and interfacial boundaries in individual-based models of multicellular biological systems. Bulletin of Mathematical Biology, 85(11):??, November 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/article/10.1007/s11538-023-01214-8.

Elzinga:2023:GSH

[534] David C. Elzinga and W. Christopher Strickland. Generalized stressors on hive and forager bee colonies. Bulletin of Mathematical Biology, 85 (11):??, November 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01219-3.

Walker:2023:VRI

[535] Benjamin J. Walker, Adam K. Townsend, Alexander K. Chudasama, and Andrew L. Krause. VisualPDE: Rapid interactive simulations of partial differential equations. Bulletin of Mathematical Biology, 85(11): ??, November 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01218-4.

Gorin:2023:AMD

[536] Gennady Gorin, Shawn Yoshida, and Lior Pachter. Assessing Markovian and delay models for single-nucleus RNA sequencing. Bulletin of Mathematical Biology, 85(11):??, November 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01213-9.

Liceaga:2023:EVE

[537] Daniel Olmos Liceaga, Sandro Filipe Nunes, and Roberto A. Saenz. Ex vivo experiments shed light on the innate immune response from influenza virus. Bulletin of Mathematical Biology, 85(11):??, November 2023. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01217-5.

Enkhtaivan:2023:PAA

[538] Enkhzaya Enkhtaivan, Joel Nishimura, and Amy Cochran. Placing approach-avoidance conflict within the framework of multi-objective reinforcement learning. Bulletin of Mathematical Biology, 85(11):??, November 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01216-6.

Eftimie:2023:MKD

[539] R. Eftimie, G. Rolin, O. E. Adebayo, S. Urcun, F. Chouly, and S. P. A. Bordas. Modelling keloids dynamics: a brief review and new mathematical perspectives. Bulletin of Mathematical Biology, 85(12):??, December 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01222-8.

Boudreau:2023:TPC

[540] Mariah C. Boudreau, Andrea J. Allen, Nicholas J. Roberts, Antoine Allard, and Laurent Hébert-Dufresne. Temporal and probabilistic comparisons of epidemic interventions. Bulletin of Mathematical Biology, 85 (12):??, December 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01220-w.

Shea:2023:RAC

[541] Jacob Shea, Lisa Davis, Bright Quaye, and Tomas Gedeon. Ribosome abundance control in prokaryotes. Bulletin of Mathematical Biology, 85 (12):??, December 2023. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01212-w.

Simmons:2023:MTM

[542] Emily S. G. Simmons, Arielle M. Cooley, Joshua R. Puzey, and Gregory D. ConradiSmith. A multigenerational Turing model reproduces transgressive petal spot phenotypes in hybrid Mimulus. Bulletin of Mathematical Biology, 85(12):??, December 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01223-7.

Li:2023:FTW

[543] Bingtuan Li and Garrett Otto. Forced traveling waves in a reaction– diffusion equation with strong Allee effect and shifting habitat. Bulletin of Mathematical Biology, 85(12):??, December 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/article/10.1007/s11538-023-01221-9.

Chumley:2023:NDM

[544] Max M. Chumley, Firas A. Khasawneh, Andreas Otto, and Tomas Gedeon. A nonlinear delay model for metabolic oscillations in yeast cells. Bulletin of Mathematical Biology, 85(12):??, December 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-023-01227-3.

Pei:2023:OTC

[545] Yongzhen Pei, Yunfei Lv, Changguo Li, and Dandan Fang. Optimization therapy by coupling intermittent androgen suppression with impulsive chemotherapy for a prostate cancer model. Bulletin of Mathematical Biology, 85(12):??, December 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01228-2.

Fung:2023:TCI

[546] Tak Fung, Hannah E. Clapham, and Ryan A. Chisholm. Temporary crossimmunity as a plausible driver of asynchronous cycles of dengue serotypes. Bulletin of Mathematical Biology, 85(12):??, December 2023. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-023-01226-4.

Katriel:2024:OAT

[547] Guy Katriel. Optimizing antimicrobial treatment schedules: Some fundamental analytical results. Bulletin of Mathematical Biology, 86(1):??, January 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01230-8.

Shuttleworth:2024:EQP

[548] Joseph G. Shuttleworth, Chon Lok Lei, Dominic G. Whittaker, Monique J. Windley, Adam P. Hill, Simon P. Preston, and Gary R. Mirams. Empirical quantification of predictive uncertainty due to model discrepancy by training with an ensemble of experimental designs: an application to ion channel kinetics. Bulletin of Mathematical Biology, 86(1):??, January 2024.

CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01224-6.

Patterson:2024:PFM

[549] Denis Patterson, Simon Levin, Ann Carla Staver, and Jonathan Touboul. Pattern formation in mesic savannas. Bulletin of Mathematical Biology, 86(1):??, January 2024. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01231-7.

FragaDelfinoKunz:2024:NAP

[550] Camile Fraga Delfino Kunz, Alf Gerisch, James Glover, Denis Headon, Kevin John Painter, and Franziska Matthäus. Novel aspects in pattern formation arise from coupling Turing reaction–diffusion and chemotaxis. Bulletin of Mathematical Biology, 86(1):??, January 2024. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-023-01225-5.

Tserunyan:2024:ITA

[551] Vardges Tserunyan and Stacey Finley. Information-theoretic analysis of a model of CAR-4-1BB-mediated NFκB activation. Bulletin of Mathematical Biology, 86(1):??, January 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01232-6.

Kiss:2024:TIN

[552] Istvan Z. Kiss, Luc Berthouze, and Wasiur R. KhudaBukhsh. Towards inferring network properties from epidemic data. Bulletin of Mathematical Biology, 86(1):??, January 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01235-3.

Dyck:2024:MII

[553] Cheryl Dyck, Kathryn V. Isaac, and Leah Edelstein-Keshet. Models for implant-induced capsular contracture post breast cancer surgery. Bulletin of Mathematical Biology, 86(1):??, January 2024. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/article/10.1007/s11538-023-01236-2.

Murphy:2024:FGC

[554] Ryan J. Murphy, Gency Gunasingh, Nikolas K. Haass, and Matthew J. Simpson. Formation and growth of co-culture tumour spheroids: New compartment-based mathematical models and experiments. Bulletin of Mathematical Biology, 86(1):??, January 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01229-1.

Bellver-Arnau:2024:DME

[555] Jesús Bellver-Arnau, Alessandro Margheri, and Carlota Rebelo. Dynamics of a mosquito egg–larvae model with seasonality. Bulletin of Mathematical Biology, 86(1):??, January 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01238-0.

Cardona:2024:GOT

[556] Gabriel Cardona, Gerard Ribas, and Joan Carles Pons. Generation of orchard and tree–child networks. Bulletin of Mathematical Biology, 86 (1):??, January 2024. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01239-z.

Bergman:2024:CAB

[557] Daniel R. Bergman, Kerri-Ann Norton, Harsh Vardhan Jain, and Trachette Jackson. Connecting agent-based models with high-dimensional parameter spaces to multidimensional data using SMoRe ParS: a surrogate modeling approach. Bulletin of Mathematical Biology, 86(1):??, January 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01240-6.

Dadashova:2024:LIA

[558] Kamala Dadashova, Ralph C. Smith, and Mansoor A. Haider. Local identifiability analysis, parameter subset selection and verification for a minimal brain PBPK model. Bulletin of Mathematical Biology, 86 (2):??, February 2024. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01234-4.

Diez:2024:TPF

[559] Antoine Diez, Andrew L. Krause, Philip K. Maini, Eamonn A. Gaffney, and Sungrim Seirin-Lee. Turing pattern formation in reaction–crossdiffusion systems with a bilayer geometry. Bulletin of Mathematical Biology, 86(2):??, February 2024. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01237-1.

Liu:2024:DLV

[560] Weiwei Liu, Jie Liu, and Shanshan Chen. Dynamics of Lotka–Volterra competition patch models in streams with two branches. Bulletin of Mathematical Biology, 86(2):??, February 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01243-3.

Han:2024:SMA

[561] Lifeng Han and John Fricks. A semi-Markov approach to study a group of kinesin motors. Bulletin of Mathematical Biology, 86(2):??, February 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01241-5.

Henley:2024:BMC

[562] Lucy Henley, Owen Jones, Fiona Mathews, and Thomas E. Woolley. Bat motion can be described by leap frogging. Bulletin of Mathematical Biology, 86(2):??, February 2024. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01233-5.

Abo:2024:SDS

[563] Stéphanie M. C. Abo, Elisa Casella, and Anita T. Layton. Sexual dimorphism in substrate metabolism during exercise. Bulletin of Mathematical Biology, 86(2):??, February 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01242-4.

Wang:2024:CTB

[564] Ren-Yi Wang and Marek Kimmel. A countable-type branching process model for the tug-of-war cancer cell dynamics. Bulletin of Mathematical Biology, 86(2):??, February 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01245-1.

Browning:2024:PRP

[565] Alexander P. Browning, Thomas D. Lewin, Ruth E. Baker, Philip K. Maini, Eduardo G. Moros, Jimmy Chaudell, Helen M. Byrne, and Heiko Enderling. Predicting radiotherapy patient outcomes with real-time clinical data using mathematical modelling. Bulletin of Mathematical Biology, 86(2):??, February 2024. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01246-0.

Arabameri:2024:UIC

[566] Abazar Arabameri and Samaneh Arab. Understanding the interplay of CAR–NK cells and triple-negative breast cancer: Insights from computational modeling. Bulletin of Mathematical Biology, 86(2):??, February 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
023-01247-z.

Krause:2024:TIE

[567] Andrew L. Krause, Eamonn A. Gaffney, Thomas Jun Jewell, Václav Klika, and Benjamin J. Walker. Turing instabilities are not enough to ensure pattern formation. Bulletin of Mathematical Biology, 86(2):??, February 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01250-4.

Vaghy:2024:KFV

[568] Mihály A. Vághy, Irene Otero-Muras, Manuel Pájaro, and Gábor Szederkényi. A kinetic finite volume discretization of the multidimensional PIDE model for gene regulatory networks. Bulletin of Mathematical Biology, 86(2):??, February 2024. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/article/10.
1007/s11538-023-01251-3.

Meyer:2024:MYO

[569] Alexander D. Meyer, Alan Hastings, and John L. Largier. Making your own luck: Weak vertical swimming improves dispersal success for coastal marine larvae. Bulletin of Mathematical Biology, 86(3):??, March 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01252-2.

Collienne:2024:RSP

[570] Lena Collienne, Chris Whidden, and Alex Gavryushkin. Ranked subtree prune and regraft. Bulletin of Mathematical Biology, 86(3):??, March 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-023-01244-2.

Husar:2024:LDM

[571] Kateryna Husar, Dana C. Pittman, Johnny Rajala, Fahad Mostafa, and Linda J. S. Allen. Lyme disease models of tick–mouse dynamics with seasonal variation in births, deaths, and tick feeding. Bulletin of Mathematical Biology, 86(3):??, March 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

article/10.1007/s11538-023-01248-y.

Hindel:2024:GKM

[572] Stefan Hindel. A generalized kinetic model of fractional order transport dynamics with transit time heterogeneity in microvascular space. Bulletin of Mathematical Biology, 86(3):??, March 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01255-z.

Hamley:2024:IDW

[573] Jonathan I. D. Hamley, Guido Beldi, and Daniel Sánchez-Taltavull. Infectious disease in the workplace: Quantifying uncertainty in transmission. Bulletin of Mathematical Biology, 86(3):??, March 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/article/10.1007/s11538-023-01249-x.

Potts:2024:DBL

[574] Jonathan R. Potts and Kevin J. Painter. Distinguishing between longtransient and asymptotic states in a biological aggregation model. Bulletin of Mathematical Biology, 86(3):??, March 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-023-01254-0.

Wifvat:2024:OCR

[575] Kathryn Wifvat, Erika T. Camacho, Matthias Kawski, Thierry Léveillard, and Stephen Wirkus. Optimal control with RdCVFL for degenerating photoreceptors. Bulletin of Mathematical Biology, 86(3):??, March 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

https://link.springer.com/article/10.1007/s11538-024-01256-6.

Ciavolella:2024:MMD

[576] Giorgia Ciavolella, Nathalie Ferrand, Michéle Sabbah, Benoît Perthame, and Roberto Natalini. A model for membrane degradation using a gelatin invadopodia assay. Bulletin of Mathematical Biology, 86(3):??, March 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11538-
024-01260-w.

Mallela:2024:IVS

[577] Abhishek Mallela, Ye Chen, Yen Ting Lin, Ely F. Miller, Jacob Neumann, Zhili He, Kathryn E. Nelson, Richard G. Posner, and William S. Hlavacek. Impacts of vaccination and severe acute respiratory syndrome coronavirus 2 variants alpha and delta on coronavirus disease 2019 transmission dynamics in four metropolitan areas of the United States. Bulletin of Mathematical Biology, 86(3):??, March 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-024-01258-4.

Wang:2024:OME

[578] Yue Wang, Blerta Shtylla, and Tom Chou. Order-of-mutation effects on cancer progression: Models for myeloproliferative neoplasm. Bulletin of Mathematical Biology, 86(3):??, March 2024. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/article/10.1007/s11538-024-01257-5.